### v1

# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

Abstract
Drones are increasingly used in forestry to capture high-resolution remote sensing data. While operations above the forest canopy are already highly automated, flying inside forests remains challenging, primarily relying on manual piloting. Inside dense forests, reliance on the Global Navigation Satellite System (GNSS) for localization is not feasible. Additionally, the drone must autonomously adjust its flight path to avoid collisions. Recently, advancements in robotics have enabled autonomous drone flights in GNSS-denied obstacle-rich areas. In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open-source methods and validating its performance for data collection inside forests. The autonomous flight capability was evaluated through multiple test flights in two boreal forest test sites. The tree parameter estimation capability was studied by conducting diameter at breast height (DBH) estimation using onboard stereo camera data and photogrammetric methods. The prototype conducted flights in selected challenging forest environments, and the experiments showed excellent performance in forest reconstruction with a miniaturized stereoscopic photogrammetric system. The stem detection algorithm managed to identify 79.31% of the stems. The DBH estimation had a root mean square error (RMSE) of 3.33 cm (12.79%) and a bias of 1.01 cm (3.87%) across all trees. For trees with a DBH less than 30 cm, the RMSE was 1.16 cm (5.74%), and the bias was 0.13 cm (0.64%). When considering the overall performance in terms of DBH accuracy, autonomy, and forest complexity, the proposed approach was superior compared to methods proposed in the scientific literature. Results provided valuable insights into autonomous forest reconstruction using drones, and several further development topics were proposed.
keywords:
Robotic; drone; forest; photogrammetry; DBH
articletype: ARTICLE PREPRINT
1 Introduction
Over the last decade, the use of drones in forestry has rapidly increased, offering a practical solution for forest data collection both above and inside the forest canopy (Puliti et al. 2015; Hyyppä et al. 2020a; Liang et al. 2019). Modern drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems. However, their positioning systems rely exclusively on Global Navigation Satellite Systems (GNSS) signals, limiting autonomous flights to open areas where these signals are reliable. In dense forests, GNSS localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al. 2010). Additionally, navigating inside forests requires the capability to autonomously avoid obstacles such as trees and bushes along the flight path. Concequently, under-canopy data collection with drones is still primarily controlled manually by human pilots (e.g. Hyyppä et al. 2020a; Liang et al. 2019; Krisanski, Taskhiri, and Turner 2020).
Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in GNSS-denied, cluttered environments. In recent years, several solutions utilizing various sensor setups and algorithms for autonomous navigation under forest canopies have been proposed (e.g. Liu et al. 2022; Loquercio et al. 2021; Zhou et al. 2022). However, these methods are typically validated by reporting a single test flight in a specific forest environment. Most studies assess the effects of varying object density only in simulators, with only a couple of the methods tested through multiple test flights inside real forests (Loquercio et al. 2021; Ren et al. 2022; Liu, Ren, and Zhang 2024). Therefore, a significant research gap remains regarding the applicability and reliablity of these algorithms in diverse forest environments.
The autonomous under-canopy drone solutions proposed in the literature typically focus on aerial robotics and primarily utilize the collected sensor data for navigation. However, recent research has begun to explore boarder applications. For example, Liu et al. (2022) employed a semantic simultaneous localization and mapping (SLAM) algorithm to calculate the number of trees in the explored area. Similarly, Liang et al. (2024) demonstrated the use of autonomous drones for in situ forest observations under the canopy. While their work estimated tree attributes based on data collected during two test flights, it did not evaluate the reliability of autonomous flight itself. Consequently, a research gap remains regarding the borader application of autonomous drones for forestry applications.
Remote sensing measurements within forests are increasingly used to capture detailed forest information. Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al. 2007) and mobile solutions e.g. handheld devices (Ryding et al. 2015; Bauwens et al. 2016; Marselis et al. 2016; Balenović et al. 2021), backpacks (Liang et al. 2014b, 2018; Hyyppä et al. 2020b), ground-based vehicles (Forsman, Holmgren, and Olofsson 2016; Liang et al. 2018; Bienert et al. 2018) or manually operated drones (Chisholm et al. 2013; Brede et al. 2017; Wieser et al. 2017; Hyyppä et al. 2020a). Similarly, the use of structure-from-motion (SFM) based point clouds for forest mapping has been demonstrated using different camera solutions (Liang et al. 2014a; Wallace et al. 2016; Kuželka and Surovỳ 2018; Piermattei et al. 2019). Combining these techniques with autonomously navigating drones inside forests represents a highly relevant research opportunity.
To address these gaps and opportunities, this study introduces a novel approach for autonomous data collection and object characterization under unstructured forest canopies. The primary objectives were to assess the performance of an autonomously flying drone in a boreal forest and to demonstrate its capability in measuring tree characteristics. A prototype autonomously flying drone was developed using state-of-the-art open-source algorithms, followed by multiple flight tests to evaluate its performance and suitability for boreal forests. Image data collected by the low-cost onboard stereo-camera was stored for further analysis. The dataset was post-processed using SfM photogrammetry to improve the georeferencing and to generate dense 3D point clouds. Finally, tree detection and diameter at breast height (DBH) estimation were performed on the point clouds, assessing the potential of the low-cost autonomous system for forest parameter estimation.
The study is structured as follows. Section 2 reviews recent open-source solutions for autonomous drone flights inside forests and state-of-the-art methods forest parameter estimation using drones flying inside forests. Section 3 describes the algorithms, hardware, experiments, and methods for autonomous navigation, tree detection, and DBH estimation. Experimental results are presented in Section 4. Section 5 discusses the results, potential improvements, and concludes the study.
2 Related work
2.1 Autonomous flying inside forests
Autonomous navigation in obstacle-rich environments has been an actively researched topic in robotics during the last decades. Whereas the majority of the research has tried to tackle the 2D path planning problem for autonomous ground vehicles, in recent years increasing number of methods for autonomous path planning and obstacle avoidance have also been published for drones. Although many of the methods proposed in the literature are only tested in simulations or indoor conditions, some of the methods have been also successfully flight-tested inside forests. Some of the authors have also open-sourced their methods to encourage the development of the field. In this section, recent open-source methods successfully flight-tested inside a forest are summarized.
Campos-Macías et al. (2021) proposed a solution that focuses on computational efficiency that can be run in real time with a low-power CPU. The solution used a stereo tracking camera to run Visual-Inertial Odometry (VIO) for pose estimation and a point cloud from an active stereo depth camera to create a 3D voxel map. Collision-free trajectories were planned to the free regions on the map. In addition to multiple indoor environments, the system was tested in a sparse mixed forest. However, the flight distance in the forest test was only 11.58 m.
Liu et al. (2022) proposed a method that aims at large-scale autonomous flights in forest environments. The proposed system used an inertial measurement unit (IMU) and a stereo camera to run VIO and 3D light detection and ranging (LiDAR) to run a semantic simultaneous localization and mapping (SLAM) algorithm. Semantic SLAM was developed to correct drift in the VIO estimate. The system was reported to fly an 800 m long autonomous flight, but the semantic SLAM algorithm was only tested in manual flight. The documented flight tests were performed in a North American pine forest. The forest was relatively dense, but it had a minimal amount of low branches and understory vegetation.
Ren et al. (2022) proposed another LiDAR-based method that generates sphere-shaped safe flight corridors from the point clouds. The method was targeting for high-speed flying in unknown environments. FAST-LIO2 (Xu et al. 2022) was utilized in the pose estimation. In total, 12 test flights with varying flight speeds were performed in a sparse Chinese deciduous forest. The maximum flight speed was 13.7 m/s and 11 of 12 test flights were successful, but the trajectory lengths were not reported. Liu, Ren, and Zhang (2024) have also proposed another method for the same LiDAR-based platform targeting especially safe navigation in the presence of dynamic obstacles and external disturbances with an integrated planning and control framework. Planning was performed within a local grid map and FAST-LIO2 (Xu et al. 2022) was utilized in the pose estimation. In addition to indoor environments, the system was also tested with 10 successful flights at varying speeds in the same forest as (Ren et al. 2022). The maximum flight speed was 5.86 m/s, and the reported trajectory length was 58.64 m.
Instead of using a single drone, the solution proposed by Ahmad et al. (2022) focused on the navigation of drone swarms. The proposed approach is fully decentralized and tries to mimic the collective navigation behavior of natural flocks and swarms. The physical drone hardware used in the tests incorporated a LiDAR for running the SLAM algorithm by Kohlbrecher et al. (2011), a laser rangefinder as an altimeter, and an Ultra Violet Direction And Ranging (UVDAR) system for relative localization between the drones of the swarm. The approach was tested in a swarm flight inside a forest, but the test forest was very clean containing no low branches or understory vegetation. The flight distance was approximately 40 m.
Zhou et al. (2022) also proposed a solution concentrating on swarm flights. In the proposed solution the drones use VIO for localization and probabilistic maps for obstacle avoidance. Swarm members also share their trajectories over a wireless network and correct the VIO drift based on the relative distances between the drones measured with Ultra Wide Band (UWB) sensors. The proposed solution was tested with palm-sized drones which used only one stereo camera to run both mapping and VIO. In the flight tests, a swarm of drones was successfully flying approximately 80 m through a dense bamboo forest.
The solution proposed by Loquercio et al. (2021) replaced different sensing, mapping, and planning subtasks with a single neural network directly mapping depth images and a pose estimate from VIO to a feasible collision-free trajectory for the drone. The system was trained entirely with simulated data, and the scope of the research was high-speed autonomous flying. Real-world flight tests were performed with a drone using an active stereo depth camera to produce the depth images, and a stereo tracking camera to run VIO. Flight tests were performed in various environments containing also Central-European mixed forests. The forest densities were not reported, but the forests were relatively sparse with a low amount of understory vegetation. The evaluated trajectories were a 40 m long straight line and a circle with a 6 m radius, and multiple test flights were conducted with flight speeds varying from 3 to 10 m/s. The flight success rate was 100 % with a flight speed of 3 m/s and 5 m/s, 80 % with a flight speed of 7 m/s, and 60 % with a flight speed of 10 m/s
Nguyen et al. (2023) proposed also a deep-learning-based method that utilizes neural networks to directly map depth images and partial pose estimates of the drone to a series of actions leading to a collision-free trajectory. The proposed system was also trained entirely in a simulator. In addition to the indoor flight tests, the performance assessment was reported for a single 60 m long trajectory through a dense Finnish pinewood forest making the test environment the most similar to the target environments in this study.
2.2 DBH estimation from under-canopy drone data
In a forest inventory, DBH is one of the most important tree attributes to be estimated since it provides a measure for the size of a tree and also enables the prediction of the total stem volume (Laasasenaho 1982). During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH. Even though several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al. (2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al. (2018); Piermattei et al. (2019); Xu, Shen, and Cao (2023)) and above-canopy photogrammetric systems (e.g., Wallace et al. (2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023). If the operation of such under-canopy drone systems can be automated in the future, they would provide a truly efficient solution for the acquisition of accurate forest data.
Kuželka and Surovỳ (2018) manually operated a drone system based on a DJI Phantom 4 Pro within two forest plots to produce low-noise photogrammetric point clouds, from which the DBH of individual trees could be estimated with an RMSE between 2.6 cm and 5.3 cm depending on the plot and the applied analysis method. The studied plots were sparse with only 270-290 trees/ha. Krisanski, Taskhiri, and Turner (2020) performed manual flights of a drone system inside two forest plots with a stem density of approximately 1000 trees/ha to acquire high-density SfM point cloud data, from which the DBH of individual trees could be measured with a RMSE of 1.1-2.1 cm. Shimabuku, Konoshima, and Ota (2023) achieved an RMSE of 0.4-0.7 cm for DBH estimation on two small forest plots using SfM point clouds acquired with an under-canopy flying drone system. The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction. They also demonstrated stem volume estimates of individual trees with a RMSE of 15.8-17.2%.
In addition to photogrammetric approaches, LiDAR-based methods have been studied extensively for the purpose of efficiently and accurately estimating various tree attributes, such as DBH, tree height and stem volume. To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al. 2015; Bauwens et al. 2016; Cabo et al. 2018; Chen et al. 2019; Hyyppä et al. 2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation. Even though a LiDAR can be mounted to a variety of other mobile platforms besides a drone, the focus is on the use of under-canopy drones for DBH estimation in this brief literature review.
Chisholm et al. (2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually-piloted drone flying under the forest canopy. They detected 12 trees in the vicinity of the drone trajectory and reported a RMSE of 25.1% for the DBH estimates of the detected trees.
Hyyppä et al. (2020a) collected high-quality point cloud data using a Kaarta Stencil-1 laser scanner mounted horizontally at the bottom of a manually-piloted drone. They demonstrated DBH estimates with a RMSE of 0.6-0.9 cm (2%-3%) on two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha. Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory. As the laser scanner was mounted horizontally in their study, the limited field of view made it difficult to obtain direct measurements of tree height. In a follow-up study by Hyyppä et al. (2021), this issue was resolved by instead integrating a rotating laser scanner to an under-canopy flying drone, which resulted in accurate measurements for both stem diameter and tree height. Wang et al. (2021) combined data from above- and under-canopy flights of a manually-piloted drone, resulting in an RMSE of 7.95 cm (33.35%) and a bias of 3.72 cm (15.61%) for the DBH.
Muhojoki et al. (2024) compared the accuracy of the estimated tree attributes from several LiDAR-based mobile systems, including two under-canopy drone systems, a system by Hovermap (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a system by Deep Forestry (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA). Both drones were flown manually in Muhojoki et al. (2024) to collect data in sparse density forest plots. The RMSE and bias of DBH estimates for the Deep Forestry’s and Hovermap’s drones, with bias removal method introduced by Muhojoki et al. (2024), averaged RMSE of 1.5 cm (6.7 %) and bias of -0.7 cm (3.0 %) and RMSE of 1.6 cm (6.9 %) and bias of -0.1 cm (-0.6 %), respectively. Without using bias removal method, RMSE and bias for both drones were significantly higher. Muhojoki et al. (2024) also demonstrated that the under-canopy drone systems can provide roughly as accurate tree attributes as other under-canopy MLS methods, such as handheld LiDARs.
Recently, Liang et al. (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha. To enable the autonomous flight, a predefined global path was planned, to which the drone system performed real-time adjustments during the flight according to real-time environment conditions. When the flight path covered most of the test plot, they achieved an RMSE of 5.1 cm for the DBH of individual trees, and a similar RMSE for the stem curve.
In addition to under-canopy flying drones, LiDARs have also been integrated to above-canopy flying drones capable of direct stem diameter measurements, though typically with a significantly higher RMSE and a lower tree detection rate due to the larger measurement distance and the upper canopy obstructing the tree stems (Wieser et al. 2017; Brede et al. 2017; Kuželka, Slavík, and Surovỳ 2020; Vandendaele et al. 2021; Feng et al. 2022).
A summary of relevant studies on under canopy forest inventory studies is presented in Table 1.
Table 1: Accuracy of the stem diameter breast height (DBH) estimations in other studies. SfM-MVS: Structure-from-motion and multi-view stereo; ULS: Under canopy laser scanning; ALS: Above canopy laser scanning. *averaged values from three plots
Dataset Type Number of trees RMSE (cm) RMSEBias (cm) Bias%
Shimabuku et al. (2023) plot I Drone SfM-MVS 7 0.4 3.2 - -
Shimabuku et al. (2023) plot II Drone SfM-MVS 18 0.7 3.6 - -
Kuželka and Surovỳ (2018) Drone SfM-MVS 119 2.63-5.31 7.01-13.3 0.27-3.77 0.69-10.11
Krisanski et al. (2020) Drone SfM-MVS 29 4.1 - - -
Liang et al. (2024) traj. I ULS 143 9.43 63.41 7.17 48.21
Liang et al. (2024) traj. II ULS 143 4.48 21.37 5.01 14.91
Hyyppä et al. (2020a) plot I ULS 42 0.6 2.3 0.34 1.3
Hyyppä et al. (2020a) plot II ULS 43 1.1 3.5 0.12 0.4
Wang et al. (2021) ULS&ALS - 7.95 33.35 3.72 15.61
Muhojoki et al. (2024) * ULS 50 1.50-1.60 6.70-6.90 0.10-0.70 0.60-3.00
3 Materials and methods
3.1 Autonomous navigation algorithms
In this study, the objective was to develop a prototype drone capable of autonomously avoiding obstacles under the forest canopy and operating without GNSS. The drone was required to autonomously fly from the takeoff location to a specified goal point while recording sensor data during the flight. This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
Among the open-source solutions presented in Section 2.1, the solution by Zhou et al. (2022) was chosen as the base of the prototype. The solution was proven to work with only one stereo camera making it possible to build a smaller and lighter platform. Furthermore, cameras are typically lighter and more power-efficient than LiDARs (Reddy Cenkeramaddi et al. 2020) making them an attractive solution for small under-canopy drones, for which the weight and battery consumption must be optimized. This enables building a compact system that is capable of operating in dense forests.
The main details of the algorithms used in the chosen solution are described briefly in the following and the complete system architecture of the prototype is presented in Figure 1. For more detailed information the readers are referred to the cited original sources.
3.1.1 EGO-Planner-v2
The source code of the planner algorithm by Zhou et al. (2022) was named in Github as EGO-Planner-v2 (ZJU FAST Lab Team 2022a). Even though the EGO-Planner-v2 is designed for swarm flying, in this study, only a single drone system was used. The planner is running on top of the Robot Operating System (ROS) (Quigley et al. 2009). The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985). The mapping module uses depth images as an input and the system adopts the fixed-sized circular buffers proposed by Usenko et al. (2017) for maintaining the local map. The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map respectively. Virtual floor and virtual ceiling restrict the altitude where the algorithm is allowed to plan paths.
In the trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al. 2022) trajectory representation designed for differentially flat systems, such as quadrotors (Mellinger and Kumar 2011). Planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (Zhou et al. 2022). The trajectories are optimized iteratively by minimizing the weighted sum of metrics defined for trajectory smoothness and flying time. The feasibility of the planned trajectory is forced by deforming the shape to avoid obstacles and by restricting the magnitudes of trajectory velocity, acceleration, and jerk to fulfill the dynamical constraints of the drone. For computational efficiency, the trajectory constraints are enforced via integrals of penalty functions with large penalty weights, and the integrals are evaluated by a finite sum of equally spaced samples along the timeline. In a simplified form, the optimization of time-dependent objectives $J$ in one piece of the trajectory can be represented by
${\min}\sum_{x}\lambda_{x}J_{x}$ (1)
where where $J_{x}$ are the penalty terms and $\lambda_{x}$ are the corresponding weights. Subscripts $x=\{s,t,d,o,u\}$ stand for the trajectory smoothness, total flying time, dynamical feasibility, obstacle avoidance, and the uniform distribution of constraint evaluation points along the trajectory, respectively. In addition to the above mentioned penalty terms, EGO-Planner-v2 includes also penalty terms for the swarm operations, but those are omitted here. A* is utilized as the global path-searching algorithm given as an input for the optimization algorithm.
The trajectory planning is continuously performed within a defined local planning distance towards a user-defined global goal until the goal is reached. Since the safety constraints are modeled as integrals in the penalty functions, the feasibility of the trajectory needs to be verified with a post-check after planning (Zhou et al. 2022). Due to the non-static probabilistic map, the post-check after trajectory planning guarantees feasibility only at a certain moment. Hence, a collision-check process is continuously running in the background. If that process detects a potential collision, a trajectory replanning is activated immediately. If the replanning fails and the predicted time to collision is under a given threshold, the planner performs an emergency stop. After the drone has stopped, the trajectory planner tries to start up again and find a new feasible trajectory.
3.1.2 VINS-Fusion
The VIO algorithm used by Zhou et al. (2022) was VINS-Fusion (Qin, Li, and Shen 2018; Qin et al. 2019b)(VINS stands for ”Visual-Inertial system”), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
VINS-Fusion detects Shi-Tomasi corner features (Shi and Tomasi 1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (Lucas and Kanade 1981). IMU data is used for preintegration, i.e. position, velocity, and orientation are integrated from IMU measurements between two frames. The state optimization is performed within a sliding window including the position, velocity, rotation, and IMU biases. To reduce the computational complexity, old measurements are marginalized into a prior of the estimation process. Certain frames are also selected as keyframes which are saved to a global pose graph after being marginalized out from the sliding window. Keyframes are selected based on the number of tracked features and the average parallax between the features in the newest frame and the latest keyframe.
VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate. The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al. 2010) to DBoW2 (Galvez-López and Tardos 2012) bag-of-word database. After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses. In addition to sliding window relocalization, also the global pose graph is optimized after loop detections to enforce the global consistency of the poses. Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (Qin and Shen 2018). The method can be used for adjusting the time offset $t_{d}$ between the camera and the IMU online.
In the swarm operations of the original article (Zhou et al. 2022), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion. However, with a single drone of this study that function is omitted.
The parameters used in the configuration of VINS-Fusion were chosen so that the real-time performance was guaranteed. In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion causing instability to the estimates, so the IMU noise values were increased more. All VINS-Fusion configuration parameters are listed in Table 2.
Table 2: VINS-Fusion parameters.
Parameter Value
Image resolution $640\times 360$
Max number of tracked features 150
Min distance between features 40 px
RANSAC threshold 1 px
Max solver iteration time 0.04 ms
Max solver iterations 8
Keyframe parallax threshold 10 px
Accelerometer noise SD 0.2
Gyroscope noise SD 0.05
Accelerometer random walk SD 0.002
Gyroscope random walk SD 0.0005
3.1.3 Tracking controller
The original EGO-Planner-v2 article (Zhou et al. 2022) did not specify the utilized method for trajectory tracking. In this study, an open-source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used. The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner, and converts those to attitude and thrust setpoint. The setpoints are forwarded to PX4 flight controller software (PX4 Community 2023) responsible for low level attitude and angular rate controls.
Figure 1: The system architecture used in the data collection. In comparison to the original EGO-Planner-v2 article by Zhou et al. (2022), UWB-based Drift correction and Wireless trajectory broadcast network modules used for swarm navigation were omitted in this study. Stereo images were downsampled for VIO to guarantee real-time estimation. Full-resolution stereo images were also saved and post-processed to point clouds used in the DBH estimation.
3.2 Drone hardware
In the physical robot drone prototype, an onboard computer NVIDIA Jetson Orin NX (Santa Clara, California, USA) was used. The used PX4-compliant autopilot was Holybro Kakute H7 (Hongkong, China), which contained also the IMU. The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the Mavros-protocol was used for communication between the onboard computer and the autopilot. As a camera, Intel RealSense D435 (Intel Realsense Development Team 2023)(Keselman et al. 2017) was used. The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
RealSense D435 is based on active stereoscopy and the camera has an infrared laser projector emitting dots that can be used for triangulation by stereo camera. This can improve the triangulation in low-light conditions and to homogeneous surfaces that contain no features. However, the dots of the laser emitter cause problems for VIO since they are moving with the camera. In the previous setup described by (Karjalainen et al. 2023), a modified version of ROS-RealSense driver was used by Zhou et al. (2021b) that turns the laser emitter of the camera constantly on and off making it possible to produce high-quality depth images in every other frame, and stereo images without the laser dots in every other frame. However, this solution comes with the drawback of halving the output frame rate of the camera.
In the first test flights described by (Karjalainen et al. 2023), difficulties were encountered in detecting small branches with a depth camera resolution of $640\times 480$, leading to the decision to increase the depth camera resolution to a maximum value of $1280\times 720$. Stereo images for VIO were downsampled to resolution of $640\times 360$ to ensure real-time performance. As the maximum frame rate of D435 at a resolution of $1280\times 720$ is 30 Hz, the modified driver version would have resulted in a low frame rate of 15 Hz. To ensure the VIO performance in quick turns, the frame rate of 30 Hz was preferred, and hence disabled the IR emitter making the camera function as a passive IR stereo camera. The choice can be further justified with the fact that the usefulness of the IR emitter in forest conditions is questionable, since sunlight contains infrared light that can interfere with the emitted artificial features(Kuan, Ee, and Wei 2019), and the forests typically do not contain large homogeneous surfaces. The Realsense D435 onboard camera was configured to use the builtin automatic exposure control.
The camera-IMU system was calibrated with an open-source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012)(Furgale, Rehder, and Siegwart 2013)(Maye, Furgale, and Siegwart 2013). In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were calculated from recorded data where Aprilgrid (Olson 2011) was used as a calibration target. The IMU noise parameters needed for the calibration were first obtained with open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the obtained values were multiplied with 10 before the calibration as suggested in the project wiki of kalibr (ETH Zürich Autonomous Systems Lab 2023).
The computer, the camera, and the autopilot were attached to a 330 mm drone frame. The propellers had two pieces of 6 cm long blades each. The measured weight of the drone including the motors and the onboard computer was 791 g without the batteries, and 1153 g with the batteries. An overview of the used drone hardware is presented in Figure 2.
Figure 2: The drone hardware used in the flight tests with a forward-facing Realsense D435 camera. Nvidia Jetson Orin NX onboard computer is hidden in the case under the drone frame. Holybro Kakute H7 Flight controller unit is located inside the frame. Batteries (not visible in the picture) are mounted on the top of the drone frame.
3.3 Test sites
The performance of the system was evaluated in three different Boreal forest areas. The flight tests of the system were performed in two test areas located in Evo, Finland (61.19°N, 25.11°E). The flight paths went through well-documented forest sample plots that have been widely used in previous studies, such as in Liang et al. (2018, 2019); Wang et al. (2019); Hyyppä et al. (2020a, c), and are classified into three complexity categories, i.e., “easy”, “medium”, and “difficult” based on, mainly tree density and amount of the understory vegetation in the plots.
The first plot has been classified as ”medium”, and the second one has been classified as ”difficult”. The medium difficulty plot had a tree density of 650 trees/ha and a mean diameter at breast height (DBH) of 28 cm. The most common tree species in the plot were Norway spruce (Picea abies (L.) H.Karst.)(81.2%) and Scots pine (Pinus sylvestris L.)(9.4%). The difficult forest had a tree density of 2000 trees/ha and a mean DBH of 17 cm. The most common tree species in the plot were Norway spruce (64.4%) and aspen (Populus tremula, L.) (18.3%). Only trees with a DBH greater than 5 cm were included in these statistics. From now on, these two test areas are called ”Evo-medium” and ”Evo-difficult”, respectively.
For georeferencing purposes, circular coded targets provided by Agisoft Meatahsape (Agisoft LLC, St. Petersburg, Russia) were placed within the forest plots. The targets had a diameter of 24.7 cm, and the middle dots had a size of 7 cm. Relative distances between the coded targets were measured with Leica Nova TS60 total station with a precision of approximately 2mm (Leica Geosystems 2020). Figure 5 shows coded targets in ”Evo-difficult” forest during the test flights.
The third test area was a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E). Aim of the experiment there was to test the loop detection performance of VINS-Fusion by walking with the drone in an environment without coded targets. Estimated average tree density in the area was approximately 2000 trees/ha.
Figure 3: View from the first takeoff location in the Evo-medium test area. The takeoff location was changed after five test flights.
Figure 4: View from the takeoff location in the Evo-difficult test forest
3.4 Flight tests
Test flights in Evo were conducted on 18th October, 2023. Multiple test flights were performed to evaluate the forest-type effects on the performance of autonomous navigation. In total, seven test flights were conducted in the ”Evo-medium” test area and nine test flights in the ”Evo-difficult” test area. In each test environment, either the takeoff location or the takeoff heading was altered during the test flight campaign to force the drone to follow different paths. In the ”Evo-medium” test area, the takeoff location was changed after five test flights. In the ”Evo-difficult” test area the takeoff heading was changed after six test flights. The target flight velocity was 1 m/s in all flight tests. The goal was set 36 m forward in the first takeoff location in the ”Evo-medium” test area, 34 m forward in the second takeoff location in the ”Evo-medium” test area, and 42 m forward in the takeoff location in the ”Evo-difficult” test area. The maximum flight height was set to 2.25 m relative to the takeoff location in the ”Evo-medium” test area. For the first test flight in the ”Evo-difficult” test area, the maximum flight height was set to 2.5 m with respect to the takeoff location and was increased to 2.75 m for the rest of the flights.
The campaign to evaluate the loop detection capability took place on 26th January 2024. Stereo images and IMU data were collected during two walks inside the forest while carrying the drone by hand. The paths of the walks were approximately 340 m and 240 m long, each containing multiple loops where the same scene was observed multiple times. The data collection took place during winter, with the forest covered in snow.
Figure 5: Control targets for reference trajectory generation in ”Evo-difficult” forest. Targets were spread widely to the test site to increase the probability that there would be targets visible in the recorded camera data regardless the path that the robot takes.
3.5 Structure-from-Motion Multi-View Stereo based reconstruction
Photogrammetric Structure-from-Motion Multi-View Stereo (SfM-MVS) processing was conducted in post-processing mode to generate 3D point clouds for DBH estimation and to generate reference trajectories for evaluating VIO performance.
Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St. Petersburg, Russia) (Metashape Development Team 2021) following similar processing steps as introduced by Viljanen et al. (2018). Every sixth stereo pair were used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m. In the image orientation processing, quality setting was set to high meaning that two times downsampled images were used; the settings for number of key points and tie points per image were 40,000 and 4,000, respectively. To enhance the quality of the post-processed reference trajectories, ground control points (GCPs) were used in the processing with an accuracy setting of ±$\pm 0.005$ m. Automatic outlier removal was performed using gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy. Subsequently, a bundle-adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP). Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering. Finally, the EOPs and dense point clouds were exported.
Point clouds were derived using datasets from four flights conducted in the ”Evo-medium” test area, each designated as follows: I: the first flight, II: the second flight, III: the third flight, IV: the fourth flight, and I-IV: combined processing of all four flights (Combined). The processing was carried out using all GCPs. Additionally, to evaluate the impacts of GCPs, point clouds were generated from the first flight trajectory processed with different GCP configurations, denoted as: I0GCP: no GCPs, I1GCP: one GCP at the beginning of the trajectory, I2GCP: two GCPs, one at the beginning and one at the end of the trajectory, and I3GCP: three GCPs, positioned at the beginning, middle, and end of the trajectory.
The experiments were conducted in dense forests, where collecting reference data using GNSS was not possible due to the forest canopy blocking the signal. Reference trajectories were generated for the flights using SfM processing. In each case, the referenece trajectory was derived from the first takeoff heading and orientation in both Evo test environments.
3.6 Tree parameter estimation
In this study, trees were detected and their DBHs estimated using an automatic stem diameter estimation algorithm originally proposed by Hyyppä et al. (2020b), and further developed by Hyyppä et al. (2020a) to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS. The method described by Hyyppä et al. (2020a) was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments, and 5-s-long temporal segments. The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus the temporal segmentation was not performed for the TLS data. The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al. 1996)), and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle. The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA). The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation. The diameters were processed into a stem curve by automatically removing clear outliers followed by fitting a smoothing spline to the remaining diameter estimates as a function of height. Finally, the DBH estimates were calculated by interpolating the stem curve at the height of 1.3 m above the ground. For a detailed description of the method, the reader is referred to Hyyppä et al. (2020b, a).
3.7 Performance assessment
3.7.1 Autonomous flying performance
The performance and reliability of the autonomous navigation of the prototype were evaluated based on two factors: the reliability of the obstacle avoidance and the accuracy of the VIO localization.
Obstacle avoidance reliability was evaluated based on the mission success. A mission was considered successful if the drone reached the end of the flight without major collisions. Minor touches to thin vegetation or branches were allowed, if the drone was able to continue the mission after that. If the given goal was inside an obstacle, the mission was considered successful if the drone managed to fly next to the goal. In addition to obstacle avoidance, the smoothness of the trajectory was assessed. The trajectory was considered smooth if the system did not perform any emergency stops during the flight.
The accuracy of the VIO localization was assessed by calculating the Absolute Trajectory Error (ATE), as proposed by Sturm et al. (2012), with respect to the SfM reference trajectories. In this study, ATE was calculated only for the position estimate, and the error in the orientation estimate was not evaluated. ATE was calculated using the trajectory evaluation toolbox by Zhang and Scaramuzza (2018), which aligns the trajectory estimates with the reference trajectories with the method by Umeyama (1991). The root mean square error (RMSE) of the individual state errors is then calculated to generate a single metric value representing the entire trajectory’s accuracy.. In the mathematical form the error can be presented as
$\text{ATE}_{\text{pos}}=\left(\frac{1}{N}\sum_{i=0}^{N-1}||\Delta\mathbf{p}_{i}||^{2}\right)^{\frac{1}{2}}$ (2)
where $\Delta\mathbf{p}_{i}$ is the error between the estimated position and reference trajectory position in a single state.
In the first tests conducted in a simulator as described by (Karjalainen et al. 2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to constant, compared to when the online estimation method (Qin and Shen 2018) was used. For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr. However, the calibration estimate is never perfect and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU. For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests. The estimates were performed with the drone’s onboard computer, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of online estimation.
Since the coded targets are easily detectable, unique, and unnatural objects, the loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets. Since there were no coded targets for forming the reference trajectories, only the ability to detect the loops inside dense natural forest was tested, and the accuracy of the loop corrected trajectory was not evaluated. VINS-Fusion estimate with the loop detection algorithm was ran for the collected data with the drone’s onboard computer, and the effect of amount of features in pictures for loop detection capability was examined.
3.7.2 DBH estimation performance
The accuracy of the DBH estimation from the drone point cloud data was evaluated using Terrestrial Laser Scanning (TLS) data as a reference. The TLS scannings were conducted in April 2020, and tree height, diameter at breast height (DBH), and other parameters were estimated from the dataset. The estimated accuracy of the DBH measurement was less than 1 cm.
Several metrics were used to evaluate the accuracy of the DBH estimation, including Root Mean Square Error (RMSE), Relative RMSE (RMSE%), Bias, Relative Bias (bias%), and Standard Deviation.
The RMSE measures the square root of the average of the squared differences between the estimated and reference DBH values. Bias indicates the average deviation between the estimated and reference DBH values.
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-d_{\text{reference},i}\right)^{2}}$ (3)
where $n$ is the total number of observations, $d_{\text{estimated},i}$ is the estimated DBH, and $d_{\text{reference},i}$ is the reference DBH.
$\text{Bias}=\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-d_{\text{reference},i}\right)$ (4)
Relative RMSE (RMSE%) is computed by dividing the RMSE by the mean of the reference DBH values and multiplying by 100 to express it as a percentage. Relative Bias (Bias%) is determined by dividing the Bias by the mean of the reference DBH values and multiplying by 100.
[…]
Additionally standard deviation was calculated with following equation:
$\text{SD}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-\bar{d}_{\text{reference}}\right)^{2}}$ (7)
where $\bar{d}_{\text{reference}}$ is the mean of the reference DBH values, and SD is the standard deviation.
Additionally, the completeness of stem detection was evaluated by comparing the total number of reference trees found from the drone-collected point cloud data to the total number of reference trees within the point cloud boundary.
$c=\frac{r}{t}\times 100$ (8)
where $c$ represents completeness (%), $r$ is the number of reference trees found, and $t$ is the total number of reference trees.
The impact of tree growth between the collection of drone data and reference measurements was not taken into account. Variability in tree growth, influenced by factors such as species, locations, health, age, is acknowledged as a source of error. Considering that the tree growth was relatively constant in the test area, forming a major component of the bias in the estimates. Additionally, some reference trees fell between the measurements and were not included in the estimations.
4 Results
4.1 Autonomous navigation performance
4.1.1 Obstacle avoidance
The task of avoiding collisions during the flight is naturally more demanding in a dense forest than in a sparse forest. Consistent with these expectations, the trajectories were smoother and the success rate was better in ”Evo-medium” than in ”Evo-difficult” test area. In ”Evo-medium” all performed test flights were successful and five out of seven test flights were performed with smooth trajectories. In ”Evo-difficult” one of nine test flights failed and none of the test flights was performed with a smooth trajectory.
The reason for emergency stops in both environments was late detection of thin and dry spruce branches. In the failed test flight in ”Evo-difficult” the reason was also a collision to a similar thin low branch.
The test flight success rates in both test forests are summarized in the Table 3. Figure 7 presents examples of performed trajectories and obtained point clouds in all test forests.
Table 3: Test flight success in the test forests
Test areas Forest type Density (trees/ha) Flight lengths (m) Successful flights Smooth trajectories
”Evo-medium” Old Spruce forest 650 34 - 36 7/7 5/7
”Evo-difficult” Young Spruce forest 2000 42 8/9 0/8
(a) ”Evo-medium”
(b) ”Evo-difficult”
Figure 7: Examples of drones own navigation grid maps with performed trajectories (left) and obtained point clouds (right) in (a) ”Evo-medium” and (b) ”Evo-difficult”
4.1.2 VIO
The trajectory estimate error against the reference trajectory was successfully calculated for five flights in the medium forest and for three flights in the difficult forest. In the difficult forest, one of the flights was performed without coded targets making the reference trajectory non-comparable to others, in one flight there was a gap in the recorded onboard data due to a data recording error during a flight, and in one flight Metashape failed to retrieve the reference path. Average $\text{ATE}_{\text{pos}}$ was 0.50 m in the medium forest and 0.34 m in the difficult forest. The standard deviation between $\text{ATE}_{\text{pos}}$ was 0.06 m in the medium forest and 0.07 m in the difficult forest.
The VINS-Fusion estimates with online $t_{d}$ estimation were calculated afterwards for the same flights. With online $t_{d}$ estimation, average $\text{ATE}_{\text{pos}}$ was 0.47 m in the medium forest and 0.33 m in the difficult forest. The standard deviation between $\text{ATE}_{\text{pos}}$ was 0.02 m in the medium forest and 0.04 m in the difficult forest.
As the small standard deviation between the $\text{ATE}_{\text{pos}}$ estimates indicates, the error between the VINS-Fusion estimates and the reference trajectories was similar in all flights. In every flight, VINS-Fusion was estimating the shape of the trajectory well but underestimating the flight distance. Figure 8 presents an example of VINS-Fusion trajectory estimate and the reference trajectory in the difficult test forest. Table 4 presents the VINS-Fusion estimate errors in all flights.
Table 4: VINS-Fusion estimation Errors in all test flights
Test area Flight No. $\text{ATE}_{\text{pos}}$ (constant $t_{d}$) $\text{ATE}_{\text{pos}}$ (online estimated $t_{d}$)
”Evo-medium” 1 0.5567 0.4816
”Evo-medium” 2 0.5599 0.4790
”Evo-medium” 3 0.4538 0.4769
”Evo-medium” 4 0.4945 0.4722
”Evo-medium” 5 0.4313 0.4400
”Evo-difficult” 1 0.3787 0.3340
”Evo-difficult” 2 0.3853 0.3727
”Evo-difficult” 3 0.2680 0.2962
Figure 9 shows an example of VINS-Fusion tracking image during a test flight in the dense test forest. The algorithm detects stable features to track from all parts of the picture. The image shows also flying dry leaves blown up by the air flow from the propellers of the drone. The moving leaves did not cause visible disturbances to the tracking of the pose.
(a) Top view
(b) Side view
Figure 8: Example of a VINS-Fusion trajectory estimate and a reference trajectory obtained with Agisoft Metashape in the difficult forest from a) top and b) side.
Figure 9: Example of VINS-Fusion tracking image visualizing the tracked features. The tracking was not disturbed by the dry leaves blown by the air flow from the propellers of the drone. The coloring of the feature markers depends on the time how long the feature has been tracked. Blue dots symbolize new features and red features have been tracked longer.
In the loop closure tests, VINS-fusion detected a loop closure between 20 image pairs in the shorter path and four image pairs in the longer path with the same feature tracking parameters (Table 2) that were used in the flight tests. Figure 10 shows examples of detected loop closure image pairs from both paths. The number of detected loops can be increased by increasing the amount of features per image in the VINS-Fusion configuration. By reducing the minimum distance between features from 40 to 35 pixels, VINS-Fusion detected a loop closure between 52 image pairs in the shorter path and 23 image pairs in the longer path. By decreasing the minimum distance between features to 30 pixels, VINS-Fusion detected a loop closure between 82 image pairs in the shorter path and 63 image pairs in the longer path. However, increasing the number of features also increases the computational burden of the VIO, and the effects on the real-time performance during flying were not tested in this study.
(a) Shorter path.
(b) Longer path.
Figure 10: Example of image pairs where VINS-Fusion detects a scene to already visited one and performs a loop closure.
4.2 Forest reconstruction and analysis
4.2.1 Photogrammetric processing
Results of the photogrammetric processing are given in Table 5. In the image alignment process, all images were aligned in each dataset, which indicated successful processing. Re-projection errors varied between 0.29 and 0.39 pixels. The estimated flight GSDs (Ground Sampling Distance) ranged between 4.08 and 6.17 mm at ground level. Distances from the trees varied between 0.55 and 7.94 m, resulting in GSDs of 0.85 - 12.34 mm at the objects of interest. Point clouds from datasets II-IV had similar point densities of 2.63-2.84 points/cm2, while dataset I provided a slightly higher point density of 3.58 points/cm2.
Results of accuracy estimation using Check Points (CPs) are shown in Table 6. The cases without GCPs or with one GCP at the beginning of the trajectory provided similar 2D and 3D RMSEs, of 11.36-11.72 cm and 11.37-11.73 cm, respectively. When two GCPs were used, the 2D and 3D RMSE improved to 5.26 cm and 5.48 cm, respectively. Adding an additional GCP resulted in a significant improvement in the 2D and 3D RMSE, reducing them to 1.41 cm and 1.48 cm, respectively.
Table 5: Photogrammetric processing results. Img. count: Number images/Number of aligned images; Number of GCPs; Completeness % (Number of trees found/Number of total trees); Reprojection error (pix); Number of tie points (in millions); Point density (points(cm2) in ”Evo-medium” test site. I: Trajectory 1, II: Trajectory 2, III: Trajectory 3, IV: Trajectory 4, Combined: Trajectories 1-4 merged, and for Trajectory 1 I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
Dataset Img. count N GCPs Completeness (%) Reproj. error (pix) Num. tie points (millions) Point density (points/cm2)
I 650/650 15 79.31 (23/29) 0.39 197132 3.58
II 648/648 15 62.07 (18/29) 0.33 151889 2.84
III 619/619 11 61.29 (19/31) 0.33 151083 2.63
IV 597/597 15 62.50 (20/32) 0.33 141716 2.66
Combined 2514/2514 15 33.33 (12/36) 0.29 1547828 8.53
I0GCP 650/650 0 72.41 (21/29) 0.39 197132 3.54
I1GCP 650/650 1 68.97 (20/29) 0.39 197132 3.54
I2GCP 650/650 2 75.86 (22/29) 0.39 197132 3.54
I3GCP 650/650 3 79.31 (23/29) 0.39 197132 3.54
Table 6: Accuracy estimation (Root Mean Square Error; RMSE) using Check Points (CPs) and Camera Location in ”Evo-medium” test site, for trajectory 1: I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
Dataset Check Points (CPs) RMSE (cm) Camera location RMSE (cm)
N CPs X Y Z 2D 3D X0 Y0 Z0 2D0 3D0
I0GCP 15 6.99 8.95 0.65 11.36 11.37 7.06 9.07 0.30 11.49 11.50
I1GCP 14 7.19 9.25 0.56 11.72 11.73 7.03 9.06 0.28 11.47 11.47
I2GCP 13 5.01 1.61 1.52 5.26 5.48 7.03 9.06 0.28 11.47 11.47
I3GCP 12 1.19 0.75 0.45 1.41 1.48 1.46 0.70 0.56 1.62 1.71
4.2.2 Tree detection
Figure 12 shows post-processed trajectories for dataset I-IV, as well as reference and detected trees, and the boundaries of the point clouds. I and II followed almost the same trajectory, while III and IV had similar trajectories. Dataset I trajectory was smoothest and it did not include any sharp turns. Other 3 trajectories had one or more sharper turns. For datasets I, II, III and IV, the total number of reference trees inside the point cloud boundary was 29, 29, 31 and 32, of which 23, 18, 19 and 20 were found from drone-collected point cloud. The resulting completeness values were 79.31%, 62.07%, 61.29% and 62.5%, respectively. The total number of different trees found was 29 out of 33 (87,88%) from dataset I-IV. However, when all the flights were processed simultaneously (combined) that resulted only $3$ completeness, where only 12 of 36 trees was found. When utilizing 0, 1, 2 and 3 gcps in dataset I completeness values were 72.41 %, 68.97 %, 75.86 % and 79.31 %, respectively (Table 5).
Figure 11: GCPs, stereo images and reference trees visualized during the point cloud processing in the Agisoft Metashape (Metashape Development Team, 2021). Detected reference trees are marked, while those not detected are highlighted with a orange box.
In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the drone’s trajectory, despite these trees being well-reconstructed in the point cloud. Similarly, number of well-reconstructed trees but not detected by algorithm was 4, 5 and 5, in dataset II, III and IV, respectively. The three extra trees that were not identified in dataset I, were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud. The datasets II trough IV each contained 7 poorly reconstructed trees. Many of these trees were located at a greater distance from the trajectory, while others were obscured from the camera’s view by intervening trees.
4.2.3 DBH estimation
RMSE and bias of the DBH estimates of datasets I-IV were similar, ranging from 3.33 to 3.86 cm (10.69-12.98%) and -0.67 to 1.40 (-2.11-4.71%), respectively (Table 7). Processing all the flights simultaneously (Combined) provided only marginally better results with an RMSE of 2.91 cm (9.14%) and a bias of 1.51 cm (4.74%).
When considering the impact of the number of GCPs, the utilization of 3, 2, 1 and 0 GCPs in photogrammetric processing yielded slightly poorer results compared to those obtained with all GCPs. The RMSE and bias values ranged from 3.93 to 4.37 cm (13.68%-14.66%) and 1.12 to 2.13 cm (3.89-7.55%), respectively (Table 7).
Tree size had an impact on the DBH estimation accuracy (Table 7). For small and medium-sized trees, with a DBH smaller than 30 cm, DBH estimations achieved superior performance in all dataset, with RMSE and bias varying from 1.16 to 2.75 cm (5.74-12.47%) and -0.01 to 0.62 cm (-0.06-3.14%), respectively. Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99-25.44%) and -0.66 to 2.08 cm (-1.53-5.65%), respectively. This is demonstrated in Figure 14, illustrating a positive correlation between tree diameter and DBH estimation error. Notably, when the tree diameter exceeded 30 cm, the DBH estimation errors became more pronounced.
Analysis of the impact of the camera distance from trees did not reveal significant correlation between tree distance and DBH estimation errors (Figure 13).
Figure 12: Post-processed trajectories, reference trees and found trees in Dataset (a) I (b) II (c) III (d) IV..
Figure 13: Scatter plots of DBH errors with respect to distance to trees. Dataset (a) I (b) II (c) III (d) IV.
Figure 14: Scatter plots of DBH errors with respect to diameter of trees. Dataset (a) I (b) II (c) III (d) IV
Table 7: Accuracy of the stem diameter breast height (DBH) in ”Evo-medium” test site. Dataset is split in 3 categories: All, DBH $<$ 30 cm and DBH $>$ 30 cm trees. Datasets: I: Flight 1, II: Flight 2, III: Flight 3, IV: Flight 4, Combined: Flights I-IV merged, and for Dataset I: I0GCP: with no GCPs, I1GCP: with 1 GCP, I2GCP: with 2 GCPs, I3GCP with 3 GCPs.
Dataset Number of trees All $<$30 cm $>$30 cm RMSE (cm) All $<$30 cm $>$30 cm RMSEAll $<$30 cm $>$30 cm Bias (cm) All $<$30 cm $>$30 cm BiasAll $<$30 cm $>$30 cm Standard Deviation All $<$30 cm $>$30 cm
I 23 14 9 3.86 2.56 5.29 12.98 12.47 11.98 1.40 0.17 1.23 4.71 0.85 2.78 3.60 2.54 4.26
II 17 12 5 3.33 2.16 5.14 12.79 11.36 12.01 1.01 0.31 0.70 3.87 1.63 1.63 3.17 2.12 4.56
III 19 9 10 3.97 1.16 5.36 12.05 5.74 12.07 0.97 0.13 0.84 2.95 0.64 1.89 3.85 1.13 5.12
IV 20 10 10 3.40 1.28 4.63 10.69 6.21 10.77 -0.67 -0.01 -0.66 -2.11 -0.06 -1.53 3.33 1.28 4.44
Combined 12 7 6 2.91 1.41 4.04 9.14 6.70 8.99 1.51 -0.12 1.63 4.74 -0.56 3.62 2.49 1.39 1.87
I0GCP 21 12 9 4.37 2.75 7.93 14.66 13.07 18.62 1.61 -0.20 2.41 5.42 -0.97 5.65 4.06 2.73 5.17
I1GCP 20 12 8 4.38 2.02 6.46 14.71 9.63 15.06 1.68 -0.01 1.69 5.65 -0.06 3.95 4.04 2.02 4.88
I2GCP 23 14 8 4.35 1.77 11.08 15.43 8.98 25.44 2.13 0.62 2.08 7.55 3.14 4.77 3.79 1.48 8.96
I3GCP 23 14 9 3.93 1.22 6.10 13.68 6.23 14.20 1.12 0.14 0.98 3.89 0.72 2.27 3.77 1.20 5.57
5 Discussion and Conclusions
5.1 Autonomous flying
The performance of autonomous navigation was promising in both test forests. The obstacle avoidance success rate was 100% in the medium forest and 87.5% in the difficult forest. However, the smoothness of the trajectories suffered from the increasing forest density. While in the medium forest, 71.4% of the flights had a smooth trajectory, in the difficult forest at least one emergency stop was needed in all of the flights. Further consideration should be given to finding the optimal sensor for detecting thin and dry branches earlier. Nonetheless, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium difficulty level boreal forests.
The accuracy of the position estimate by VINS-Fusion was satisfying for short flights performed in this study, but in the longer flights, the error would start to accumulate making the position errors high. However, the standard deviation between the errors between the test flights was small, and the VINS-Fusion systematically estimated the flight paths shorter than in reality (as can be seen from Figure 8). This suggests, that a notable part of the position error could be eliminated by improving the calibration of the sensor setup. The results suggest that the estimates by VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration. However, further development of the prototype should include implementing hardware-level synchronization between the camera and the IMU of the autopilot.
The position error can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests showed that VINS-Fusion could detect loop closures even in a homogeneous forest environment. Increasing the number of features tracked by VINS-Fusion increases the number of detected loops. However, increasing the number of features increases also the computational burden of the system potentially disrupting the real-time estimation. Finding the optimal value for the tracked features as well as measuring the actual improvement of the loop closures to the position estimate requires further studies.
5.2 Accuracy of forest measurements
The proposed system was assessed by comparing it to other relevant studies (Section 2.2, Table 1) considering the accuracy, system autonomy and weight, and complexity of test environment.
The highest accuracy UAV SfM-MVS results were presented by Shimabuku, Konoshima, and Ota (2023) with an RMSE of 0.4 cm (3.2%). However, their test site included only 7 reference trees. Kuželka and Surovỳ (2018) and Krisanski, Taskhiri, and Turner (2020) had more reference trees and achieved similar DBH accuracy compared to that in this study, but the flights were piloted manually. Notably, the SfM-MVS approaches involved light-weight systems.
The fully automated under-canopy UAV Lidar system (ULS) by Liang et al. (2024) weighted 5 kg and provided an RMSE of 4.48 cm (21.37%) and bias of 5.01 (14.91%) at best. Wang et al. (2021) combined data from above- and under-canopy drone with a Riegl miniVUX-1UAV lidar (Riegl GmbH, Horn, Austria), resulting in an RMSE of 7.95 cm (33.35%) and bias of 3.72 cm (15.61%). Their under-canopy flight was carried out manually by a pilot and the system weighted approximately 10 kg. Muhojoki et al. (2024) carried out manually flown flights in three sparse density forest plots with the Deep Forest drone utilized an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA) and the Hovermap drone equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA), resulting in an averaged RMSE of 1.5 cm (6.7 %) and 1.6 cm (6.9 %) and bias of -0.7 cm (3.0 %) and -0.1 cm (-0.6 %). However, without utilizing their bias removal method RMSE and bias were significantly higher. The highest accuracy DBH estimates with under-canopy LiDAR were presented by Hyyppä et al. (2020a) resulting in an RMSE of 0.6 cm (2.3%) and bias of 0.34 cm (1.3%) in a sparse forest and 1.1 cm (3.5%) and 0.12 cm (0.4%) in a medium density forest. Their system was also manually piloted and weighted approximately 10 kg, a Tarot 960 UAV equipped with Kaarta Stencil-1 (VLP-16) LiDAR (Kaarta, Pittsburgh, Pennsylvania, USA).
Summarizing, the highest accuracy levels were obtained with the relatively heavy ULS systems, which were operated manually in sparse forest environments. Miniaturized lightweight sensors were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), and Kuželka and Surovỳ (2018), and Liang et al. (2024). Results from the SfM-MVS systems were comparable to each other but ULS provided slightly poorer accuracy. Only studies by Shimabuku, Konoshima, and Ota (2023) and Liang et al. (2024) had flights with a high level of automation. Considering the complexity of the environments, the forest by Liang et al. (2024) (1000 trees/ha) was the closest to that in this study (650 and 2000 trees/ha), while other studies were conducted in remarkably sparser forests.
Based on the previous, it can be concluded that when considering the combination of accuracy, autonomy of flying, and environmental complexity, the proposed system introduced a new level of performance for under-canopy forest reconstruction. The results show the great promise of miniaturized camera SfM-MVS in autonomous under-canopy forest reconstruction.
5.3 Potential improvements and future work
During the evaluation of the system, some areas for improvement were identified, which can help develop the method towards a tool in practical forest measurements.
The reliability of the navigation could be improved by performing further research on potential sensors and methods for detecting the thin branches. Blindsides in the mapping could be tackled either by sensors with 360° FOV (e.g. Lidar, 360° camera or multi-camera system) or storing historical data to occupancy grid maps in addition to the currently visible area.
The virtual floor and ceiling of Ego-planner-V2 (Zhou et al. 2022) are also defined as static with respect to the initial altitude of the drone limiting the low-altitude flying to relatively flat areas. To comprehensively map uneven forest plots, the altitude limits should be modified to be dynamic and with respect to the local terrain level.
The drift compensation is one of the important future development goals. Loop closure tests provided promising results and potential future tests could evaluate the approach by performing block flights. In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate. The potential methods for GNSS-Fusion include using the global fusion features in VINS-Fusion (Qin et al. 2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g. GVINS (Cao, Lu, and Shen 2022). Even though the dense forest canopy blocks the GNSS signals, during long-term operations the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forests.
It is also notable that the rigorous photogrammetric post-processing effectively compensated for the systematic errors. While the real-time trajectory error was 50 cm, the post-processed 3D error was approximately 11 cm without any GCPs, this correlates with the trajectory error. Therefore, for applications relying on post-processing, drift is not expected to pose significant issues, especially when covering small areas such as forest test plots.
In this study, the performed flight paths were straight flights through the sample plots, and the DBH estimation was performed only for the tree trunks recorded during that path. For more comprehensive forest characterization, the trajectory should cover the whole test plot, and observe the trees from multiple sides. The global path covering the whole test plot could be implemented either with a simple waypoint generation algorithm as Liang et al. (2024) or with more advanced exploration methods such as proposed by Zhou et al. (2021a) or Cieslewski, Kaufmann, and Scaramuzza (2017). Another way to achieve area coverage would be utilizing a swarm of drones as in the original article by Zhou et al. (2022). Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage.
While this study only demonstrated the potential of forest characteristics estimation through DBH estimation and tree detection, the scope of data collection is not limited to that. Having a second camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds. With a developed camera-based prototype, the data collection is not limited to point clouds, but also applications of image based mapping and object detection can be developed on top of it. Potential applications could include e.g. classifying and counting species of trees and other vegetation as well as monitoring damages in the bark of the trees due to insect pests.
The prototype opens also possibilities for research of real-time forest characteristics analysis. In this study, the point clouds were obtained with post-processing, but the utilized onboard camera Intel Realsense D435 can also produce 3D point clouds in real-time. Image-based object detection can also be performed in real-time e.g. with YOLO. The analysis can be performed either with the onboard computer or with a remote computer or cloud service over the Internet connection.
6 Acknowledgements
This research was funded by the Research Council of Finland within projects “Learning techniques for autonomous drone based hyperspectral analysis of forest vegetation” (decision no. 357380), Fireman (decision no. 346710), and the Research Council of Finland Flagship Forest–Human–Machine Interplay—Building Resilience, Redefining Value Networks and Enabling Meaningful Experiences (UNITE) (decision no. 357908).

### v2

@@
 Abstract
-Drones are increasingly used in forestry to capture high-resolution remote sensing data.
+Drones are increasingly used in forestry to capture high-resolution remote sensing data, supporting enhanced monitoring, assessment, and decision-making processes.
 While operations above the forest canopy are already highly automated, flying inside forests remains challenging, primarily relying on manual piloting.
@@
 In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open-source methods and validating its performance for data collection inside forests.
-The autonomous flight capability was evaluated through multiple test flights in two boreal forest test sites.
-The tree parameter estimation capability was studied by conducting diameter at breast height (DBH) estimation using onboard stereo camera data and photogrammetric methods.
-The prototype conducted flights in selected challenging forest environments, and the experiments showed excellent performance in forest reconstruction with a miniaturized stereoscopic photogrammetric system.
-The stem detection algorithm managed to identify 79.31% of the stems.
-The DBH estimation had a root mean square error (RMSE) of 3.33 cm (12.79%) and a bias of 1.01 cm (3.87%) across all trees.
-For trees with a DBH less than 30 cm, the RMSE was 1.16 cm (5.74%), and the bias was 0.13 cm (0.64%).
-When considering the overall performance in terms of DBH accuracy, autonomy, and forest complexity, the proposed approach was superior compared to methods proposed in the scientific literature.
-Results provided valuable insights into autonomous forest reconstruction using drones, and several further development topics were proposed.
+Specifically, the study focused on camera-based autonomous flight under the forest canopy and photogrammetric post-processing of the data collected with the low-cost onboard stereo camera.
+The autonomous flight capability of the prototype was evaluated through multiple test flights at boreal forests.
+The tree parameter estimation capability was studied by performing diameter at breast height (DBH) estimation.
+The prototype successfully carried out flights in selected challenging forest environments, and the experiments showed excellent performance in forest 3D modeling with a miniaturized stereoscopic photogrammetric system.
+The DBH estimation achieved a root mean square error (RMSE) of 3.33 cm (12.79 %) across all trees.
+For trees with a DBH less than 30 cm, the RMSE was 1.16 cm (5.74 %).
+The results provide valuable insights into autonomous under-canopy forest mapping and highlight the critical next steps for advancing lightweight robotic drone systems for mapping complex forest environments.
 keywords:
 Robotic; drone; forest; photogrammetry; DBH
-articletype: ARTICLE PREPRINT
+articletype: ARTICLE PREPRINT - Word Count 10048
 1 Introduction
-Over the last decade, the use of drones in forestry has rapidly increased, offering a practical solution for forest data collection both above and inside the forest canopy (Puliti et al.
-2015; Hyyppä et al.
-2020a; Liang et al.
-2019).
-Modern drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems.
-However, their positioning systems rely exclusively on Global Navigation Satellite Systems (GNSS) signals, limiting autonomous flights to open areas where these signals are reliable.
-In dense forests, GNSS localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al.
-2010).
-Additionally, navigating inside forests requires the capability to autonomously avoid obstacles such as trees and bushes along the flight path.
-Concequently, under-canopy data collection with drones is still primarily controlled manually by human pilots (e.g.
-Hyyppä et al.
-2020a; Liang et al.
-2019; Krisanski, Taskhiri, and Turner 2020).
-Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in GNSS-denied, cluttered environments.
-In recent years, several solutions utilizing various sensor setups and algorithms for autonomous navigation under forest canopies have been proposed (e.g.
-Liu et al.
-2022; Loquercio et al.
-2021; Zhou et al.
-2022).
-However, these methods are typically validated by reporting a single test flight in a specific forest environment.
-Most studies assess the effects of varying object density only in simulators, with only a couple of the methods tested through multiple test flights inside real forests (Loquercio et al.
-2021; Ren et al.
-2022; Liu, Ren, and Zhang 2024).
-Therefore, a significant research gap remains regarding the applicability and reliablity of these algorithms in diverse forest environments.
-The autonomous under-canopy drone solutions proposed in the literature typically focus on aerial robotics and primarily utilize the collected sensor data for navigation.
-However, recent research has begun to explore boarder applications.
-For example, Liu et al.
-(2022) employed a semantic simultaneous localization and mapping (SLAM) algorithm to calculate the number of trees in the explored area.
-Similarly, Liang et al.
-(2024) demonstrated the use of autonomous drones for in situ forest observations under the canopy.
-While their work estimated tree attributes based on data collected during two test flights, it did not evaluate the reliability of autonomous flight itself.
-Consequently, a research gap remains regarding the borader application of autonomous drones for forestry applications.
-Remote sensing measurements within forests are increasingly used to capture detailed forest information.
+Over the last decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information.
 Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al.
-2007) and mobile solutions e.g.
-handheld devices (Ryding et al.
+2007) and mobile solutions, e.g., handheld devices (Ryding et al.
 2015; Bauwens et al.
@@
 2018; Bienert et al.
-2018) or manually operated drones (Chisholm et al.
+2018) or drones (Chisholm et al.
 2013; Brede et al.
@@
 2020a).
-Similarly, the use of structure-from-motion (SFM) based point clouds for forest mapping has been demonstrated using different camera solutions (Liang et al.
+Similarly, the use of structure-from-motion (SFM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., Liang et al.
 2014a; Wallace et al.
@@
 2019).
-Combining these techniques with autonomously navigating drones inside forests represents a highly relevant research opportunity.
-To address these gaps and opportunities, this study introduces a novel approach for autonomous data collection and object characterization under unstructured forest canopies.
-The primary objectives were to assess the performance of an autonomously flying drone in a boreal forest and to demonstrate its capability in measuring tree characteristics.
-A prototype autonomously flying drone was developed using state-of-the-art open-source algorithms, followed by multiple flight tests to evaluate its performance and suitability for boreal forests.
-Image data collected by the low-cost onboard stereo-camera was stored for further analysis.
-The dataset was post-processed using SfM photogrammetry to improve the georeferencing and to generate dense 3D point clouds.
-Finally, tree detection and diameter at breast height (DBH) estimation were performed on the point clouds, assessing the potential of the low-cost autonomous system for forest parameter estimation.
-The study is structured as follows.
-Section 2 reviews recent open-source solutions for autonomous drone flights inside forests and state-of-the-art methods forest parameter estimation using drones flying inside forests.
-Section 3 describes the algorithms, hardware, experiments, and methods for autonomous navigation, tree detection, and DBH estimation.
+Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
+Hence, the use of drones in forestry research has rapidly increased, and they offer a practical solution for forest data collection above the forest canopy (Puliti et al.
+2015; Tuominen et al.
+2017).
+Modern commercial drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems.
+However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable.
+In dense forests, GNSS localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al.
+2010).
+Additionally, navigating inside forests requires the capability to autonomously avoid obstacles such as trees and bushes along the flight path.
+Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., Hyyppä et al.
+2020a; Liang et al.
+2019; Krisanski, Taskhiri, and Turner 2020; Prabhu et al.
+2024).
+Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in GNSS-denied, cluttered environments.
+Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., Liu et al.
+2022; Loquercio et al.
+2021; Zhou et al.
+2022).
+However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment.
+Most studies assess the effects of varying object density only in simulators, with only a couple of the methods tested through multiple test flights inside real forests.
+Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation.
+Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (Liu et al.
+2022; Liang et al.
+2024; Cheng et al.
+2024).
+Notably, all of these studies have utilized Ouster OS1 3D LiDAR (Ouster, San Francisco, CA, USA) as the main sensor.
+Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (Reddy Cenkeramaddi et al.
+2020).
+In the case of drone development, where optimizing the weight and battery consumption play a key role, these factors lead to bigger and heavier drone platforms and flight batteries requiring more space between the obstacles to fly safely inside the forest.
+Moreover, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost.
+The data collected with the LiDARs is also limited to point clouds only, whereas the images collected with cameras can also be used for image-based analysis.
+Hence, although some LiDAR-based robotic drone systems have been successfully applied to under-canopy mapping of forest parameters, the potential of smaller and cheaper camera-based systems remains largely unexplored.
+Furthermore, the robotic drone systems proposed in the literature are typically not evaluated through multiple flight tests in different forest environments leaving the reliability and suitability to diverse forest environments as an open question.
+Therefore, in this study, a prototype of a small camera-based robotic drone system that can fly autonomously in challenging GNSS-denied forests was built and developed using custom hardware and state-of-the-art robotic methods.
+The study aimed to tackle the research gaps in flight reliability of camera-based robotic drone systems in various boreal forest environments, as well as the potential of photogrammetric DBH estimation with autonomous drones in such environments.
+The performance and the reliability of autonomous flight capabilities were evaluated by performing multiple test flights in boreal forests with varying difficulty.
+In addition, the image data collected with the low-cost onboard stereo camera were processed into point clouds, and the DBH values were estimated for the trees detected from the dataset.
+To the best of our knowledge, this study is the first to empirically validate a miniaturized, autonomous, camera-based robotic drone for under-canopy forest reconstruction.
+Specifically, the study focused on three research questions.
+What level of autonomous flight reliability can be achieved in challenging boreal forests with current methods for camera-based robotic drone navigation?
+What level of DBH estimation accuracy can be achieved with the low-cost onboard stereo-camera data recorded during the flight?
+What are the remaining bottlenecks to be tackled in follow-up studies in order to achieve large-scale autonomous photogrammetric forest data collecting with miniaturized under-canopy robotic drone systems?
+The main contributions of this study can be highlighted as follows:
+  • Build and development of a lightweight robotic drone prototype capable of flying within GNSS-denied forest environments, using custom hardware and state-of-the-art robotic algorithms.
+The reliability of the system was comprehensively evaluated through multiple test flights in boreal forests with varying tree densities and difficulty levels.
+  • Combining autonomous flight capabilities to under-canopy photogrammetric forest reconstruction by processing the data from a low-cost onboard stereo-camera to 3D point clouds.
+  • Validation of autonomous flight data collection based on stereo cameras for the estimation of forest parameters through DBH estimation, including an evaluation of the impact of various factors on the accuracy of the estimate.
+  • Comprehensive discussion of current limitations to large-scale, fully autonomous photogrammetric under-canopy forest monitoring, along with proposed solutions and further research directions.
+The rest of this study is structured as follows.
+Section 2 reviews the latest studies that estimate DBH using drones flying inside forests and recent open-source solutions for autonomous drone flights inside forests.
+Section 3 describes the algorithms, hardware, experiments, and methods utilized in this study.
 Experimental results are presented in Section 4.
-Section 5 discusses the results, potential improvements, and concludes the study.
+Section 5 discusses the results, and potential improvements, and concludes the study.
 2 Related work
-2.1 Autonomous flying inside forests
-Autonomous navigation in obstacle-rich environments has been an actively researched topic in robotics during the last decades.
-Whereas the majority of the research has tried to tackle the 2D path planning problem for autonomous ground vehicles, in recent years increasing number of methods for autonomous path planning and obstacle avoidance have also been published for drones.
-Although many of the methods proposed in the literature are only tested in simulations or indoor conditions, some of the methods have been also successfully flight-tested inside forests.
-Some of the authors have also open-sourced their methods to encourage the development of the field.
-In this section, recent open-source methods successfully flight-tested inside a forest are summarized.
-Campos-Macías et al.
-(2021) proposed a solution that focuses on computational efficiency that can be run in real time with a low-power CPU.
-The solution used a stereo tracking camera to run Visual-Inertial Odometry (VIO) for pose estimation and a point cloud from an active stereo depth camera to create a 3D voxel map.
-Collision-free trajectories were planned to the free regions on the map.
-In addition to multiple indoor environments, the system was tested in a sparse mixed forest.
-However, the flight distance in the forest test was only 11.58 m.
-Liu et al.
-(2022) proposed a method that aims at large-scale autonomous flights in forest environments.
-The proposed system used an inertial measurement unit (IMU) and a stereo camera to run VIO and 3D light detection and ranging (LiDAR) to run a semantic simultaneous localization and mapping (SLAM) algorithm.
-Semantic SLAM was developed to correct drift in the VIO estimate.
-The system was reported to fly an 800 m long autonomous flight, but the semantic SLAM algorithm was only tested in manual flight.
-The documented flight tests were performed in a North American pine forest.
-The forest was relatively dense, but it had a minimal amount of low branches and understory vegetation.
-Ren et al.
-(2022) proposed another LiDAR-based method that generates sphere-shaped safe flight corridors from the point clouds.
-The method was targeting for high-speed flying in unknown environments.
-FAST-LIO2 (Xu et al.
-2022) was utilized in the pose estimation.
-In total, 12 test flights with varying flight speeds were performed in a sparse Chinese deciduous forest.
-The maximum flight speed was 13.7 m/s and 11 of 12 test flights were successful, but the trajectory lengths were not reported.
-Liu, Ren, and Zhang (2024) have also proposed another method for the same LiDAR-based platform targeting especially safe navigation in the presence of dynamic obstacles and external disturbances with an integrated planning and control framework.
-Planning was performed within a local grid map and FAST-LIO2 (Xu et al.
-2022) was utilized in the pose estimation.
-In addition to indoor environments, the system was also tested with 10 successful flights at varying speeds in the same forest as (Ren et al.
-2022).
-The maximum flight speed was 5.86 m/s, and the reported trajectory length was 58.64 m.
-Instead of using a single drone, the solution proposed by Ahmad et al.
-(2022) focused on the navigation of drone swarms.
-The proposed approach is fully decentralized and tries to mimic the collective navigation behavior of natural flocks and swarms.
-The physical drone hardware used in the tests incorporated a LiDAR for running the SLAM algorithm by Kohlbrecher et al.
-(2011), a laser rangefinder as an altimeter, and an Ultra Violet Direction And Ranging (UVDAR) system for relative localization between the drones of the swarm.
-The approach was tested in a swarm flight inside a forest, but the test forest was very clean containing no low branches or understory vegetation.
-The flight distance was approximately 40 m.
-Zhou et al.
-(2022) also proposed a solution concentrating on swarm flights.
-In the proposed solution the drones use VIO for localization and probabilistic maps for obstacle avoidance.
-Swarm members also share their trajectories over a wireless network and correct the VIO drift based on the relative distances between the drones measured with Ultra Wide Band (UWB) sensors.
-The proposed solution was tested with palm-sized drones which used only one stereo camera to run both mapping and VIO.
-In the flight tests, a swarm of drones was successfully flying approximately 80 m through a dense bamboo forest.
-The solution proposed by Loquercio et al.
-(2021) replaced different sensing, mapping, and planning subtasks with a single neural network directly mapping depth images and a pose estimate from VIO to a feasible collision-free trajectory for the drone.
-The system was trained entirely with simulated data, and the scope of the research was high-speed autonomous flying.
-Real-world flight tests were performed with a drone using an active stereo depth camera to produce the depth images, and a stereo tracking camera to run VIO.
-Flight tests were performed in various environments containing also Central-European mixed forests.
-The forest densities were not reported, but the forests were relatively sparse with a low amount of understory vegetation.
-The evaluated trajectories were a 40 m long straight line and a circle with a 6 m radius, and multiple test flights were conducted with flight speeds varying from 3 to 10 m/s.
-The flight success rate was 100 % with a flight speed of 3 m/s and 5 m/s, 80 % with a flight speed of 7 m/s, and 60 % with a flight speed of 10 m/s
-Nguyen et al.
-(2023) proposed also a deep-learning-based method that utilizes neural networks to directly map depth images and partial pose estimates of the drone to a series of actions leading to a collision-free trajectory.
-The proposed system was also trained entirely in a simulator.
-In addition to the indoor flight tests, the performance assessment was reported for a single 60 m long trajectory through a dense Finnish pinewood forest making the test environment the most similar to the target environments in this study.
 2.1 DBH estimation from under-canopy drone data
-In a forest inventory, DBH is one of the most important tree attributes to be estimated since it provides a measure for the size of a tree and also enables the prediction of the total stem volume (Laasasenaho 1982).
-During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH.
-Even though several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al.
-(2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al.
-(2018); Piermattei et al.
-(2019); Xu, Shen, and Cao (2023)) and above-canopy photogrammetric systems (e.g., Wallace et al.
-(2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023).
-If the operation of such under-canopy drone systems can be automated in the future, they would provide a truly efficient solution for the acquisition of accurate forest data.
-Kuželka and Surovỳ (2018) manually operated a drone system based on a DJI Phantom 4 Pro within two forest plots to produce low-noise photogrammetric point clouds, from which the DBH of individual trees could be estimated with an RMSE between 2.6 cm and 5.3 cm depending on the plot and the applied analysis method.
-The studied plots were sparse with only 270-290 trees/ha.
-Krisanski, Taskhiri, and Turner (2020) performed manual flights of a drone system inside two forest plots with a stem density of approximately 1000 trees/ha to acquire high-density SfM point cloud data, from which the DBH of individual trees could be measured with a RMSE of 1.1-2.1 cm.
-Shimabuku, Konoshima, and Ota (2023) achieved an RMSE of 0.4-0.7 cm for DBH estimation on two small forest plots using SfM point clouds acquired with an under-canopy flying drone system.
-The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
-They also demonstrated stem volume estimates of individual trees with a RMSE of 15.8-17.2%.
-In addition to photogrammetric approaches, LiDAR-based methods have been studied extensively for the purpose of efficiently and accurately estimating various tree attributes, such as DBH, tree height and stem volume.
+In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total stem volume either alone or combined with the three height (Laasasenaho 1982).
+Although some of the relevant studies have also estimated tree volume by flying the flight path multiple times with different altitudes, with rotating sensors, or with Lidars with high vertical FOV, only DBH estimation accuracies and methods are included in this brief literature review.
+In the existing literature, LiDAR-based methods have been studied extensively for the purpose of efficiently and accurately estimating various tree attributes, such as DBH, tree height, and stem volume.
 To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al.
@@
 2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation.
-Even though a LiDAR can be mounted to a variety of other mobile platforms besides a drone, the focus is on the use of under-canopy drones for DBH estimation in this brief literature review.
 Chisholm et al.
-(2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually-piloted drone flying under the forest canopy.
-They detected 12 trees in the vicinity of the drone trajectory and reported a RMSE of 25.1% for the DBH estimates of the detected trees.
+(2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy.
+The utilized custom-built drone system included Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan).
+They detected 12 trees in the vicinity of the drone trajectory and reported an RMSE of 25.1 % for the DBH estimates of the detected trees.
 Hyyppä et al.
-(2020a) collected high-quality point cloud data using a Kaarta Stencil-1 laser scanner mounted horizontally at the bottom of a manually-piloted drone.
-They demonstrated DBH estimates with a RMSE of 0.6-0.9 cm (2%-3%) on two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha.
+(2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone.
+They demonstrated DBH estimates with an RMSE of 0.69-0.92 cm (2.2 %-3.1 %) and a bias of 0.29-0.31 cm (1.0 %-1.1 %) on two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha.
 Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory.
-As the laser scanner was mounted horizontally in their study, the limited field of view made it difficult to obtain direct measurements of tree height.
-In a follow-up study by Hyyppä et al.
-(2021), this issue was resolved by instead integrating a rotating laser scanner to an under-canopy flying drone, which resulted in accurate measurements for both stem diameter and tree height.
+In a follow-up study, Hyyppä et al.
+(2020c) compared various mobile laser-scanning methods, including backpack, handheld, Under-Canopy UAV, and Above-Canopy UAV, for forest data collection.
+The same manually piloted under-canopy drone system was utilized in this study in the same forest plots as in Hyyppä et al.
+(2020a).
+In this study, the manually piloted under-canopy drone system yielded an RMSE of 0.6-1.1 cm (2.3 %-3.5 %) and a bias of 0.34-0.12 cm (1.3 %-0.4 %).
 Wang et al.
@@
 5 laser scanner (Ouster, San Francisco, CA, USA).
-Both drones were flown manually in Muhojoki et al.
-(2024) to collect data in sparse density forest plots.
-The RMSE and bias of DBH estimates for the Deep Forestry’s and Hovermap’s drones, with bias removal method introduced by Muhojoki et al.
+Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between the trees was at least 3 m.
+Both drone systems were used to collect point clouds under the canopy of three sparse forest plots.
+The RMSE and bias of DBH estimates for Deep Forestry and Hovermap drones, with bias removal method introduced by Muhojoki et al.
 (2024), averaged RMSE of 1.5 cm (6.7 %) and bias of -0.7 cm (3.0 %) and RMSE of 1.6 cm (6.9 %) and bias of -0.1 cm (-0.6 %), respectively.
-Without using bias removal method, RMSE and bias for both drones were significantly higher.
-Muhojoki et al.
-(2024) also demonstrated that the under-canopy drone systems can provide roughly as accurate tree attributes as other under-canopy MLS methods, such as handheld LiDARs.
-Recently, Liang et al.
+Without using the bias removal method, RMSE and bias for both drones were significantly higher.
+In addition to the manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation.
+Liang et al.
 (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha.
-To enable the autonomous flight, a predefined global path was planned, to which the drone system performed real-time adjustments during the flight according to real-time environment conditions.
-When the flight path covered most of the test plot, they achieved an RMSE of 5.1 cm for the DBH of individual trees, and a similar RMSE for the stem curve.
-In addition to under-canopy flying drones, LiDARs have also been integrated to above-canopy flying drones capable of direct stem diameter measurements, though typically with a significantly higher RMSE and a lower tree detection rate due to the larger measurement distance and the upper canopy obstructing the tree stems (Wieser et al.
+To enable autonomous flight, a predefined global path was planned, to which the drone system performed real-time adjustments during the flight according to real-time environmental conditions.
+When the flight path covered most of the test plot, they achieved an RMSE of 5.13 cm (22.01 %) and bias of 3.40 cm (14.62 %) for the DBH of individual trees.
+Liu et al.
+(2022) employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (Chen et al.
+2020) algorithm to calculate the number of trees in the explored area.
+Although in this study the autonomous flight was demonstrated without the SLOAM algorithm and the tree number calculation was only demonstrated with one manually-piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation.
+Prabhu et al.
+(2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data.
+The DBH values were estimated for 20 trees located in two different test plots with RMSE of 1.67 cm (7.42 %).
+However, the DBH data-collecting flights were performed with manual flight mode.
+Cheng et al.
+(2024) used the same drone platform to collect a dataset for LiDAR-based mapping of trees.
+In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas.
+Although the drone was manually piloted in one of these forests, the data collection at the other forest consisting of loblolly pine trees was performed in autonomous flight mode.
+They also estimated the DBH value for 77 trees from the LiDAR data collected during the autonomous flights achieving an RMSE of 3.8 cm (12.3 %) by offline afterprocessing and 9.8 cm (31.8 %) by online estimation with SLOAM.
+With manually piloted data collection in the second test forest consisting of 20 trees with varying species, the RMSEs were 2.0 cm (7.5 %) with offline processing and 7.5 cm (27.5 %) with online processing.
+Besides the under-canopy flying drones, LiDARs have also been integrated to above-canopy flying drones capable of direct stem diameter measurements, though typically with a significantly higher RMSE and a lower tree detection rate due to the larger measurement distance and the upper canopy obstructing the tree stems (Wieser et al.
 2017; Brede et al.
@@
 2022).
-A summary of relevant studies on under canopy forest inventory studies is presented in Table 1.
+During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH.
+Even though several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al.
+(2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al.
+(2018); Piermattei et al.
+(2019); Xu, Shen, and Cao (2023); Gao and Kan (2022)) and above-canopy photogrammetric systems (e.g., Wallace et al.
+(2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023; He et al.
+2025).
+Kuželka and Surovỳ (2018) manually operated a drone system based on a DJI Phantom 4 Pro (DJI, Shenzhen, China) within two forest plots to produce low-noise photogrammetric point clouds, from which the DBH of individual trees could be estimated with an RMSE between 2.6 cm and 5.3 cm depending on the plot and the applied analysis method.
+The studied plots were sparse with only 270-290 trees/ha.
+Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro inside two forest plots with a stem density of approximately 1000 trees/ha to acquire high-density SfM point cloud data, from which the DBH of individual trees could be measured with an RMSE of 1.1-2.1 cm.
+Shimabuku, Konoshima, and Ota (2023) achieved an RMSE of 0.4-0.7 cm for DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China).
+The first plot containing 18 trees was clean containing a minimal amount of low-hanging branches and understory vegetation.
+The second test plot was denser and closer to a natural subtropical forest environment but contained only 7 trees.
+The drone system was able to detect obstacles in the horizontal direction but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
+Recently, He et al.
+(2025) mounted an additional 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) to the same LiDAR-based autonomous drone platform used in Liu et al.
+(2022) and post-processed the photogrammetric data to 3D point clouds.
+Whether the flights were performed in autonomous flight mode or with manually piloted mode was not mentioned in the study.
+Their study evaluated DBH for 43 trees with varying tree species and the reported median relative DBH error was 11.5 % estimated from the photogrammetric point clouds.
+A summary of the results obtained in the relevant studies on under-canopy forest inventory is presented in Table 1.
+In summary of the related work, most of the studies for under-canopy drone-based DBH estimation have been performed with manually piloted drones.
+Although a few studies have explored the potential of custom robotic drone systems capable of flying without commands from the pilot during the flight, all of these systems have been based on LiDARs.
+All of the studies based on photogrammetry and SfM point clouds have utilized commercial drone systems, which have required manual piloting during the data collection.
 Table 1: Accuracy of the stem diameter breast height (DBH) estimations in other studies.
 SfM-MVS: Structure-from-motion and multi-view stereo; ULS: Under canopy laser scanning; ALS: Above canopy laser scanning.
-*averaged values from three plots
-Dataset Type Number of trees RMSE (cm) RMSEBias (cm) Bias%
-Shimabuku et al. (2023) plot I Drone SfM-MVS 7 0.4 3.2 - -
-Shimabuku et al. (2023) plot II Drone SfM-MVS 18 0.7 3.6 - -
+*averaged values from three plots.
+The study by He et al.
+(2025) was omitted from this table since they reported only median errors instead of RMSE.
+Dataset Type Number of trees RMSE (cm) RMSE % Bias (cm) Bias %
+Chisholm et al. (2013) ULS 12 - 25.1 - -
+Liang et al. (2024) traj. I ULS 143 9.79 61.41 7.18 45.03
+Liang et al. (2024) traj. II ULS 143 5.13 22.01 3.40 14.62
+Hyyppä et al. (2020a) plot I ULS 42 0.69 2.2 0.31 1.1
+Hyyppä et al. (2020a) plot II ULS 43 0.92 3.1 0.29 1.0
+Hyyppä et al. (2020c) plot I ULS 42 0.6 2.3 0.34 1.3
+Hyyppä et al. (2020c) plot II ULS 43 1.1 3.5 0.12 0.4
+Wang et al. (2021) ULS&ALS - 7.95 33.35 3.72 15.61
+Muhojoki et al. (2024) * ULS 50 1.50-1.60 6.70-6.90 0.10-0.70 0.60-3.00
+Prabhu et al. (2024) ULS 20 1.67 7.42 - -
+Cheng et al. (2024) Plot I ULS 77 3.8 12.8 - -
+Cheng et al. (2024) Plot II ULS 20 2.0 7.5 - -
+Shimabuku et al. (2023) plot I Drone SfM-MVS 18 0.7 3.6 - -
+Shimabuku et al. (2023) plot II Drone SfM-MVS 7 0.4 3.2 - -
 Kuželka and Surovỳ (2018) Drone SfM-MVS 119 2.63-5.31 7.01-13.3 0.27-3.77 0.69-10.11
 Krisanski et al. (2020) Drone SfM-MVS 29 4.1 - - -
-Liang et al. (2024) traj. I ULS 143 9.43 63.41 7.17 48.21
-Liang et al. (2024) traj. II ULS 143 4.48 21.37 5.01 14.91
-Hyyppä et al. (2020a) plot I ULS 42 0.6 2.3 0.34 1.3
-Hyyppä et al. (2020a) plot II ULS 43 1.1 3.5 0.12 0.4
-Wang et al. (2021) ULS&ALS - 7.95 33.35 3.72 15.61
-Muhojoki et al. (2024) * ULS 50 1.50-1.60 6.70-6.90 0.10-0.70 0.60-3.00
+The highest estimation accuracies have been obtained with large systems unable to fly in dense forests or with test plots containing very few reference trees, but miniaturized systems have still yielded promising results.
+Notably, all studies utilizing SfM point clouds have been performed in environments that differ significantly from boreal forests.
+2.2 Autonomous flying inside forests
+Even though only a few studies have yet utilized robotic under-canopy drones in DBH estimation, drone autonomy itself has been an actively researched topic in the robotics community during the past few years.
+Recent advances in mobile robotics and the enhanced computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in GNSS-denied cluttered environments.
+Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments.
+Most importantly, the authors of these methods have also shared their source code to be openly available for the research community to encourage more development in the field.
+While the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al.
+2022; Prabhu et al.
+2024; Ren et al.
+2022; Liu, Ren, and Zhang 2024; Ahmad et al.
+2022) and camera-based methods (Campos-Macías et al.
+2021; Zhou et al.
+2022; Loquercio et al.
+2021; Nguyen et al.
+2023) based on the sensor utilized for the sensing of the obstacles in the flight path.
+Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system.
+Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by Loquercio et al.
+(2021) and Nguyen et al.
+(2023) replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories.
+In addition, the methods proposed by Zhou et al.
+(2022) and Ahmad et al.
+(2022) concentrated a swarm of drones instead of autonomous flying of an individual drone.
+However, most of these methods are typically validated by a single successful autonomous test flight in a specific forest environment.
+Only four of the articles reported multiple test flights inside real forests (Loquercio et al.
+2021; Ren et al.
+2022; Liu, Ren, and Zhang 2024; Prabhu et al.
+2024).
+From these, (Loquercio et al.
+2021) and (Prabhu et al.
+2024) were the only methods tested at least in two different forest types.
+Moreover, since methods by (Loquercio et al.
+2021), (Ren et al.
+2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in those studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation.
+The autonomous trajectory planning method by Prabhu et al.
+(2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data collecting flights in denser forests were manually piloted.
+Therefore, open questions remain regarding the applicability and reliability of these algorithms in diverse forest environments, especially in challenging boreal forests.
 3 Materials and methods
 3.1 Autonomous navigation algorithms
-In this study, the objective was to develop a prototype drone capable of autonomously avoiding obstacles under the forest canopy and operating without GNSS.
-The drone was required to autonomously fly from the takeoff location to a specified goal point while recording sensor data during the flight.
+In this study, the objective was to develop a prototype lightweight camera-based drone capable of autonomously avoiding obstacles under the forest canopy and operating without GNSS.
+The drone was required to fly autonomously from the takeoff location to a specified goal point while recording stereo-camera data during the flight.
 This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
@@
 (2022) was chosen as the base of the prototype.
-The solution was proven to work with only one stereo camera making it possible to build a smaller and lighter platform.
-Furthermore, cameras are typically lighter and more power-efficient than LiDARs (Reddy Cenkeramaddi et al.
-2020) making them an attractive solution for small under-canopy drones, for which the weight and battery consumption must be optimized.
-This enables building a compact system that is capable of operating in dense forests.
-The main details of the algorithms used in the chosen solution are described briefly in the following and the complete system architecture of the prototype is presented in Figure 1.
-For more detailed information the readers are referred to the cited original sources.
+The solution was the only open-sourced camera-based solution proven to work with only one stereo camera making it possible to build a smaller and lighter platform.
+Additionally, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, even though it had a highly different structure than the typical boreal forests.
+The main details of the algorithms used in the chosen solution are briefly described in the following, and the complete system architecture of the prototype is presented in Figure 1.
+For more detailed information, the readers are referred to the cited original sources.
+Figure 1: The system architecture used in the data collection.
+In comparison to the original EGO-Planner-v2 article by Zhou et al.
+(2022), UWB-based Drift correction and Wireless trajectory broadcast network modules used for swarm navigation were omitted in this study.
+Stereo images were downsampled for VIO to guarantee real-time estimation.
+Full-resolution stereo images were also saved and post-processed to point clouds used in the DBH estimation.
 3.1.1 EGO-Planner-v2
@@
 (2022) was named in Github as EGO-Planner-v2 (ZJU FAST Lab Team 2022a).
-Even though the EGO-Planner-v2 is designed for swarm flying, in this study, only a single drone system was used.
-The planner is running on top of the Robot Operating System (ROS) (Quigley et al.
+Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments.
+The planner runs on top of the Robot Operating System (ROS) (Quigley et al.
 2009).
 The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985).
-The mapping module uses depth images as an input and the system adopts the fixed-sized circular buffers proposed by Usenko et al.
+The mapping module uses depth images as an input, and the system adopts the fixed-sized circular buffers proposed by Usenko et al.
 (2017) for maintaining the local map.
-The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map respectively.
-Virtual floor and virtual ceiling restrict the altitude where the algorithm is allowed to plan paths.
+The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively.
+The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
 In the trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al.
@@
 ${\min}\sum_{x}\lambda_{x}J_{x}$ (1)
-where where $J_{x}$ are the penalty terms and $\lambda_{x}$ are the corresponding weights.
+where $J_{x}$ are the penalty terms and $\lambda_{x}$ are the corresponding weights.
 Subscripts $x=\{s,t,d,o,u\}$ stand for the trajectory smoothness, total flying time, dynamical feasibility, obstacle avoidance, and the uniform distribution of constraint evaluation points along the trajectory, respectively.
-In addition to the above mentioned penalty terms, EGO-Planner-v2 includes also penalty terms for the swarm operations, but those are omitted here.
-A* is utilized as the global path-searching algorithm given as an input for the optimization algorithm.
+In addition to the above-mentioned penalty terms, EGO-Planner-v2 also includes penalty terms for the swarm operations, but those are omitted here.
+A* is utilized as the global path-search algorithm given as input for the optimization algorithm.
 The trajectory planning is continuously performed within a defined local planning distance towards a user-defined global goal until the goal is reached.
@@
 3.1.2 VINS-Fusion
-The VIO algorithm used by Zhou et al.
+The Visual-inertial odometry (VIO) algorithm used by Zhou et al.
 (2022) was VINS-Fusion (Qin, Li, and Shen 2018; Qin et al.
-2019b)(VINS stands for ”Visual-Inertial system”), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
+2019b)(VINS stands for ’Visual-Inertial system’), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
 VINS-Fusion detects Shi-Tomasi corner features (Shi and Tomasi 1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (Lucas and Kanade 1981).
-IMU data is used for preintegration, i.e.
-position, velocity, and orientation are integrated from IMU measurements between two frames.
-The state optimization is performed within a sliding window including the position, velocity, rotation, and IMU biases.
+IMU data is used for preintegration, i.e., position, velocity, and orientation are integrated from IMU measurements between two frames.
+The state optimization is performed within a sliding window, including the position, velocity, rotation, and IMU biases.
 To reduce the computational complexity, old measurements are marginalized into a prior of the estimation process.
-Certain frames are also selected as keyframes which are saved to a global pose graph after being marginalized out from the sliding window.
-Keyframes are selected based on the number of tracked features and the average parallax between the features in the newest frame and the latest keyframe.
+Certain frames are also selected as keyframes, which are saved to a global pose graph after being marginalized out from the sliding window.
+Keyframes are selected by considering the number of tracked features and the average parallax between these features in the most recent camera frame and the latest keyframe.
 VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate.
@@
 After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses.
-In addition to sliding window relocalization, also the global pose graph is optimized after loop detections to enforce the global consistency of the poses.
+In addition to sliding window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses.
 Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
 VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (Qin and Shen 2018).
-The method can be used for adjusting the time offset $t_{d}$ between the camera and the IMU online.
+The method can be used to adjust the time offset $t_{d}$ between the camera and the IMU online.
 In the swarm operations of the original article (Zhou et al.
 2022), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion.
-However, with a single drone of this study that function is omitted.
-The parameters used in the configuration of VINS-Fusion were chosen so that the real-time performance was guaranteed.
+However, with a single drone system in this study, that function is omitted.
+The parameters used in the configuration of VINS-Fusion were chosen so that real-time performance was guaranteed.
 In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion causing instability to the estimates, so the IMU noise values were increased more.
@@
 In this study, an open-source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used.
-The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner, and converts those to attitude and thrust setpoint.
-The setpoints are forwarded to PX4 flight controller software (PX4 Community 2023) responsible for low level attitude and angular rate controls.
-Figure 1: The system architecture used in the data collection.
-In comparison to the original EGO-Planner-v2 article by Zhou et al.
-(2022), UWB-based Drift correction and Wireless trajectory broadcast network modules used for swarm navigation were omitted in this study.
-Stereo images were downsampled for VIO to guarantee real-time estimation.
-Full-resolution stereo images were also saved and post-processed to point clouds used in the DBH estimation.
+The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints.
+The setpoints are forwarded to the PX4 flight controller software (PX4 Community 2023) responsible for the low-level attitude and angular rate controls.
 3.2 Drone hardware
 In the physical robot drone prototype, an onboard computer NVIDIA Jetson Orin NX (Santa Clara, California, USA) was used.
-The used PX4-compliant autopilot was Holybro Kakute H7 (Hongkong, China), which contained also the IMU.
-The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the Mavros-protocol was used for communication between the onboard computer and the autopilot.
+The used PX4-compliant autopilot was Holybro Kakute H7 (Hongkong, China), which also contained the IMU.
+The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the Mavros protocol was used for communication between the onboard computer and the autopilot.
 As a camera, Intel RealSense D435 (Intel Realsense Development Team 2023)(Keselman et al.
@@
 The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
-RealSense D435 is based on active stereoscopy and the camera has an infrared laser projector emitting dots that can be used for triangulation by stereo camera.
+RealSense D435 is based on active stereoscopy, and the camera has an infrared laser projector emitting dots that can be used for triangulation by the stereo camera.
 This can improve the triangulation in low-light conditions and to homogeneous surfaces that contain no features.
 However, the dots of the laser emitter cause problems for VIO since they are moving with the camera.
-In the previous setup described by (Karjalainen et al.
-2023), a modified version of ROS-RealSense driver was used by Zhou et al.
-(2021b) that turns the laser emitter of the camera constantly on and off making it possible to produce high-quality depth images in every other frame, and stereo images without the laser dots in every other frame.
+In the previous set-up described by (Karjalainen et al.
+2023), a modified version of the ROS-RealSense driver by Zhou et al.
+(2021b) was used.
+The modified driver turns the laser emitter of the camera constantly on and off, making it possible to produce high-quality depth images in every other frame and stereo images without the laser dots in every other frame.
 However, this solution comes with the drawback of halving the output frame rate of the camera.
@@
 2023), difficulties were encountered in detecting small branches with a depth camera resolution of $640\times 480$, leading to the decision to increase the depth camera resolution to a maximum value of $1280\times 720$.
-Stereo images for VIO were downsampled to resolution of $640\times 360$ to ensure real-time performance.
+The stereo images for VIO were downsampled at a resolution of $640\times 360$ to ensure real-time performance.
 As the maximum frame rate of D435 at a resolution of $1280\times 720$ is 30 Hz, the modified driver version would have resulted in a low frame rate of 15 Hz.
-To ensure the VIO performance in quick turns, the frame rate of 30 Hz was preferred, and hence disabled the IR emitter making the camera function as a passive IR stereo camera.
-The choice can be further justified with the fact that the usefulness of the IR emitter in forest conditions is questionable, since sunlight contains infrared light that can interfere with the emitted artificial features(Kuan, Ee, and Wei 2019), and the forests typically do not contain large homogeneous surfaces.
-The Realsense D435 onboard camera was configured to use the builtin automatic exposure control.
-The camera-IMU system was calibrated with an open-source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012)(Furgale, Rehder, and Siegwart 2013)(Maye, Furgale, and Siegwart 2013).
+To ensure the VIO performance in quick turns, the frame rate of 30 Hz was preferred, and hence, the IR emitter was disabled, making the camera function as a passive IR stereo camera.
+The choice can be further justified with the fact that the usefulness of the IR emitter in forest conditions is questionable since sunlight contains infrared light that can interfere with the emitted artificial features(Kuan, Ee, and Wei 2019), and the forests typically do not contain large homogeneous surfaces.
+The Realsense D435 onboard camera was configured to use the built-in automatic exposure control.
+The camera-IMU system was calibrated with an open-source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013).
 In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were calculated from recorded data where Aprilgrid (Olson 2011) was used as a calibration target.
-The IMU noise parameters needed for the calibration were first obtained with open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the obtained values were multiplied with 10 before the calibration as suggested in the project wiki of kalibr (ETH Zürich Autonomous Systems Lab 2023).
+The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (ETH Zürich Autonomous Systems Lab 2023).
 The computer, the camera, and the autopilot were attached to a 330 mm drone frame.
 The propellers had two pieces of 6 cm long blades each.
-The measured weight of the drone including the motors and the onboard computer was 791 g without the batteries, and 1153 g with the batteries.
-An overview of the used drone hardware is presented in Figure 2.
+The measured weight of the drone, including the motors and the onboard computer, was 791 g without batteries and 1153 g with the batteries.
+An overview of drone hardware is presented in Figure 2.
 Figure 2: The drone hardware used in the flight tests with a forward-facing Realsense D435 camera.
@@
 (2019); Hyyppä et al.
-(2020a, c), and are classified into three complexity categories, i.e., “easy”, “medium”, and “difficult” based on, mainly tree density and amount of the understory vegetation in the plots.
-The first plot has been classified as ”medium”, and the second one has been classified as ”difficult”.
-The medium difficulty plot had a tree density of 650 trees/ha and a mean diameter at breast height (DBH) of 28 cm.
+(2020a, c), and are classified into three complexity categories, i.e., ’easy’, ’medium’, and ’difficult’ based on, mainly tree density and amount of vegetation under the canopy in the plots.
+The third test area, a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E) was used only to test the loop detection performance of VINS-Fusion.
+The first plot in Evo has been classified as ’medium’ (Figure 3), and the second has been classified as ’difficult’ (Figure 4).
+The medium-difficulty plot had a tree density of 650 trees/ha and a mean DBH of 28 cm.
 The most common tree species in the plot were Norway spruce (Picea abies (L.) H.Karst.)(81.2 %) and Scots pine (Pinus sylvestris L.)(9.4 %).
@@
 Only trees with a DBH greater than 5 cm were included in these statistics.
-From now on, these two test areas are called ”Evo-medium” and ”Evo-difficult”, respectively.
-For georeferencing purposes, circular coded targets provided by Agisoft Meatahsape (Agisoft LLC, St.
-Petersburg, Russia) were placed within the forest plots.
-The targets had a diameter of 24.7 cm, and the middle dots had a size of 7 cm.
-Relative distances between the coded targets were measured with Leica Nova TS60 total station with a precision of approximately 2mm (Leica Geosystems 2020).
-Figure 5 shows coded targets in ”Evo-difficult” forest during the test flights.
-The third test area was a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E).
-Aim of the experiment there was to test the loop detection performance of VINS-Fusion by walking with the drone in an environment without coded targets.
-Estimated average tree density in the area was approximately 2000 trees/ha.
-Figure 3: View from the first takeoff location in the Evo-medium test area. The takeoff location was changed after five test flights.
-Figure 4: View from the takeoff location in the Evo-difficult test forest
+From now on, these two test areas are called ’Evo-medium’ and ’Evo-difficult’, respectively.
+Figure 3: View from the first takeoff location in the ’Evo-medium’ test area. The takeoff location was changed after five test flights.
+Figure 4: View from the takeoff location in the ’Evo-difficult’ test forest.
+For reference purposes, circular GCPs provided by Agisoft Metashape (Agisoft LLC, St.
+Petersburg, Russia) were placed within the forest plots used in the flight tests.
+The targets had a diameter of 24.7 cm and the middle dots had a size of 7 cm.
+Relative distances between the GCPs were measured with Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020).
+Figure 5 shows GCPs in ’Evo-difficult’ during the test flights.
+Figure 5: GCPs for reference trajectory generation in ’Evo-difficult’ forest.
+Targets were spread widely to the test site to increase the probability that there would be targets visible in the recorded camera data regardless the path that the robot takes.
+GCPs were not used in the third test forest, ’Paloheinä’.
+The experiment there aimed to test the loop detection performance of VINS-Fusion in a homogeneous forest environment without any artificial targets.
+The data were collected by walking with the drone.
+The estimated average tree density in the area was approximately 2000 trees/ha.
 3.4 Flight tests
-Test flights in Evo were conducted on 18th October, 2023.
+Test flights in Evo were conducted on 18th October 2023.
 Multiple test flights were performed to evaluate the forest-type effects on the performance of autonomous navigation.
-In total, seven test flights were conducted in the ”Evo-medium” test area and nine test flights in the ”Evo-difficult” test area.
+In total, seven test flights were conducted in the ’Evo-medium’ test area and nine test flights in the ’Evo-difficult’ test area.
 In each test environment, either the takeoff location or the takeoff heading was altered during the test flight campaign to force the drone to follow different paths.
-In the ”Evo-medium” test area, the takeoff location was changed after five test flights.
-In the ”Evo-difficult” test area the takeoff heading was changed after six test flights.
+In the ’Evo-medium’ test area, the takeoff location was changed after five test flights.
+In the ’Evo-difficult’ test area, the takeoff heading was changed after six test flights.
 The target flight velocity was 1 m/s in all flight tests.
-The goal was set 36 m forward in the first takeoff location in the ”Evo-medium” test area, 34 m forward in the second takeoff location in the ”Evo-medium” test area, and 42 m forward in the takeoff location in the ”Evo-difficult” test area.
-The maximum flight height was set to 2.25 m relative to the takeoff location in the ”Evo-medium” test area.
-For the first test flight in the ”Evo-difficult” test area, the maximum flight height was set to 2.5 m with respect to the takeoff location and was increased to 2.75 m for the rest of the flights.
+The goal was set 36 m forward in the first takeoff location in the ’Evo-medium’ test area, 34 m forward in the second takeoff location in the ’Evo-medium’ test area, and 42 m forward in the takeoff location in the ’Evo-difficult’ test area.
+The maximum flight height was set to 2.25 m relative to the takeoff location in the ’Evo-medium’ test area.
+For the first test flight in the ’Evo-difficult’ test area, the maximum flight height was set to 2.5 m with respect to the takeoff location and was increased to 2.75 m for the rest of the flights.
 The campaign to evaluate the loop detection capability took place on 26th January 2024.
@@
 The data collection took place during winter, with the forest covered in snow.
-Figure 5: Control targets for reference trajectory generation in ”Evo-difficult” forest.
-Targets were spread widely to the test site to increase the probability that there would be targets visible in the recorded camera data regardless the path that the robot takes.
 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
@@
 (2018).
-Every sixth stereo pair were used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m.
-In the image orientation processing, quality setting was set to high meaning that two times downsampled images were used; the settings for number of key points and tie points per image were 40,000 and 4,000, respectively.
-To enhance the quality of the post-processed reference trajectories, ground control points (GCPs) were used in the processing with an accuracy setting of ±$\pm 0.005$ m.
+Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m.
+In the image orientation processing, the quality setting was set to high meaning that two times downsampled images were used.
+The settings for the number of key points and tie points per image were 40,000 and 4,000, respectively.
+To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of ±$\pm 0.005$ m.
 Automatic outlier removal was performed using gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy.
-Subsequently, a bundle-adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP).
+Subsequently, a bundle adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP).
 Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering.
 Finally, the EOPs and dense point clouds were exported.
-Point clouds were derived using datasets from four flights conducted in the ”Evo-medium” test area, each designated as follows: I: the first flight, II: the second flight, III: the third flight, IV: the fourth flight, and I-IV: combined processing of all four flights (Combined).
+Point clouds for DBH estimation were derived using datasets from four flights conducted in the ’Evo-medium’ test area, each designated as follows: I: the first flight, II: the second flight, III: the third flight, IV: the fourth flight, and I-IV: combined processing of all four flights (Combined).
 The processing was carried out using all GCPs.
@@
 Reference trajectories were generated for the flights using SfM processing.
-In each case, the referenece trajectory was derived from the first takeoff heading and orientation in both Evo test environments.
+In each case, the reference trajectory was derived from the first takeoff heading and orientation in both Evo test environments.
+Figure 6 visualizes the GCPs, stereo images, and reference trees during the point cloud processing in the Agisoft Metashape.
+Figure 6: GCPs, stereo images and reference trees visualized during the point cloud processing in the Agisoft Metashape (Metashape Development Team, 2021).
+Detected reference trees are marked, while those not detected are highlighted with a orange box.
 3.6 Tree parameter estimation
-In this study, trees were detected and their DBHs estimated using an automatic stem diameter estimation algorithm originally proposed by Hyyppä et al.
-(2020b), and further developed by Hyyppä et al.
+In this study, trees were detected, and their DBHs were estimated using an automatic stem diameter estimation algorithm proposed by Hyyppä et al.
+(2020b) and further developed by Hyyppä et al.
 (2020a) to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS.
@@
 The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments, and 5-s-long temporal segments.
-The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus the temporal segmentation was not performed for the TLS data.
+The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data.
 The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al.
-1996)), and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle.
+1996)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle.
 The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA).
 The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation.
-The diameters were processed into a stem curve by automatically removing clear outliers followed by fitting a smoothing spline to the remaining diameter estimates as a function of height.
+The diameters were processed into a stem curve by automatically removing clear outliers, followed by fitting a smoothing spline to the remaining diameter estimates as a function of height.
 Finally, the DBH estimates were calculated by interpolating the stem curve at the height of 1.3 m above the ground.
@@
 3.7.1 Autonomous flying performance
-The performance and reliability of the autonomous navigation of the prototype were evaluated based on two factors: the reliability of the obstacle avoidance and the accuracy of the VIO localization.
+The performance and reliability of the autonomous flight of the prototype were evaluated based on two factors: the reliability of the obstacle avoidance and the accuracy of the VIO localization.
 Obstacle avoidance reliability was evaluated based on the mission success.
 A mission was considered successful if the drone reached the end of the flight without major collisions.
-Minor touches to thin vegetation or branches were allowed, if the drone was able to continue the mission after that.
+Minor touches to thin vegetation or branches were allowed if the drone was able to continue the mission after that.
 If the given goal was inside an obstacle, the mission was considered successful if the drone managed to fly next to the goal.
@@
 The trajectory was considered smooth if the system did not perform any emergency stops during the flight.
-The accuracy of the VIO localization was assessed by calculating the Absolute Trajectory Error (ATE), as proposed by Sturm et al.
-(2012), with respect to the SfM reference trajectories.
+The accuracy of the VIO localization was evaluated by calculating the Absolute Trajectory Error (ATE), as proposed by Sturm et al.
+(2012), with respect to the post-processed SfM reference trajectories.
 In this study, ATE was calculated only for the position estimate, and the error in the orientation estimate was not evaluated.
 ATE was calculated using the trajectory evaluation toolbox by Zhang and Scaramuzza (2018), which aligns the trajectory estimates with the reference trajectories with the method by Umeyama (1991).
-The root mean square error (RMSE) of the individual state errors is then calculated to generate a single metric value representing the entire trajectory’s accuracy..
-In the mathematical form the error can be presented as
+After aligning the trajectories, the toolbox calculates the RMSE of the individual state errors to generate a single metric error value representing the accuracy of the entire trajectory.
+In the mathematical form, the error can be presented as
 $\text{ATE}_{\text{pos}}=\left(\frac{1}{N}\sum_{i=0}^{N-1}||\Delta\mathbf{p}_{i}||^{2}\right)^{\frac{1}{2}}$ (2)
@@
 For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr.
-However, the calibration estimate is never perfect and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU.
+However, the calibration estimate is never perfect, and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU.
 For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests.
-The estimates were performed with the drone’s onboard computer, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of online estimation.
-Since the coded targets are easily detectable, unique, and unnatural objects, the loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets.
-Since there were no coded targets for forming the reference trajectories, only the ability to detect the loops inside dense natural forest was tested, and the accuracy of the loop corrected trajectory was not evaluated.
-VINS-Fusion estimate with the loop detection algorithm was ran for the collected data with the drone’s onboard computer, and the effect of amount of features in pictures for loop detection capability was examined.
+The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of online estimation.
+The loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets that could affect the system performance.
+Since there were no GCPs for forming the reference trajectories, only the ability to detect the loops inside dense natural forests was tested, and the accuracy of the loop-corrected trajectory was not evaluated.
+VINS-Fusion estimate with the loop detection algorithm was run for the collected data with the onboard computer of the drone, and the effect of a maximum number of features in pictures for loop detection capability was examined.
 3.7.2 DBH estimation performance
-The accuracy of the DBH estimation from the drone point cloud data was evaluated using Terrestrial Laser Scanning (TLS) data as a reference.
-The TLS scannings were conducted in April 2020, and tree height, diameter at breast height (DBH), and other parameters were estimated from the dataset.
-The estimated accuracy of the DBH measurement was less than 1 cm.
-Several metrics were used to evaluate the accuracy of the DBH estimation, including Root Mean Square Error (RMSE), Relative RMSE (RMSE%), Bias, Relative Bias (bias%), and Standard Deviation.
+The accuracy of the DBH estimation from the drone point cloud data was evaluated using TLS data as a reference.
+The TLS scannings were conducted in April 2020, and tree height, DBH, and other parameters were estimated from the dataset.
+The estimated error of the DBH measurement was less than 1 cm.
+Several metrics were used to evaluate the accuracy of the DBH estimation, including RMSE, Relative RMSE (RMSE %), Bias, Relative Bias (bias %), and Standard Deviation.
 The RMSE measures the square root of the average of the squared differences between the estimated and reference DBH values.
@@
 where $\bar{d}_{\text{reference}}$ is the mean of the reference DBH values, and SD is the standard deviation.
-Additionally, the completeness of stem detection was evaluated by comparing the total number of reference trees found from the drone-collected point cloud data to the total number of reference trees within the point cloud boundary.
+The completeness of stem detection was evaluated by comparing the total number of reference trees found from the drone-collected point cloud data to the total number of reference trees within the point cloud boundary.
 $c=\frac{r}{t}\times 100$ (8)
 where $c$ represents completeness (%), $r$ is the number of reference trees found, and $t$ is the total number of reference trees.
-The impact of tree growth between the collection of drone data and reference measurements was not taken into account.
-Variability in tree growth, influenced by factors such as species, locations, health, age, is acknowledged as a source of error.
+The impact of tree growth between the collection of drone data and reference measurements (four growth seasons) was not taken into account.
+Variability in tree growth, influenced by factors such as species, locations, health, and age, is acknowledged as a source of error.
 Considering that the tree growth was relatively constant in the test area, forming a major component of the bias in the estimates.
@@
 The task of avoiding collisions during the flight is naturally more demanding in a dense forest than in a sparse forest.
-Consistent with these expectations, the trajectories were smoother and the success rate was better in ”Evo-medium” than in ”Evo-difficult” test area.
-In ”Evo-medium” all performed test flights were successful and five out of seven test flights were performed with smooth trajectories.
-In ”Evo-difficult” one of nine test flights failed and none of the test flights was performed with a smooth trajectory.
-The reason for emergency stops in both environments was late detection of thin and dry spruce branches.
-In the failed test flight in ”Evo-difficult” the reason was also a collision to a similar thin low branch.
-The test flight success rates in both test forests are summarized in the Table 3. Figure 7 presents examples of performed trajectories and obtained point clouds in all test forests.
+Consistent with these expectations, the trajectories were smoother and the success rate was better in ’Evo-medium’ than in ’Evo-difficult’ test area.
+In ’Evo-medium’, all performed test flights were successful and five out of seven test flights were performed with smooth trajectories without any emergency stops during the flight.
+In ’Evo-difficult’, one of nine test flights failed.
+None of the test flights in ’Evo-difficult’ was performed with a smooth trajectory.
+However, despite the emergency stops during the flight, the system was still able to continue the mission after the stops, leading to a successful mission in eight of the nine test flights.
+The reason for emergency stops in both environments was the late detection of thin and dry spruce branches.
+In the failed test flight in ’Evo-difficult’ the reason was also a collision to a similar thin low branch.
+The test flight success rates in both test forests are summarized in Table 3. Figure 8 presents examples of performed trajectories and obtained point clouds in both test forests.
 Table 3: Test flight success in the test forests
 Test areas Forest type Density (trees/ha) Flight lengths (m) Successful flights Smooth trajectories
-”Evo-medium” Old Spruce forest 650 34 - 36 7/7 5/7
-”Evo-difficult” Young Spruce forest 2000 42 8/9 0/8
-(a) ”Evo-medium”
-(b) ”Evo-difficult”
-Figure 7: Examples of drones own navigation grid maps with performed trajectories (left) and obtained point clouds (right) in (a) ”Evo-medium” and (b) ”Evo-difficult”
+’Evo-medium’ Old Spruce forest 650 34 - 36 7/7 5/7
+’Evo-difficult’ Young Spruce forest 2000 42 8/9 0/8
+(a) ’Evo-medium’
+(b) ’Evo-difficult’
+Figure 8: Examples of drones’ own navigation grid maps with performed trajectories (left) and obtained point clouds (right) in (a) ’Evo-medium’ and (b) ’Evo-difficult’
 4.1.2 VIO
-The trajectory estimate error against the reference trajectory was successfully calculated for five flights in the medium forest and for three flights in the difficult forest.
-In the difficult forest, one of the flights was performed without coded targets making the reference trajectory non-comparable to others, in one flight there was a gap in the recorded onboard data due to a data recording error during a flight, and in one flight Metashape failed to retrieve the reference path.
-Average $\text{ATE}_{\text{pos}}$ was 0.50 m in the medium forest and 0.34 m in the difficult forest.
-The standard deviation between $\text{ATE}_{\text{pos}}$ was 0.06 m in the medium forest and 0.07 m in the difficult forest.
-The VINS-Fusion estimates with online $t_{d}$ estimation were calculated afterwards for the same flights.
-With online $t_{d}$ estimation, average $\text{ATE}_{\text{pos}}$ was 0.47 m in the medium forest and 0.33 m in the difficult forest.
-The standard deviation between $\text{ATE}_{\text{pos}}$ was 0.02 m in the medium forest and 0.04 m in the difficult forest.
-As the small standard deviation between the $\text{ATE}_{\text{pos}}$ estimates indicates, the error between the VINS-Fusion estimates and the reference trajectories was similar in all flights.
-In every flight, VINS-Fusion was estimating the shape of the trajectory well but underestimating the flight distance.
-Figure 8 presents an example of VINS-Fusion trajectory estimate and the reference trajectory in the difficult test forest.
-Table 4 presents the VINS-Fusion estimate errors in all flights.
+The trajectory estimate error against the reference trajectory was calculated for the first five flights in ’Evo-medium’ and for three flights in the difficult forest.
+The flights from the second takeoff location in ’Evo-medium’ and from the second takeoff orientation in ’Evo-difficult’ were not used in the VIO estimation since GCPs used in the reference trajectory processing were positioned along the first flight trajectories in both environments.
+In ’Evo-difficult’, one of the flights from the first takeoff orientation was performed without GCPs, leading to a non-comparable reference trajectory, one of the flights had a data recording error during a flight, and in one flight, Metashape failed to retrieve the reference path.
+Figure 9 shows an example of the VINS-Fusion trajectory estimate compared to the reference in ’Evo-difficult’ while Table 4 summarizes the VINS-Fusion estimation errors across all flights.
+The average $\text{ATE}_{\text{pos}}$ was 0.50 m in ’Evo-medium’ and 0.34 m in ’Evo-difficult’ with standard deviations of 0.06 m and 0.07 m, respectively, without online $t_{d}$ estimation.
+The VINS-Fusion estimates with online $t_{d}$ estimation were computed post-flight for the same trajectories, yielding an average $\text{ATE}_{\text{pos}}$ of 0.47 m in ’Evo-medium’ and 0.33 m in ’Evo-difficult’ with standard deviations of 0.02 m and 0.04 m, respectively.
+(a) Top view.
+(b) Side view.
+Figure 9: Example of a VINS-Fusion trajectory estimate and a reference trajectory obtained with Agisoft Metashape in the ’Evo-difficult’ forest from a) top and b) side.
 Table 4: VINS-Fusion estimation Errors in all test flights
 Test area Flight No. $\text{ATE}_{\text{pos}}$ (constant $t_{d}$) $\text{ATE}_{\text{pos}}$ (online estimated $t_{d}$)
-”Evo-medium” 1 0.5567 0.4816
-”Evo-medium” 2 0.5599 0.4790
-”Evo-medium” 3 0.4538 0.4769
-”Evo-medium” 4 0.4945 0.4722
-”Evo-medium” 5 0.4313 0.4400
-”Evo-difficult” 1 0.3787 0.3340
-”Evo-difficult” 2 0.3853 0.3727
-”Evo-difficult” 3 0.2680 0.2962
-Figure 9 shows an example of VINS-Fusion tracking image during a test flight in the dense test forest.
+’Evo-medium’ 1 0.5567 0.4816
+’Evo-medium’ 2 0.5599 0.4790
+’Evo-medium’ 3 0.4538 0.4769
+’Evo-medium’ 4 0.4945 0.4722
+’Evo-medium’ 5 0.4313 0.4400
+’Evo-difficult’ 1 0.3787 0.3340
+’Evo-difficult’ 2 0.3853 0.3727
+’Evo-difficult’ 3 0.2680 0.2962
+The small standard deviations between the $\text{ATE}_{\text{pos}}$ estimates indicated that the error between the VINS-Fusion trajectories and the reference trajectories was similar across all flights.
+In each case, VINS-Fusion accurately captured the trajectory shape but systematically underestimated the total flight distance.
+Figure 10 shows an example of a VINS-Fusion tracking image during a test flight in ’Evo-difficult’.
 The algorithm detects stable features to track from all parts of the picture.
-The image shows also flying dry leaves blown up by the air flow from the propellers of the drone.
+The image also shows flying dry leaves blown up by the airflow from the propellers of the drone.
 The moving leaves did not cause visible disturbances to the tracking of the pose.
-(a) Top view
-(b) Side view
-Figure 8: Example of a VINS-Fusion trajectory estimate and a reference trajectory obtained with Agisoft Metashape in the difficult forest from a) top and b) side.
 Figure 10: Example of VINS-Fusion tracking image visualizing the tracked features.
@@
 Figure 11 shows examples of detected loop closure image pairs from both paths.
-The number of detected loops can be increased by increasing the amount of features per image in the VINS-Fusion configuration.
+The number of detected loops can be increased by increasing the number of features per image in the VINS-Fusion configuration.
 By reducing the minimum distance between features from 40 to 35 pixels, VINS-Fusion detected a loop closure between 52 image pairs in the shorter path and 23 image pairs in the longer path.
@@
 Re-projection errors varied between 0.29 and 0.39 pixels.
-The estimated flight GSDs (Ground Sampling Distance) ranged between 4.08 and 6.17 mm at ground level.
+The estimated GSDs (Ground Sampling Distance) ranged between 4.08 and 6.17 mm at ground level.
 Distances from the trees varied between 0.55 and 7.94 m, resulting in GSDs of 0.85-12.34 mm at the objects of interest.
 Point clouds from datasets II-IV had similar point densities of 2.63-2.84 points/cm2, while dataset I provided a slightly higher point density of 3.58 points/cm2.
-Results of accuracy estimation using Check Points (CPs) are shown in Table 6.
-The cases without GCPs or with one GCP at the beginning of the trajectory provided similar 2D and 3D RMSEs, of 11.36-11.72 cm and 11.37-11.73 cm, respectively.
-When two GCPs were used, the 2D and 3D RMSE improved to 5.26 cm and 5.48 cm, respectively.
-Adding an additional GCP resulted in a significant improvement in the 2D and 3D RMSE, reducing them to 1.41 cm and 1.48 cm, respectively.
 Table 5: Photogrammetric processing results.
 Img.
-count: Number images/Number of aligned images; Number of GCPs; Completeness % (Number of trees found/Number of total trees); Reprojection error (pix); Number of tie points (in millions); Point density (points(cm2) in ”Evo-medium” test site.
+count: Number images/Number of aligned images; Number of GCPs; Completeness % (Number of trees found/Number of total trees); Reprojection error (pix); Number of tie points (in millions); Point density (points/cm2) in ’Evo-medium’ test site.
 I: Trajectory 1, II: Trajectory 2, III: Trajectory 3, IV: Trajectory 4, Combined: Trajectories 1-4 merged, and for Trajectory 1 I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
@@
 I3GCP 650/650 3 79.31 (23/29) 0.39 197132 3.54
-Table 6: Accuracy estimation (Root Mean Square Error; RMSE) using Check Points (CPs) and Camera Location in ”Evo-medium” test site, for trajectory 1: I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
+Results of accuracy estimation using Check Points (CPs) are shown in Table 6.
+The cases without GCPs or with one GCP at the beginning of the trajectory provided similar 2D and 3D RMSEs of 11.36-11.72 cm and 11.37-11.73 cm, respectively.
+When two GCPs were used, the 2D and 3D RMSEs improved to 5.26 cm and 5.48 cm, respectively.
+Adding an additional GCP resulted in a significant improvement in the 2D and 3D RMSE, reducing them to 1.41 cm and 1.48 cm, respectively.
+Table 6: Accuracy estimation (RMSE) using Check Points (CPs) and Camera Location in ’Evo-medium’ test site, for trajectory 1: I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
 Dataset Check Points (CPs) RMSE (cm) Camera location RMSE (cm)
@@
 4.2.2 Tree detection
-Figure 12 shows post-processed trajectories for dataset I-IV, as well as reference and detected trees, and the boundaries of the point clouds.
+Figure 12 shows post-processed trajectories for dataset I-IV, as well as reference and detected trees and the boundaries of the point clouds.
 I and II followed almost the same trajectory, while III and IV had similar trajectories.
-Dataset I trajectory was smoothest and it did not include any sharp turns.
-Other 3 trajectories had one or more sharper turns.
-For datasets I, II, III and IV, the total number of reference trees inside the point cloud boundary was 29, 29, 31 and 32, of which 23, 18, 19 and 20 were found from drone-collected point cloud.
+Dataset I trajectory was the smoothest, and it did not include any sharp turns.
+The other three trajectories had one or more sharper turns.
+For datasets I, II, III, and IV, the number of reference trees inside the point cloud boundary was 29, 29, 31, and 32, of which 23, 18, 19, and 20 were found from the drone-collected point clouds, respectively.
 The resulting completeness values were 79.31 %, 62.07 %, 61.29 % and 62.5 %, respectively.
 The total number of different trees found was 29 out of 33 (87,88 %) from dataset I-IV.
-However, when all the flights were processed simultaneously (combined) that resulted only $3$ completeness, where only 12 of 36 trees was found.
-When utilizing 0, 1, 2 and 3 gcps in dataset I completeness values were 72.41 %, 68.97 %, 75.86 % and 79.31 %, respectively (Table 5).
-Figure 11: GCPs, stereo images and reference trees visualized during the point cloud processing in the Agisoft Metashape (Metashape Development Team, 2021).
-Detected reference trees are marked, while those not detected are highlighted with a orange box.
-In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the drone’s trajectory, despite these trees being well-reconstructed in the point cloud.
-Similarly, number of well-reconstructed trees but not detected by algorithm was 4, 5 and 5, in dataset II, III and IV, respectively.
-The three extra trees that were not identified in dataset I, were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud.
-The datasets II trough IV each contained 7 poorly reconstructed trees.
-Many of these trees were located at a greater distance from the trajectory, while others were obscured from the camera’s view by intervening trees.
+However, when all the flights were processed simultaneously (combined) that resulted in only $3$ completeness, where only 12 of 36 trees were found.
+When utilizing 0, 1, 2, and 3 GCPs in the dataset I, completeness values were 72.41 %, 68.97 %, 75.86 %, and 79.31 %, respectively (Table 5).
+Figure 12: Post-processed trajectories, reference trees and found trees in Dataset (a) I (b) II (c) III (d) IV
+In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the flight trajectory despite these trees being well-reconstructed in the point cloud.
+Similarly, the number of well-reconstructed trees but not detected by the algorithm was 4, 5, and 5, in datasets II, III, and IV, respectively.
+The three extra trees that were not identified in dataset I were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud.
+The datasets II through IV each contained seven poorly reconstructed trees.
+Many of these trees were located at a greater distance from the trajectory, while others were obscured from the view of the camera by intervening trees.
 4.2.3 DBH estimation
-RMSE and bias of the DBH estimates of datasets I-IV were similar, ranging from 3.33 to 3.86 cm (10.69-12.98%) and -0.67 to 1.40 (-2.11-4.71%), respectively (Table 7).
-Processing all the flights simultaneously (Combined) provided only marginally better results with an RMSE of 2.91 cm (9.14%) and a bias of 1.51 cm (4.74%).
-When considering the impact of the number of GCPs, the utilization of 3, 2, 1 and 0 GCPs in photogrammetric processing yielded slightly poorer results compared to those obtained with all GCPs.
-The RMSE and bias values ranged from 3.93 to 4.37 cm (13.68%-14.66%) and 1.12 to 2.13 cm (3.89-7.55%), respectively (Table 7).
+RMSE and bias of the DBH estimates of datasets I-IV were similar, ranging from 3.33 to 3.86 cm (10.69 %-12.98 %) and -0.67 to 1.40 (-2.11 %-4.71 %), respectively (Table 7).
+Processing four flights simultaneously (Combined) provided only marginally better results with an RMSE of 2.91 cm (9.14 %) and a bias of 1.51 cm (4.74 %).
+When considering the impact of the number of GCPs, the utilization of 3, 2, 1, and 0 GCPs in photogrammetric processing yielded slightly poorer results compared to those obtained with all GCPs.
+The RMSE and bias values ranged from 3.93 to 4.37 cm (13.68 %-14.66 %) and 1.12 to 2.13 cm (3.89 %-7.55 %), respectively (Table 7).
 Tree size had an impact on the DBH estimation accuracy (Table 7).
-For small and medium-sized trees, with a DBH smaller than 30 cm, DBH estimations achieved superior performance in all dataset, with RMSE and bias varying from 1.16 to 2.75 cm (5.74-12.47%) and -0.01 to 0.62 cm (-0.06-3.14%), respectively.
-Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99-25.44%) and -0.66 to 2.08 cm (-1.53-5.65%), respectively.
-This is demonstrated in Figure 14, illustrating a positive correlation between tree diameter and DBH estimation error.
-Notably, when the tree diameter exceeded 30 cm, the DBH estimation errors became more pronounced.
-Analysis of the impact of the camera distance from trees did not reveal significant correlation between tree distance and DBH estimation errors (Figure 13).
-Figure 12: Post-processed trajectories, reference trees and found trees in Dataset (a) I (b) II (c) III (d) IV..
-Figure 13: Scatter plots of DBH errors with respect to distance to trees. Dataset (a) I (b) II (c) III (d) IV.
-Figure 14: Scatter plots of DBH errors with respect to diameter of trees. Dataset (a) I (b) II (c) III (d) IV
-Table 7: Accuracy of the stem diameter breast height (DBH) in ”Evo-medium” test site.
+For small and medium-sized trees with a DBH smaller than 30 cm, DBH estimations achieved superior performance in all datasets, with RMSE and bias varying from 1.16 to 2.75 cm (5.74 %-12.47 %) and -0.01 to 0.62 cm (-0.06 %-3.14 %), respectively.
+Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99 %-25.44 %) and -0.66 to 2.08 cm (-1.53 %-5.65 %), respectively.
+Analysis of the impact of the camera distance from trees did not reveal a significant correlation between tree distance and DBH estimation errors.
+Table 7: Accuracy of the stem diameter breast height (DBH) in ’Evo-medium’ test site.
 Dataset is split in 3 categories: All, DBH $<$ 30 cm and DBH $>$ 30 cm trees.
 Datasets: I: Flight 1, II: Flight 2, III: Flight 3, IV: Flight 4, Combined: Flights I-IV merged, and for Dataset I: I0GCP: with no GCPs, I1GCP: with 1 GCP, I2GCP: with 2 GCPs, I3GCP with 3 GCPs.
-Dataset Number of trees All $<$30 cm $>$30 cm RMSE (cm) All $<$30 cm $>$30 cm RMSEAll $<$30 cm $>$30 cm Bias (cm) All $<$30 cm $>$30 cm BiasAll $<$30 cm $>$30 cm Standard Deviation All $<$30 cm $>$30 cm
+Dataset Number of trees All $<$30 cm $>$30 cm RMSE (cm) All $<$30 cm $>$30 cm RMSE % All $<$30 cm $>$30 cm Bias (cm) All $<$30 cm $>$30 cm Bias % All $<$30 cm $>$30 cm Standard Deviation All $<$30 cm $>$30 cm
 I 23 14 9 3.86 2.56 5.29 12.98 12.47 11.98 1.40 0.17 1.23 4.71 0.85 2.78 3.60 2.54 4.26
@@
 The performance of autonomous navigation was promising in both test forests.
-The obstacle avoidance success rate was 100% in the medium forest and 87.5% in the difficult forest.
+The obstacle avoidance success rate was 100 % in the ’Evo-medium’ forest and 87.5 % in the ’Evo-difficult’ forest.
 However, the smoothness of the trajectories suffered from the increasing forest density.
-While in the medium forest, 71.4% of the flights had a smooth trajectory, in the difficult forest at least one emergency stop was needed in all of the flights.
+While in ’Evo-medium’ 71.4 % of the flights had a smooth trajectory, in ’Evo-difficult’ at least one emergency stop was needed in all of the flights.
+However, apart from the one failed flight in ’Evo-difficult’, the system was able to do a successful trajectory re-planning and finish the mission after the emergency stops.
 Further consideration should be given to finding the optimal sensor for detecting thin and dry branches earlier.
-Nonetheless, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium difficulty level boreal forests.
-The accuracy of the position estimate by VINS-Fusion was satisfying for short flights performed in this study, but in the longer flights, the error would start to accumulate making the position errors high.
+Nonetheless, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium-difficulty level boreal forests.
+The accuracy of the position estimate by VINS-Fusion was satisfying for short flights performed in this study, but in the longer flights, the error would start to accumulate, making the position errors high.
 However, the standard deviation between the errors between the test flights was small, and the VINS-Fusion systematically estimated the flight paths shorter than in reality (as can be seen from Figure 9).
-This suggests, that a notable part of the position error could be eliminated by improving the calibration of the sensor setup.
-The results suggest that the estimates by VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration.
+This indicates that a notable part of the position error could be eliminated by improving the system calibration.
+The results indicate that the estimates by VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration.
 However, further development of the prototype should include implementing hardware-level synchronization between the camera and the IMU of the autopilot.
-The position error can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests showed that VINS-Fusion could detect loop closures even in a homogeneous forest environment.
-Increasing the number of features tracked by VINS-Fusion increases the number of detected loops.
-However, increasing the number of features increases also the computational burden of the system potentially disrupting the real-time estimation.
-Finding the optimal value for the tracked features as well as measuring the actual improvement of the loop closures to the position estimate requires further studies.
+The position error can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests show that VINS-Fusion could detect loop closures even in a homogeneous forest environment.
+Increasing the number of features tracked by VINS-Fusion increased the number of detected loops.
+However, increasing the number of features also increases the computational burden of the system, potentially disrupting the real-time estimation.
+Finding the optimal value for the tracked features, as well as measuring the actual improvement of the loop closures to the position estimate, requires further studies.
 5.2 Accuracy of forest measurements
-The proposed system was assessed by comparing it to other relevant studies (Section 2.2, Table 1) considering the accuracy, system autonomy and weight, and complexity of test environment.
-The highest accuracy UAV SfM-MVS results were presented by Shimabuku, Konoshima, and Ota (2023) with an RMSE of 0.4 cm (3.2%).
-However, their test site included only 7 reference trees.
-Kuželka and Surovỳ (2018) and Krisanski, Taskhiri, and Turner (2020) had more reference trees and achieved similar DBH accuracy compared to that in this study, but the flights were piloted manually.
-Notably, the SfM-MVS approaches involved light-weight systems.
-The fully automated under-canopy UAV Lidar system (ULS) by Liang et al.
+The proposed system was assessed by comparing it to other relevant studies (Section 2.1, Table 1) considering the accuracy, system autonomy, weight, and complexity of the test environment.
+The RMSEs reported in the relevant studies varied from 3.2 % to 61.41 %.
+The highest accuracy UAV SfM-MVS results were presented by Shimabuku, Konoshima, and Ota (2023) with RMSEs of 0.4 cm (3.2 %) and 0.7 cm (3.6 %) in two test forest sites.
+However, their test sites included only 7 and 18 reference trees, and the semi-autonomous commercial drone system needed manual control to avoid horizontal obstacles.
+Kuželka and Surovỳ (2018), He et al.
+(2025), and Krisanski, Taskhiri, and Turner (2020) had more reference trees and achieved similar DBH accuracy compared to that in this study, but the flights were piloted manually.
+Notably, the SfM-MVS approaches involved commercial lightweight drone systems, except He et al.
+(2025) where 360° camera was mounted as an additional payload to heavier LiDAR-based custom drone.
+The most accurate under-canopy LiDAR-based DBH estimates were presented by Hyyppä et al.
+(2020a), achieving an RMSE of 0.6 cm (2.3 %) and bias of 0.34 cm (1.3 %) and 1.1 cm (3.5 %) and 0.12 cm (0.4 %) in two sparse forest plots.
+However, their system weighed around 10 kg and was piloted manually.
+Autonomously operating under-canopy LiDAR-based drones were utilized in the studies by Muhojoki et al.
+(2024), Liang et al.
+(2024), and Cheng et al.
+(2024).
+While the commercial Hovermap drone system utilized in Muhojoki et al.
+(2024) achieved the lowest RMSE of these, the system was able to fly only in sparse and clean forests due to its size.
+The automated under-canopy UAV Lidar system by Liang et al.
 (2024) weighted 5 kg and provided an RMSE of 4.48 cm (21.37 %) and bias of 5.01 (14.91 %) at best.
-Wang et al.
-(2021) combined data from above- and under-canopy drone with a Riegl miniVUX-1UAV lidar (Riegl GmbH, Horn, Austria), resulting in an RMSE of 7.95 cm (33.35%) and bias of 3.72 cm (15.61%).
-Their under-canopy flight was carried out manually by a pilot and the system weighted approximately 10 kg.
-Muhojoki et al.
-(2024) carried out manually flown flights in three sparse density forest plots with the Deep Forest drone utilized an Ouster OS0-32 Rev.
-5 laser scanner (Ouster, San Francisco, CA, USA) and the Hovermap drone equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA), resulting in an averaged RMSE of 1.5 cm (6.7 %) and 1.6 cm (6.9 %) and bias of -0.7 cm (3.0 %) and -0.1 cm (-0.6 %).
-However, without utilizing their bias removal method RMSE and bias were significantly higher.
-The highest accuracy DBH estimates with under-canopy LiDAR were presented by Hyyppä et al.
-(2020a) resulting in an RMSE of 0.6 cm (2.3%) and bias of 0.34 cm (1.3%) in a sparse forest and 1.1 cm (3.5%) and 0.12 cm (0.4%) in a medium density forest.
-Their system was also manually piloted and weighted approximately 10 kg, a Tarot 960 UAV equipped with Kaarta Stencil-1 (VLP-16) LiDAR (Kaarta, Pittsburgh, Pennsylvania, USA).
-Summarizing, the highest accuracy levels were obtained with the relatively heavy ULS systems, which were operated manually in sparse forest environments.
-Miniaturized lightweight sensors were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), and Kuželka and Surovỳ (2018), and Liang et al.
-(2024).
-Results from the SfM-MVS systems were comparable to each other but ULS provided slightly poorer accuracy.
-Only studies by Shimabuku, Konoshima, and Ota (2023) and Liang et al.
-(2024) had flights with a high level of automation.
-Considering the complexity of the environments, the forest by Liang et al.
-(2024) (1000 trees/ha) was the closest to that in this study (650 and 2000 trees/ha), while other studies were conducted in remarkably sparser forests.
-Based on the previous, it can be concluded that when considering the combination of accuracy, autonomy of flying, and environmental complexity, the proposed system introduced a new level of performance for under-canopy forest reconstruction.
-The results show the great promise of miniaturized camera SfM-MVS in autonomous under-canopy forest reconstruction.
+The autonomous system utilized in Cheng et al.
+(2024) weighed 4.2 kg and provided an RMSE of 3.8 cm (12.8 %) with the data collected in autonomous flight mode.
+The same LiDAR-based drone platform was used in studies by Prabhu et al.
+(2024) and He et al.
+(2025), but based on the articles, it seems that data collecting flights were performed in manual flight mode in these studies.
+Prabhu et al.
+(2024) reported RMSE of 1.67 cm (7.42 %) with LiDAR data, and He et al.
+(2025) reported median relative error of 11.5 % with photogrammetric point clouds from 360° camera data.
+In summary, the highest accuracy levels were obtained with the relatively heavy ULS systems operated in sparse forest environments or in environments with few reference trees.
+Miniaturized lightweight sensors were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al.
+(2024), Cheng et al.
+(2024), He et al.
+(2025), and Liang et al.
+(2024), though the LiDAR-based systems (Prabhu et al.
+2024; Cheng et al.
+2024; He et al.
+2025; Liang et al.
+2024) were still considerably heavier and larger than the camera-based drone systems.
+From these, only studies by Shimabuku, Konoshima, and Ota (2023), Cheng et al.
+(2024), and Liang et al.
+(2024) reported flights with a high level of automation, and LiDAR-based flights in Cheng et al.
+(2024) and Liang et al.
+(2024) were the only ones where the systems were able to fly without commands from the pilot during the flight.
+Considering the complexity of the environments, the density of the forest in Liang et al.
+(2024) (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha).
+The natural-like subtropical forest in Shimabuku, Konoshima, and Ota (2023) had seven trees in 5 m $\times$ 5 m area, leading to a density of 2000 trees/ha, but the test plot had minimal size.
+In Cheng et al.
+(2024), the forest density was not reported.
+The approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a similar level of estimation accuracy to most of the manually piloted UAV SfM-MVS systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al.
+2025).
+Compared to the studies with LiDAR-based autonomous UAV collection, the accuracy was higher than in Liang et al.
+(2024) and similar than in Cheng et al.
+(2024) even without any GCPs.
+However, the system utilized in Cheng et al.
+(2024) has yielded slightly better accuracy with manually piloted flights (Cheng et al.
+2024; Prabhu et al.
+2024), suggesting that their LiDAR-based system has the potential for higher estimation accuracy by optimizing the autonomous flight trajectories.
+Although the estimation accuracy obtained in this study was in line with most of the relevant studies even without any GCPs, the reconstruction quality (completeness and RMSE) can be further improved with GCPs.
+Based on the previous, it can be concluded that camera-based autonomous miniaturized drones hold the potential to become a practical tool in forest parameter estimation.The potential estimation accuracy with low-cost stereo-cameras is near to the quality of small LiDARs suitable for miniaturized drones, but the weight and the price of the sensor are considerably lower.
+This allows the building of smaller platforms capable of flying in denser forests, as well as reduces the cost of the system.
 5.3 Potential improvements and future work
-During the evaluation of the system, some areas for improvement were identified, which can help develop the method towards a tool in practical forest measurements.
-The reliability of the navigation could be improved by performing further research on potential sensors and methods for detecting the thin branches.
-Blindsides in the mapping could be tackled either by sensors with 360° FOV (e.g.
-Lidar, 360° camera or multi-camera system) or storing historical data to occupancy grid maps in addition to the currently visible area.
+The empirical evaluation of the system revealed areas for improvement, which can inform further development toward a practical tool for operational forest measurements.
+The reliability of navigation and the smoothness of the trajectory could be further improved by investigating alternative camera systems and image processing methods (e.g., image super-resolution) with a better capability of detecting thin branches.
+Blindsides in the obstacle avoidance map could be tackled either by sensors with 360° FOV (e.g., 360° camera or multi-camera system) or storing historical data to occupancy grid maps in addition to the currently visible area.
 The virtual floor and ceiling of Ego-planner-V2 (Zhou et al.
-2022) are also defined as static with respect to the initial altitude of the drone limiting the low-altitude flying to relatively flat areas.
-To comprehensively map uneven forest plots, the altitude limits should be modified to be dynamic and with respect to the local terrain level.
-The drift compensation is one of the important future development goals.
-Loop closure tests provided promising results and potential future tests could evaluate the approach by performing block flights.
+2022) are also defined as static with respect to the takeoff location of the drone, limiting the low-altitude flying to relatively flat areas.
+To comprehensively map uneven forest plots, the altitude limits should be modified to be dynamic and concerning the local terrain level.
+Drift compensation remains an important objective for future development.
+Initial loop closure tests yielded promising results, and future developments could assess the approach using block flights with a lawnmower pattern.
 In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate.
@@
 GVINS (Cao, Lu, and Shen 2022).
-Even though the dense forest canopy blocks the GNSS signals, during long-term operations the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forests.
+Even though the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forests.
 It is also notable that the rigorous photogrammetric post-processing effectively compensated for the systematic errors.
-While the real-time trajectory error was 50 cm, the post-processed 3D error was approximately 11 cm without any GCPs, this correlates with the trajectory error.
-Therefore, for applications relying on post-processing, drift is not expected to pose significant issues, especially when covering small areas such as forest test plots.
+While the real-time trajectory error was 50 cm, the post-processed 3D error was approximately 11 cm without any GCPs.Therefore, for applications relying on post-processing, drift is unlikely to present significant challenges.
 In this study, the performed flight paths were straight flights through the sample plots, and the DBH estimation was performed only for the tree trunks recorded during that path.
-For more comprehensive forest characterization, the trajectory should cover the whole test plot, and observe the trees from multiple sides.
+For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides.
+Reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the higher DBH estimation error obtained with the trees with large diameters.
 The global path covering the whole test plot could be implemented either with a simple waypoint generation algorithm as Liang et al.
@@
 Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage.
-While this study only demonstrated the potential of forest characteristics estimation through DBH estimation and tree detection, the scope of data collection is not limited to that.
-Having a second camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds.
-With a developed camera-based prototype, the data collection is not limited to point clouds, but also applications of image based mapping and object detection can be developed on top of it.
-Potential applications could include e.g.
-classifying and counting species of trees and other vegetation as well as monitoring damages in the bark of the trees due to insect pests.
-The prototype opens also possibilities for research of real-time forest characteristics analysis.
-In this study, the point clouds were obtained with post-processing, but the utilized onboard camera Intel Realsense D435 can also produce 3D point clouds in real-time.
-Image-based object detection can also be performed in real-time e.g.
-with YOLO.
-The analysis can be performed either with the onboard computer or with a remote computer or cloud service over the Internet connection.
-6 Acknowledgements
-This research was funded by the Research Council of Finland within projects “Learning techniques for autonomous drone based hyperspectral analysis of forest vegetation” (decision no.
+While this study focused on demonstrating the potential for forest characteristics estimation through DBH estimation and tree detection, the scope of data collection extends well beyond these applications.
+Including an additional lightweight camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds.
+With the developed camera-based prototype, data collection extends beyond point clouds to encompass applications such as image-based mapping and object detection.
+This opens up a variety of potential uses, including classifying and counting tree species and other vegetation, as well as monitoring bark damages caused by insect pests.
+The prototype also opens possibilities for research of real-time forest characteristics analysis.
+In this study, the point clouds were obtained through post-processing.
+However, the onboard Intel Realsense D435 camera is also capable of generating 3D point clouds in real time.
+Additionally, image-based object detection can be performed in real-time, e.g., with YOLO (Redmon et al.
+2016).
+While only computationally light analysis can be conducted on the onboard computer, real-time analysis with higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.
+Acknowledgements
+This research was funded by the Research Council of Finland within projects ’Learning techniques for autonomous drone based hyperspectral analysis of forest vegetation’ (decision no.
 357380), Fireman (decision no.
@@
 357908).
+Declaration of Competing Interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
+Data Availability Statement
+Part of the data recorded during the field tests is available from the corresponding author upon reasonable request.

### v3

@@
 While operations above the forest canopy are already highly automated, flying inside forests remains challenging, primarily relying on manual piloting.
-Inside dense forests, reliance on the Global Navigation Satellite System (GNSS) for localization is not feasible.
-Additionally, the drone must autonomously adjust its flight path to avoid collisions.
+In dense forests, relying on the Global Navigation Satellite System (GNSS) for localization is not feasible.
+In addition, the drone must autonomously adjust its flight path to avoid collisions.
 Recently, advancements in robotics have enabled autonomous drone flights in GNSS-denied obstacle-rich areas.
-In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open-source methods and validating its performance for data collection inside forests.
+In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open source methods and validating its performance for data collection inside forests.
 Specifically, the study focused on camera-based autonomous flight under the forest canopy and photogrammetric post-processing of the data collected with the low-cost onboard stereo camera.
-The autonomous flight capability of the prototype was evaluated through multiple test flights at boreal forests.
+The autonomous flight capability of the prototype was evaluated through multiple test flights in boreal forests.
 The tree parameter estimation capability was studied by performing diameter at breast height (DBH) estimation.
-The prototype successfully carried out flights in selected challenging forest environments, and the experiments showed excellent performance in forest 3D modeling with a miniaturized stereoscopic photogrammetric system.
-The DBH estimation achieved a root mean square error (RMSE) of 3.33 cm (12.79 %) across all trees.
-For trees with a DBH less than 30 cm, the RMSE was 1.16 cm (5.74 %).
+The prototype successfully carried out flights in selected challenging forest environments, and the experiments showed promising performance in forest 3D modeling with a miniaturized stereoscopic photogrammetric system.
+The DBH estimation achieved a root mean square error (RMSE) of 3.33 - 3.97 cm (10.69 - 12.98 %) across all trees.
+For trees with a DBH less than 30 cm, the RMSE was 1.16 - 2.56 cm (5.74 - 12.47 %).
 The results provide valuable insights into autonomous under-canopy forest mapping and highlight the critical next steps for advancing lightweight robotic drone systems for mapping complex forest environments.
@@
 Robotic; drone; forest; photogrammetry; DBH
-articletype: ARTICLE PREPRINT - Word Count 10048
+articletype: ARTICLE PREPRINT
 1 Introduction
-Over the last decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information.
+In recent decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information.
 Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al.
@@
 Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
-Hence, the use of drones in forestry research has rapidly increased, and they offer a practical solution for forest data collection above the forest canopy (Puliti et al.
+Hence, the use of drones in forest research has increased rapidly and offers a practical solution to collect forest data above the forest canopy (Puliti et al.
 2015; Tuominen et al.
@@
 However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable.
-In dense forests, GNSS localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al.
+In dense forests, GNSS-based localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al.
 2010).
-Additionally, navigating inside forests requires the capability to autonomously avoid obstacles such as trees and bushes along the flight path.
+Furthermore, navigating within forests requires the ability to autonomously avoid obstacles such as trees and bushes along the flight path.
 Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., Hyyppä et al.
@@
 2024).
-Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in GNSS-denied, cluttered environments.
+Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in cluttered GNSS-denied environments.
 Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., Liu et al.
@@
 However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment.
-Most studies assess the effects of varying object density only in simulators, with only a couple of the methods tested through multiple test flights inside real forests.
+Most studies assess the effects of varying object density only in simulators, with only a couple of methods tested through multiple test flights inside real forests.
 Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation.
@@
 2020).
-In the case of drone development, where optimizing the weight and battery consumption play a key role, these factors lead to bigger and heavier drone platforms and flight batteries requiring more space between the obstacles to fly safely inside the forest.
-Moreover, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost.
+In the case of drone development, where optimizing the weight and battery consumption plays a key role, these factors lead to larger and heavier drone platforms and flight batteries that require more space between obstacles to fly safely inside the forest.
+Furthermore, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost.
 The data collected with the LiDARs is also limited to point clouds only, whereas the images collected with cameras can also be used for image-based analysis.
 Hence, although some LiDAR-based robotic drone systems have been successfully applied to under-canopy mapping of forest parameters, the potential of smaller and cheaper camera-based systems remains largely unexplored.
-Furthermore, the robotic drone systems proposed in the literature are typically not evaluated through multiple flight tests in different forest environments leaving the reliability and suitability to diverse forest environments as an open question.
+Furthermore, the robotic drone systems proposed in the literature are typically not evaluated through multiple flight tests in different forest environments, leaving the reliability and suitability to diverse forest environments as an open question.
 Therefore, in this study, a prototype of a small camera-based robotic drone system that can fly autonomously in challenging GNSS-denied forests was built and developed using custom hardware and state-of-the-art robotic methods.
@@
 What level of autonomous flight reliability can be achieved in challenging boreal forests with current methods for camera-based robotic drone navigation?
-What level of DBH estimation accuracy can be achieved with the low-cost onboard stereo-camera data recorded during the flight?
+What level of DBH estimation accuracy can be achieved with the low-cost onboard stereo camera data recorded during the flight?
 What are the remaining bottlenecks to be tackled in follow-up studies in order to achieve large-scale autonomous photogrammetric forest data collecting with miniaturized under-canopy robotic drone systems?
@@
 The reliability of the system was comprehensively evaluated through multiple test flights in boreal forests with varying tree densities and difficulty levels.
-  • Combining autonomous flight capabilities to under-canopy photogrammetric forest reconstruction by processing the data from a low-cost onboard stereo-camera to 3D point clouds.
+  • Combining autonomous flight capabilities to under-canopy photogrammetric forest reconstruction by processing the data from a low-cost onboard stereo camera to 3D point clouds.
   • Validation of autonomous flight data collection based on stereo cameras for the estimation of forest parameters through DBH estimation, including an evaluation of the impact of various factors on the accuracy of the estimate.
@@
 The rest of this study is structured as follows.
-Section 2 reviews the latest studies that estimate DBH using drones flying inside forests and recent open-source solutions for autonomous drone flights inside forests.
+Section 2 reviews the latest studies that estimate DBH using drones flying inside forests and recent open source solutions for autonomous drone flights inside forests.
 Section 3 describes the algorithms, hardware, experiments, and methods utilized in this study.
 Experimental results are presented in Section 4.
-Section 5 discusses the results, and potential improvements, and concludes the study.
+Section 5 discusses the results and potential improvements and concludes the study.
 2 Related work
 2.1 DBH estimation from under-canopy drone data
-In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total stem volume either alone or combined with the three height (Laasasenaho 1982).
-Although some of the relevant studies have also estimated tree volume by flying the flight path multiple times with different altitudes, with rotating sensors, or with Lidars with high vertical FOV, only DBH estimation accuracies and methods are included in this brief literature review.
-In the existing literature, LiDAR-based methods have been studied extensively for the purpose of efficiently and accurately estimating various tree attributes, such as DBH, tree height, and stem volume.
+In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total volume of the stem alone or in combination with the three heights (Laasasenaho 1982).
+In the existing literature, LiDAR-based methods have been extensively studied to efficiently and accurately estimate various tree attributes, such as DBH, tree height, and stem volume.
 To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al.
@@
 (2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy.
-The utilized custom-built drone system included Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan).
-They detected 12 trees in the vicinity of the drone trajectory and reported an RMSE of 25.1 % for the DBH estimates of the detected trees.
+The utilized custom-built drone system included a Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan).
 Hyyppä et al.
-(2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone.
-They demonstrated DBH estimates with an RMSE of 0.69-0.92 cm (2.2 %-3.1 %) and a bias of 0.29-0.31 cm (1.0 %-1.1 %) on two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha.
+(2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone in two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha.
 Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory.
 In a follow-up study, Hyyppä et al.
-(2020c) compared various mobile laser-scanning methods, including backpack, handheld, Under-Canopy UAV, and Above-Canopy UAV, for forest data collection.
-The same manually piloted under-canopy drone system was utilized in this study in the same forest plots as in Hyyppä et al.
+(2020c) compared various mobile laser scanning methods, including backpack, handheld, Under-Canopy drone, and Above-Canopy drone, for forest data collection.
+In this study, the same manually piloted under-canopy drone system was utilized in the same forest plots as in Hyyppä et al.
 (2020a).
-In this study, the manually piloted under-canopy drone system yielded an RMSE of 0.6-1.1 cm (2.3 %-3.5 %) and a bias of 0.34-0.12 cm (1.3 %-0.4 %).
 Wang et al.
-(2021) combined data from above- and under-canopy flights of a manually-piloted drone, resulting in an RMSE of 7.95 cm (33.35 %) and a bias of 3.72 cm (15.61 %) for the DBH.
+(2021) combined data from above and under canopy flights of a manually piloted drone equipped with a Riegl miniVUX laser scanner (RIEGL Laser Measurement Systems GmbH, Horn, Austria).
+The study area was a boreal forest plot with an approximate tree density of 200 trees/ha.
 Muhojoki et al.
-(2024) compared the accuracy of the estimated tree attributes from several LiDAR-based mobile systems, including two under-canopy drone systems, a system by Hovermap (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a system by Deep Forestry (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev.
+(2024) compared the accuracy of the estimated tree attributes of several LiDAR-based mobile systems, including two commercial under-canopy drone systems, a Hovermap system (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a Deep Forestry system (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev.
 5 laser scanner (Ouster, San Francisco, CA, USA).
-Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between the trees was at least 3 m.
+The Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between trees was at least 3 m.
 Both drone systems were used to collect point clouds under the canopy of three sparse forest plots.
-The RMSE and bias of DBH estimates for Deep Forestry and Hovermap drones, with bias removal method introduced by Muhojoki et al.
-(2024), averaged RMSE of 1.5 cm (6.7 %) and bias of -0.7 cm (3.0 %) and RMSE of 1.6 cm (6.9 %) and bias of -0.1 cm (-0.6 %), respectively.
-Without using the bias removal method, RMSE and bias for both drones were significantly higher.
-In addition to the manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation.
+In addition to manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation.
 Liang et al.
 (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha.
-To enable autonomous flight, a predefined global path was planned, to which the drone system performed real-time adjustments during the flight according to real-time environmental conditions.
-When the flight path covered most of the test plot, they achieved an RMSE of 5.13 cm (22.01 %) and bias of 3.40 cm (14.62 %) for the DBH of individual trees.
+To enable autonomous flight, a predefined global path was planned, in which the drone system made real-time adjustments during the flight according to real-time environmental conditions.
 Liu et al.
@@
 2020) algorithm to calculate the number of trees in the explored area.
-Although in this study the autonomous flight was demonstrated without the SLOAM algorithm and the tree number calculation was only demonstrated with one manually-piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation.
+Although in this study autonomous flight was demonstrated without the SLOAM algorithm and tree number calculation was only demonstrated with one manually piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation.
 Prabhu et al.
-(2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data.
-The DBH values were estimated for 20 trees located in two different test plots with RMSE of 1.67 cm (7.42 %).
-However, the DBH data-collecting flights were performed with manual flight mode.
+(2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones, as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data.
+However, DBH data collection flights were performed in manual flight mode.
 Cheng et al.
-(2024) used the same drone platform to collect a dataset for LiDAR-based mapping of trees.
+(2024) used the same drone platform to collect a dataset for tree mapping based on LiDAR.
 In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas.
-Although the drone was manually piloted in one of these forests, the data collection at the other forest consisting of loblolly pine trees was performed in autonomous flight mode.
-They also estimated the DBH value for 77 trees from the LiDAR data collected during the autonomous flights achieving an RMSE of 3.8 cm (12.3 %) by offline afterprocessing and 9.8 cm (31.8 %) by online estimation with SLOAM.
-With manually piloted data collection in the second test forest consisting of 20 trees with varying species, the RMSEs were 2.0 cm (7.5 %) with offline processing and 7.5 cm (27.5 %) with online processing.
-Besides the under-canopy flying drones, LiDARs have also been integrated to above-canopy flying drones capable of direct stem diameter measurements, though typically with a significantly higher RMSE and a lower tree detection rate due to the larger measurement distance and the upper canopy obstructing the tree stems (Wieser et al.
-2017; Brede et al.
-2017; Kuželka, Slavík, and Surovỳ 2020; Vandendaele et al.
-2021; Feng et al.
-2022).
+Although the drone was manually piloted in one of these forests, data collection in the other forest consisting of loblolly pine trees was carried out in autonomous flight mode.
+They also estimated DBH values for the LiDAR data collected during both autonomous and manual flights.
 During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH.
-Even though several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al.
+Although several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al.
 (2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al.
@@
 2025).
-Kuželka and Surovỳ (2018) manually operated a drone system based on a DJI Phantom 4 Pro (DJI, Shenzhen, China) within two forest plots to produce low-noise photogrammetric point clouds, from which the DBH of individual trees could be estimated with an RMSE between 2.6 cm and 5.3 cm depending on the plot and the applied analysis method.
-The studied plots were sparse with only 270-290 trees/ha.
-Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro inside two forest plots with a stem density of approximately 1000 trees/ha to acquire high-density SfM point cloud data, from which the DBH of individual trees could be measured with an RMSE of 1.1-2.1 cm.
-Shimabuku, Konoshima, and Ota (2023) achieved an RMSE of 0.4-0.7 cm for DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China).
-The first plot containing 18 trees was clean containing a minimal amount of low-hanging branches and understory vegetation.
-The second test plot was denser and closer to a natural subtropical forest environment but contained only 7 trees.
-The drone system was able to detect obstacles in the horizontal direction but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
+Kuželka and Surovỳ (2018) evaluated two manually operated drone systems, a DJI Phantom 4 Pro (DJI, Shenzhen, China) and a DJI Mavic Pro (DJI, Shenzhen, China), and various analysis methods to estimate DBH from photogrammetric point clouds.
+The systems were evaluated in two sparse forest plots with tree densities of 270-290 trees/ha.
+Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro in two forest plots with stem densities of 885 trees/ha and 980 to acquire high-density SfM point cloud data and estimated the DBH of individual trees from the point clouds.
+Shimabuku, Konoshima, and Ota (2023) performed DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China).
+The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
 Recently, He et al.
-(2025) mounted an additional 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) to the same LiDAR-based autonomous drone platform used in Liu et al.
+(2025) mounted a 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) as an additional payload to the same LiDAR-based autonomous drone platform used in Liu et al.
 (2022) and post-processed the photogrammetric data to 3D point clouds.
-Whether the flights were performed in autonomous flight mode or with manually piloted mode was not mentioned in the study.
-Their study evaluated DBH for 43 trees with varying tree species and the reported median relative DBH error was 11.5 % estimated from the photogrammetric point clouds.
-A summary of the results obtained in the relevant studies on under-canopy forest inventory is presented in Table 1.
-In summary of the related work, most of the studies for under-canopy drone-based DBH estimation have been performed with manually piloted drones.
-Although a few studies have explored the potential of custom robotic drone systems capable of flying without commands from the pilot during the flight, all of these systems have been based on LiDARs.
-All of the studies based on photogrammetry and SfM point clouds have utilized commercial drone systems, which have required manual piloting during the data collection.
+However, in the study, it was not mentioned whether the flights were performed in autonomous flight mode or with manual pilot mode.
+The RMSEs of the DBH estimates in the relevant studies are presented in Table 1.
+From the studies where multiple analysis methods were evaluated, the lowest RMSEs are presented.
+The study by He et al.
+(2025) was omitted from the table, as the study reported only the median relative error (11.5 %) of the accuracy of the DBH estimation instead of the RMSE.
+Although some of the relevant studies also estimated tree volumes by flying the flight path several times with different altitudes, with rotating sensors, or with LiDARs with high vertical FOV, only DBH estimation accuracies are included in the table.
 Table 1: Accuracy of the stem diameter breast height (DBH) estimations in other studies.
@@
 Krisanski et al. (2020) Drone SfM-MVS 29 4.1 - - -
+In summary of the related work, most of the studies for under-canopy drone-based DBH estimation have been conducted with manually piloted drones.
+Although a few studies have explored the potential of custom robotic drone systems capable of flying without commands from the pilot during the flight, all of these systems have been based on LiDARs.
+All of the studies based on photogrammetry and SfM point clouds have utilized commercial drone systems, which have required manual piloting during data collection.
 The highest estimation accuracies have been obtained with large systems unable to fly in dense forests or with test plots containing very few reference trees, but miniaturized systems have still yielded promising results.
-Notably, all studies utilizing SfM point clouds have been performed in environments that differ significantly from boreal forests.
+Furthermore, all studies utilizing SfM point clouds have been performed in environments that differ significantly from boreal forests.
 2.2 Autonomous flying inside forests
 Even though only a few studies have yet utilized robotic under-canopy drones in DBH estimation, drone autonomy itself has been an actively researched topic in the robotics community during the past few years.
-Recent advances in mobile robotics and the enhanced computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in GNSS-denied cluttered environments.
-Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments.
-Most importantly, the authors of these methods have also shared their source code to be openly available for the research community to encourage more development in the field.
-While the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al.
+Recent advances in mobile robotics and the improved computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in cluttered GNSS-denied environments.
+Although technology is still in the early stages of development, some recent studies have proposed various methods that have been successfully tested in forest environments.
+Most importantly, the authors of these methods have also shared their source codes to make them openly available to the research community to encourage more development in the field.
+Although the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al.
 2022; Prabhu et al.
@@
 2021; Nguyen et al.
-2023) based on the sensor utilized for the sensing of the obstacles in the flight path.
+2023) based on the sensor utilized for the detection of obstacles in the flight path.
 Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system.
@@
 (2022) and Ahmad et al.
-(2022) concentrated a swarm of drones instead of autonomous flying of an individual drone.
-However, most of these methods are typically validated by a single successful autonomous test flight in a specific forest environment.
+(2022) concentrated a swarm of drones instead of the autonomous flying of an individual drone.
+However, most of these methods were validated by a single successful autonomous test flight in a specific forest environment.
 Only four of the articles reported multiple test flights inside real forests (Loquercio et al.
@@
 2024) were the only methods tested at least in two different forest types.
-Moreover, since methods by (Loquercio et al.
+Furthermore, since the methods by (Loquercio et al.
 2021), (Ren et al.
-2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in those studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation.
+2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in these studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation.
 The autonomous trajectory planning method by Prabhu et al.
-(2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data collecting flights in denser forests were manually piloted.
-Therefore, open questions remain regarding the applicability and reliability of these algorithms in diverse forest environments, especially in challenging boreal forests.
+(2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data-collecting flights in denser forests were manually piloted.
+Therefore, open questions remain regarding the applicability and reliability of these algorithms in diverse forest environments, especially challenging boreal forests.
 3 Materials and methods
 3.1 Autonomous navigation algorithms
-In this study, the objective was to develop a prototype lightweight camera-based drone capable of autonomously avoiding obstacles under the forest canopy and operating without GNSS.
-The drone was required to fly autonomously from the takeoff location to a specified goal point while recording stereo-camera data during the flight.
+In this study, the objective was to develop a prototype lightweight camera-based drone capable of avoiding obstacles autonomously under the forest canopy and operating without GNSS.
+The drone was required to fly autonomously from the takeoff location to a specified goal point while recording stereo camera data during the flight.
 This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
-Among the open-source solutions presented in Section 2.2, the solution by Zhou et al.
-(2022) was chosen as the base of the prototype.
-The solution was the only open-sourced camera-based solution proven to work with only one stereo camera making it possible to build a smaller and lighter platform.
-Additionally, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, even though it had a highly different structure than the typical boreal forests.
+Among the open source solutions presented in Section 2.2, the solution by Zhou et al.
+(2022) was chosen as the basis of the prototype.
+The solution was the only open source camera-based solution that was proven to work with only one stereo camera, making it possible to build a smaller and lighter platform.
+Furthermore, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, although it had a highly different structure from typical boreal forests.
 The main details of the algorithms used in the chosen solution are briefly described in the following, and the complete system architecture of the prototype is presented in Figure 1.
@@
 The source code of the planner algorithm by Zhou et al.
-(2022) was named in Github as EGO-Planner-v2 (ZJU FAST Lab Team 2022a).
+(2022) was named on GitHub as EGO-Planner-v2 (ZJU FAST Lab Team 2022a).
 Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments.
@@
 The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985).
-The mapping module uses depth images as an input, and the system adopts the fixed-sized circular buffers proposed by Usenko et al.
-(2017) for maintaining the local map.
+The mapping module uses depth images as input and the system adopts the fixed-size circular buffers proposed by Usenko et al.
+(2017) to maintain the local map.
 The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively.
 The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
-In the trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al.
+In trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al.
 2022) trajectory representation designed for differentially flat systems, such as quadrotors (Mellinger and Kumar 2011).
-Planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (Zhou et al.
+The planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (Zhou et al.
 2022).
 The trajectories are optimized iteratively by minimizing the weighted sum of metrics defined for trajectory smoothness and flying time.
-The feasibility of the planned trajectory is forced by deforming the shape to avoid obstacles and by restricting the magnitudes of trajectory velocity, acceleration, and jerk to fulfill the dynamical constraints of the drone.
-For computational efficiency, the trajectory constraints are enforced via integrals of penalty functions with large penalty weights, and the integrals are evaluated by a finite sum of equally spaced samples along the timeline.
+The feasibility of the planned trajectory is forced by deforming the shape to avoid obstacles and by restricting the magnitudes of the trajectory velocity, acceleration, and jerk to meet the dynamical constraints of the drone.
+For computational efficiency, trajectory constraints are enforced through integrals of penalty functions with large penalty weights, and the integrals are evaluated by a finite sum of equally spaced samples along the timeline.
 In a simplified form, the optimization of time-dependent objectives $J$ in one piece of the trajectory can be represented by
@@
 The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al.
-2010) to DBoW2 (Galvez-López and Tardos 2012) bag-of-word database.
+2010) to a DBoW2 (Galvez-López and Tardos 2012) bag-of-word database.
 After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses.
-In addition to sliding window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses.
+In addition to sliding-window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses.
 Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
@@
 The parameters used in the configuration of VINS-Fusion were chosen so that real-time performance was guaranteed.
-In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion causing instability to the estimates, so the IMU noise values were increased more.
+In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion, causing instability to the estimates, so the IMU noise values were increased further.
 All VINS-Fusion configuration parameters are listed in Table 2.
@@
 The original EGO-Planner-v2 article (Zhou et al.
-2022) did not specify the utilized method for trajectory tracking.
-In this study, an open-source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used.
+2022) did not specify the method for trajectory tracking.
+In this study, an open source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used.
 The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints.
@@
 The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
-RealSense D435 is based on active stereoscopy, and the camera has an infrared laser projector emitting dots that can be used for triangulation by the stereo camera.
-This can improve the triangulation in low-light conditions and to homogeneous surfaces that contain no features.
+The RealSense D435 onboard camera was configured to the maximum image resolution ($1280\times 720$) and the update frequency (30 Hz).
+The exposure value was set to use the built-in automatic exposure control.
+The stereo images for VIO were downsampled at a resolution of $640\times 360$ to ensure real-time performance.
+Stereoscopic RealSense D435 has an infrared laser projector emitting dots, improving the triangulation in low light conditions and for homogeneous surfaces.
 However, the dots of the laser emitter cause problems for VIO since they are moving with the camera.
-In the previous set-up described by (Karjalainen et al.
-2023), a modified version of the ROS-RealSense driver by Zhou et al.
-(2021b) was used.
-The modified driver turns the laser emitter of the camera constantly on and off, making it possible to produce high-quality depth images in every other frame and stereo images without the laser dots in every other frame.
-However, this solution comes with the drawback of halving the output frame rate of the camera.
-In the first test flights described by (Karjalainen et al.
-2023), difficulties were encountered in detecting small branches with a depth camera resolution of $640\times 480$, leading to the decision to increase the depth camera resolution to a maximum value of $1280\times 720$.
-The stereo images for VIO were downsampled at a resolution of $640\times 360$ to ensure real-time performance.
-As the maximum frame rate of D435 at a resolution of $1280\times 720$ is 30 Hz, the modified driver version would have resulted in a low frame rate of 15 Hz.
-To ensure the VIO performance in quick turns, the frame rate of 30 Hz was preferred, and hence, the IR emitter was disabled, making the camera function as a passive IR stereo camera.
-The choice can be further justified with the fact that the usefulness of the IR emitter in forest conditions is questionable since sunlight contains infrared light that can interfere with the emitted artificial features(Kuan, Ee, and Wei 2019), and the forests typically do not contain large homogeneous surfaces.
-The Realsense D435 onboard camera was configured to use the built-in automatic exposure control.
-The camera-IMU system was calibrated with an open-source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013).
-In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were calculated from recorded data where Aprilgrid (Olson 2011) was used as a calibration target.
+Therefore, in this study, the infrared emitter was disabled.
+Since sunlight contains infrared light, which interferes with the artificial features emitted (Kuan, Ee, and Wei 2019), and forests typically do not contain homogeneous surfaces, disabling the emitter should not affect the camera performance in outdoor conditions.
+The camera-IMU system was calibrated with an open source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013).
+In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were estimated from a recorded data sequence, where Aprilgrid (Olson 2011) was used as a calibration target.
 The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (ETH Zürich Autonomous Systems Lab 2023).
@@
 Petersburg, Russia) were placed within the forest plots used in the flight tests.
-The targets had a diameter of 24.7 cm and the middle dots had a size of 7 cm.
-Relative distances between the GCPs were measured with Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020).
-Figure 5 shows GCPs in ’Evo-difficult’ during the test flights.
+The targets had a diameter of 24.7 cm and the middle dots had a diameter of 7 cm.
+Relative distances between the GCPs were measured with a Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020).
+Figure 5 shows the GCPs in ’Evo-difficult’ during the test flights.
 Figure 5: GCPs for reference trajectory generation in ’Evo-difficult’ forest.
@@
 GCPs were not used in the third test forest, ’Paloheinä’.
-The experiment there aimed to test the loop detection performance of VINS-Fusion in a homogeneous forest environment without any artificial targets.
-The data were collected by walking with the drone.
+The experiment there aimed to test the loop detection performance of VINS-Fusion in a homogeneous forest environment without artificial targets.
+The data sequences were collected by walking with the drone.
 The estimated average tree density in the area was approximately 2000 trees/ha.
@@
 Test flights in Evo were conducted on 18th October 2023.
-Multiple test flights were performed to evaluate the forest-type effects on the performance of autonomous navigation.
+Multiple test flights were conducted to evaluate forest-type effects on autonomous navigation performance.
 In total, seven test flights were conducted in the ’Evo-medium’ test area and nine test flights in the ’Evo-difficult’ test area.
@@
 Stereo images and IMU data were collected during two walks inside the forest while carrying the drone by hand.
-The paths of the walks were approximately 340 m and 240 m long, each containing multiple loops where the same scene was observed multiple times.
-The data collection took place during winter, with the forest covered in snow.
+The paths of the walks were approximately 340 m long and 240 m long, each containing multiple loops where the same scene was observed multiple times.
+Data collection was carried out during winter, with the forest covered in snow.
 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
@@
 Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m.
-In the image orientation processing, the quality setting was set to high meaning that two times downsampled images were used.
+In the image orientation processing, the quality setting was set to high, meaning that two times downsampled images were used.
 The settings for the number of key points and tie points per image were 40,000 and 4,000, respectively.
-To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of ±$\pm 0.005$ m.
-Automatic outlier removal was performed using gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy.
+To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of $\pm 0.005$ m.
+Automatic outlier removal was performed using the gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy.
 Subsequently, a bundle adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP).
@@
 (2020a) was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
-The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments, and 5-s-long temporal segments.
+The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments and 5-s-long temporal segments.
 The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data.
@@
 The diameters were processed into a stem curve by automatically removing clear outliers, followed by fitting a smoothing spline to the remaining diameter estimates as a function of height.
-Finally, the DBH estimates were calculated by interpolating the stem curve at the height of 1.3 m above the ground.
+Finally, the DBH estimates were calculated by interpolating the stem curve at a height of 1.3 m above the ground.
 For a detailed description of the method, the reader is referred to Hyyppä et al.
@@
 The performance and reliability of the autonomous flight of the prototype were evaluated based on two factors: the reliability of the obstacle avoidance and the accuracy of the VIO localization.
-Obstacle avoidance reliability was evaluated based on the mission success.
+The reliability of obstacle avoidance was evaluated on the basis of mission success.
 A mission was considered successful if the drone reached the end of the flight without major collisions.
@@
 In the mathematical form, the error can be presented as
-$\text{ATE}_{\text{pos}}=\left(\frac{1}{N}\sum_{i=0}^{N-1}||\Delta\mathbf{p}_{i}||^{2}\right)^{\frac{1}{2}}$ (2)
+[…]
 where $\Delta\mathbf{p}_{i}$ is the error between the estimated position and reference trajectory position in a single state.
 In the first tests conducted in a simulator as described by (Karjalainen et al.
-2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to constant, compared to when the online estimation method (Qin and Shen 2018) was used.
+2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to a constant, compared to when the online estimation method (Qin and Shen 2018) was used.
 For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr.
@@
 For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests.
-The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of online estimation.
+The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of the online estimation.
 The loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets that could affect the system performance.
-Since there were no GCPs for forming the reference trajectories, only the ability to detect the loops inside dense natural forests was tested, and the accuracy of the loop-corrected trajectory was not evaluated.
+Since there were no GCPs to form the reference trajectories, only the ability to detect loops inside dense natural forests was tested and the accuracy of the loop-corrected trajectory was not evaluated.
 VINS-Fusion estimate with the loop detection algorithm was run for the collected data with the onboard computer of the drone, and the effect of a maximum number of features in pictures for loop detection capability was examined.
@@
 The accuracy of the DBH estimation from the drone point cloud data was evaluated using TLS data as a reference.
-The TLS scannings were conducted in April 2020, and tree height, DBH, and other parameters were estimated from the dataset.
+The TLS scans were collected in April 2020, and tree height, DBH, and other parameters were estimated from the dataset.
 The estimated error of the DBH measurement was less than 1 cm.
@@
 Relative Bias (Bias %) is determined by dividing the Bias by the mean of the reference DBH values and multiplying by 100.
-[…]
-[…]
-Additionally standard deviation was calculated with following equation:
+$\text{Relative RMSE (\%)}=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-d_{\text{reference},i}\right)^{2}}}{\bar{d}_{\text{reference}}}\times 100$ (5)
+$\text{Relative Bias (\%)}=\frac{\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-d_{\text{reference},i}\right)}{\bar{d}_{\text{reference}}}\times 100$ (6)
+Additionally, standard deviation was calculated with the following equation:
 $\text{SD}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-\bar{d}_{\text{reference}}\right)^{2}}$ (7)
@@
 Variability in tree growth, influenced by factors such as species, locations, health, and age, is acknowledged as a source of error.
-Considering that the tree growth was relatively constant in the test area, forming a major component of the bias in the estimates.
+Considering that the tree growth was relatively constant in the test area, it formed a major component of the bias in the estimates.
 Additionally, some reference trees fell between the measurements and were not included in the estimations.
@@
 Consistent with these expectations, the trajectories were smoother and the success rate was better in ’Evo-medium’ than in ’Evo-difficult’ test area.
-In ’Evo-medium’, all performed test flights were successful and five out of seven test flights were performed with smooth trajectories without any emergency stops during the flight.
+In ’Evo-medium’, all performed test flights were successful, and five out of seven test flights were performed with smooth trajectories without any emergency stops during the flight.
 In ’Evo-difficult’, one of nine test flights failed.
 None of the test flights in ’Evo-difficult’ was performed with a smooth trajectory.
-However, despite the emergency stops during the flight, the system was still able to continue the mission after the stops, leading to a successful mission in eight of the nine test flights.
+However, despite emergency stops during the flight, the system was still able to continue the mission after stops, leading to a successful mission on eight of the nine test flights.
 The reason for emergency stops in both environments was the late detection of thin and dry spruce branches.
-In the failed test flight in ’Evo-difficult’ the reason was also a collision to a similar thin low branch.
+In the failed test flight in ’Evo-difficult’, the reason was also a collision with a similar thin low branch.
 The test flight success rates in both test forests are summarized in Table 3. Figure 8 presents examples of performed trajectories and obtained point clouds in both test forests.
@@
 In ’Evo-difficult’, one of the flights from the first takeoff orientation was performed without GCPs, leading to a non-comparable reference trajectory, one of the flights had a data recording error during a flight, and in one flight, Metashape failed to retrieve the reference path.
-Figure 9 shows an example of the VINS-Fusion trajectory estimate compared to the reference in ’Evo-difficult’ while Table 4 summarizes the VINS-Fusion estimation errors across all flights.
+Figure 9 shows an example of the VINS-Fusion trajectory estimate compared to the reference in ’Evo-difficult’, while Table 4 summarizes the VINS-Fusion estimation errors across all flights.
 The average $\text{ATE}_{\text{pos}}$ was 0.50 m in ’Evo-medium’ and 0.34 m in ’Evo-difficult’ with standard deviations of 0.06 m and 0.07 m, respectively, without online $t_{d}$ estimation.
@@
 In the loop closure tests, VINS-fusion detected a loop closure between 20 image pairs in the shorter path and four image pairs in the longer path with the same feature tracking parameters (Table 2) that were used in the flight tests.
-Figure 11 shows examples of detected loop closure image pairs from both paths.
+Figure 11 shows examples of pairs of detected loop closure images from both paths.
 The number of detected loops can be increased by increasing the number of features per image in the VINS-Fusion configuration.
@@
 I and II followed almost the same trajectory, while III and IV had similar trajectories.
-Dataset I trajectory was the smoothest, and it did not include any sharp turns.
+The dataset I trajectory was the smoothest, and it did not include any sharp turns.
 The other three trajectories had one or more sharper turns.
@@
 The total number of different trees found was 29 out of 33 (87,88 %) from dataset I-IV.
-However, when all the flights were processed simultaneously (combined) that resulted in only $3$ completeness, where only 12 of 36 trees were found.
+However, when all flights were processed simultaneously (combined), that resulted in only 33.33 % completeness, where only 12 of 36 trees were found.
 When utilizing 0, 1, 2, and 3 GCPs in the dataset I, completeness values were 72.41 %, 68.97 %, 75.86 %, and 79.31 %, respectively (Table 7).
 Figure 12: Post-processed trajectories, reference trees and found trees in Dataset (a) I (b) II (c) III (d) IV
-In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the flight trajectory despite these trees being well-reconstructed in the point cloud.
-Similarly, the number of well-reconstructed trees but not detected by the algorithm was 4, 5, and 5, in datasets II, III, and IV, respectively.
+In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the flight trajectory, despite these trees being well reconstructed in the point cloud.
+Similarly, the number of trees well reconstructed but not detected by the algorithm was 4, 5, and 5, in datasets II, III, and IV, respectively.
 The three extra trees that were not identified in dataset I were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud.
@@
 The RMSE and bias values ranged from 3.93 to 4.37 cm (13.68 %-14.66 %) and 1.12 to 2.13 cm (3.89 %-7.55 %), respectively (Table 7).
-Tree size had an impact on the DBH estimation accuracy (Table 7).
-For small and medium-sized trees with a DBH smaller than 30 cm, DBH estimations achieved superior performance in all datasets, with RMSE and bias varying from 1.16 to 2.75 cm (5.74 %-12.47 %) and -0.01 to 0.62 cm (-0.06 %-3.14 %), respectively.
+Tree size had an impact on the accuracy of DBH estimation (Table 7).
+For small and medium-sized trees with a DBH smaller than 30 cm, DBH estimations achieved good performance in all datasets, with RMSE and bias varying from 1.16 to 2.75 cm (5.74 %-12.47 %) and -0.01 to 0.62 cm (-0.06 %-3.14 %), respectively.
 Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99 %-25.44 %) and -0.66 to 2.08 cm (-1.53 %-5.65 %), respectively.
@@
 Table 7: Accuracy of the stem diameter breast height (DBH) in ’Evo-medium’ test site.
-Dataset is split in 3 categories: All, DBH $<$ 30 cm and DBH $>$ 30 cm trees.
+Dataset is split in 3 categories: All, DBH << 30 cm and DBH >> 30 cm trees.
 Datasets: I: Flight 1, II: Flight 2, III: Flight 3, IV: Flight 4, Combined: Flights I-IV merged, and for Dataset I: I0GCP: with no GCPs, I1GCP: with 1 GCP, I2GCP: with 2 GCPs, I3GCP with 3 GCPs.
-Dataset Number of trees All $<$30 cm $>$30 cm RMSE (cm) All $<$30 cm $>$30 cm RMSE % All $<$30 cm $>$30 cm Bias (cm) All $<$30 cm $>$30 cm Bias % All $<$30 cm $>$30 cm Standard Deviation All $<$30 cm $>$30 cm
+Dataset Number of trees All <<30 cm >>30 cm RMSE (cm) All <<30 cm >>30 cm RMSE % All <<30 cm >>30 cm Bias (cm) All <<30 cm >>30 cm Bias % All <<30 cm >>30 cm Standard Deviation All <<30 cm >>30 cm
 I 23 14 9 3.86 2.56 5.29 12.98 12.47 11.98 1.40 0.17 1.23 4.71 0.85 2.78 3.60 2.54 4.26
@@
 The obstacle avoidance success rate was 100 % in the ’Evo-medium’ forest and 87.5 % in the ’Evo-difficult’ forest.
-However, the smoothness of the trajectories suffered from the increasing forest density.
-While in ’Evo-medium’ 71.4 % of the flights had a smooth trajectory, in ’Evo-difficult’ at least one emergency stop was needed in all of the flights.
-However, apart from the one failed flight in ’Evo-difficult’, the system was able to do a successful trajectory re-planning and finish the mission after the emergency stops.
+However, the smoothness of the trajectories suffered from the increase in forest density.
+While in ’Evo-medium’ 71.4 % of the flights had a smooth trajectory, in ’Evo-difficult’, at least one emergency stop was needed in all of the flights.
+However, apart from the one failed flight in ’Evo-difficult’, the system was able to do a successful trajectory replanning and finish the mission after the emergency stops.
 Further consideration should be given to finding the optimal sensor for detecting thin and dry branches earlier.
-Nonetheless, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium-difficulty level boreal forests.
-The accuracy of the position estimate by VINS-Fusion was satisfying for short flights performed in this study, but in the longer flights, the error would start to accumulate, making the position errors high.
-However, the standard deviation between the errors between the test flights was small, and the VINS-Fusion systematically estimated the flight paths shorter than in reality (as can be seen from Figure 9).
+However, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium difficulty level boreal forests.
+The accuracy of the position estimate of VINS-Fusion was satisfactory for the short flights performed in this study, but during longer flights, the error would start to accumulate, resulting in high position errors.
+However, the standard deviation between the errors of the test flights was low and VINS-Fusion systematically estimated that the flight paths were shorter than in reality (as can be seen in Figure 9).
 This indicates that a notable part of the position error could be eliminated by improving the system calibration.
-The results indicate that the estimates by VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration.
-However, further development of the prototype should include implementing hardware-level synchronization between the camera and the IMU of the autopilot.
-The position error can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests show that VINS-Fusion could detect loop closures even in a homogeneous forest environment.
+The results indicate that the estimates of VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration.
+However, further development of the prototype should include implementation of hardware-level synchronization between the camera and the IMU of the autopilot.
+Position errors can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests show that VINS-Fusion can detect loop closures even in a homogeneous forest environment.
 Increasing the number of features tracked by VINS-Fusion increased the number of detected loops.
-However, increasing the number of features also increases the computational burden of the system, potentially disrupting the real-time estimation.
+However, increasing the number of features also increases the computational burden of the system, potentially disrupting real-time estimation.
 Finding the optimal value for the tracked features, as well as measuring the actual improvement of the loop closures to the position estimate, requires further studies.
 5.2 Accuracy of forest measurements
-The proposed system was assessed by comparing it to other relevant studies (Section 2.1, Table 1) considering the accuracy, system autonomy, weight, and complexity of the test environment.
-The RMSEs reported in the relevant studies varied from 3.2 % to 61.41 %.
-The highest accuracy UAV SfM-MVS results were presented by Shimabuku, Konoshima, and Ota (2023) with RMSEs of 0.4 cm (3.2 %) and 0.7 cm (3.6 %) in two test forest sites.
-However, their test sites included only 7 and 18 reference trees, and the semi-autonomous commercial drone system needed manual control to avoid horizontal obstacles.
-Kuželka and Surovỳ (2018), He et al.
-(2025), and Krisanski, Taskhiri, and Turner (2020) had more reference trees and achieved similar DBH accuracy compared to that in this study, but the flights were piloted manually.
-Notably, the SfM-MVS approaches involved commercial lightweight drone systems, except He et al.
-(2025) where 360° camera was mounted as an additional payload to heavier LiDAR-based custom drone.
-The most accurate under-canopy LiDAR-based DBH estimates were presented by Hyyppä et al.
-(2020a), achieving an RMSE of 0.6 cm (2.3 %) and bias of 0.34 cm (1.3 %) and 1.1 cm (3.5 %) and 0.12 cm (0.4 %) in two sparse forest plots.
-However, their system weighed around 10 kg and was piloted manually.
-Autonomously operating under-canopy LiDAR-based drones were utilized in the studies by Muhojoki et al.
-(2024), Liang et al.
-(2024), and Cheng et al.
-(2024).
-While the commercial Hovermap drone system utilized in Muhojoki et al.
-(2024) achieved the lowest RMSE of these, the system was able to fly only in sparse and clean forests due to its size.
-The automated under-canopy UAV Lidar system by Liang et al.
-(2024) weighted 5 kg and provided an RMSE of 4.48 cm (21.37 %) and bias of 5.01 (14.91 %) at best.
-The autonomous system utilized in Cheng et al.
-(2024) weighed 4.2 kg and provided an RMSE of 3.8 cm (12.8 %) with the data collected in autonomous flight mode.
-The same LiDAR-based drone platform was used in studies by Prabhu et al.
-(2024) and He et al.
-(2025), but based on the articles, it seems that data collecting flights were performed in manual flight mode in these studies.
-Prabhu et al.
-(2024) reported RMSE of 1.67 cm (7.42 %) with LiDAR data, and He et al.
-(2025) reported median relative error of 11.5 % with photogrammetric point clouds from 360° camera data.
-In summary, the highest accuracy levels were obtained with the relatively heavy ULS systems operated in sparse forest environments or in environments with few reference trees.
-Miniaturized lightweight sensors were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al.
+The proposed methodology and the accuracy of the DBH estimates obtained were evaluated by comparing them with other relevant studies (Section 2.1, Table 1) considering the precision, autonomy of the system, weight, and complexity of the test environment.
+The highest estimation accuracies in the literature have been obtained in studies using either considerably heavier and larger ULS systems Hyyppä et al.
+(2020a, c); Muhojoki et al.
+(2024) operated in sparse forest environments or in environments with few reference trees Shimabuku, Konoshima, and Ota (2023).
+The estimation results obtained in this study were close to those obtained in most other studies utilizing camera-based under-canopy drone systems (Kuželka and Surovỳ 2018; He et al.
+2025; Krisanski, Taskhiri, and Turner 2020).
+Furthermore, other studies based on SfM (Shimabuku, Konoshima, and Ota 2023; Kuželka and Surovỳ 2018; He et al.
+2025; Krisanski, Taskhiri, and Turner 2020) had more complex flight paths that allowed the full reconstruction most of the tree trunks, whereas in this study, the trunks were only partially reconstructed from image data collected from a single flight line.
+Considering the size of the systems, miniaturized lightweight systems were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al.
 (2024), Cheng et al.
@@
 (2025), and Liang et al.
-(2024), though the LiDAR-based systems (Prabhu et al.
+(2024), though the LiDAR-based custom drone systems (Prabhu et al.
 2024; Cheng et al.
@@
 2024) were still considerably heavier and larger than the camera-based drone systems.
-From these, only studies by Shimabuku, Konoshima, and Ota (2023), Cheng et al.
-(2024), and Liang et al.
-(2024) reported flights with a high level of automation, and LiDAR-based flights in Cheng et al.
+From these, LiDAR-based flights in Cheng et al.
 (2024) and Liang et al.
-(2024) were the only ones where the systems were able to fly without commands from the pilot during the flight.
+(2024) were the only ones where the systems were reported to fly without commands from the pilot during the flight.
 Considering the complexity of the environments, the density of the forest in Liang et al.
 (2024) (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha).
-The natural-like subtropical forest in Shimabuku, Konoshima, and Ota (2023) had seven trees in 5 m $\times$ 5 m area, leading to a density of 2000 trees/ha, but the test plot had minimal size.
 In Cheng et al.
-(2024), the forest density was not reported.
-The approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a similar level of estimation accuracy to most of the manually piloted UAV SfM-MVS systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al.
+(2024), the density of the forest was not reported.
+In summary, the approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a level of estimation close to the accuracy of most of the studies with SfM drone systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al.
 2025).
-Compared to the studies with LiDAR-based autonomous UAV collection, the accuracy was higher than in Liang et al.
-(2024) and similar than in Cheng et al.
+In particular, existing studies based on SfM of drones were using commercial drone systems piloted manually, except He et al.
+(2025) where the camera was added as an additional payload to a custom LiDAR-based drone platform, and the flight mode was not reported.
+Compared to studies with LiDAR-based miniaturized autonomous drone data collection, the accuracy of the estimation of DBH obtained in this study was higher than in Liang et al.
+(2024) and close to that in Cheng et al.
 (2024) even without any GCPs.
@@
 2024; Prabhu et al.
-2024), suggesting that their LiDAR-based system has the potential for higher estimation accuracy by optimizing the autonomous flight trajectories.
-Although the estimation accuracy obtained in this study was in line with most of the relevant studies even without any GCPs, the reconstruction quality (completeness and RMSE) can be further improved with GCPs.
-Based on the previous, it can be concluded that camera-based autonomous miniaturized drones hold the potential to become a practical tool in forest parameter estimation.The potential estimation accuracy with low-cost stereo-cameras is near to the quality of small LiDARs suitable for miniaturized drones, but the weight and the price of the sensor are considerably lower.
-This allows the building of smaller platforms capable of flying in denser forests, as well as reduces the cost of the system.
+2024), suggesting that their LiDAR-based system has the potential for higher accuracy of estimation by optimizing autonomous flight trajectories.
+Based on the results, the potential estimation accuracy with a low-cost stereo camera is considerably lower than that with heavy and expensive survey LiDARs but close to the quality of smaller LiDARs suitable for miniaturized drones capable of flying inside dense forests.
+However, the weight and price of the sensor are considerably lower, allowing the building of smaller platforms capable of flying in even denser forests, as well as reducing the cost of the system.
+The results show that GCPs can be used to improve the estimation quality (completeness and RMSE).
+However, the increase in estimation error was relatively low when GCPs were not used in postprocessing (RMSE from 12.98 % to 14.66 %) suggesting that the approach has the potential to work even without additional GCPs.
+Based on previous findings, it can be concluded that camera-based autonomous miniaturized drones have the potential to become a practical and cost-effective tool for forest parameter estimation in the future.
+However, further studies are required to increase the completeness of the reconstruction and the reconstruction coverage of the trees to decrease the RMSE of the estimation, especially for trees with large diameters.
 5.3 Potential improvements and future work
@@
 The reliability of navigation and the smoothness of the trajectory could be further improved by investigating alternative camera systems and image processing methods (e.g., image super-resolution) with a better capability of detecting thin branches.
-Blindsides in the obstacle avoidance map could be tackled either by sensors with 360° FOV (e.g., 360° camera or multi-camera system) or storing historical data to occupancy grid maps in addition to the currently visible area.
+Blind sides in the obstacle avoidance map could be tackled either by sensors with 360° FOV (e.g., 360° camera or multi-camera system) or by storing historical data in the occupancy grid map in addition to the current visible area.
 The virtual floor and ceiling of Ego-planner-V2 (Zhou et al.
-2022) are also defined as static with respect to the takeoff location of the drone, limiting the low-altitude flying to relatively flat areas.
-To comprehensively map uneven forest plots, the altitude limits should be modified to be dynamic and concerning the local terrain level.
-Drift compensation remains an important objective for future development.
+2022) are also defined as static with respect to the takeoff location of the drone, limiting low-altitude flying to relatively flat areas.
+To capture images at the DBH estimation altitude in uneven forest plots, the flight altitude limits should be modified to be dynamic with respect to the local terrain level.
+The drift compensation remains also an important objective for future development.
 Initial loop closure tests yielded promising results, and future developments could assess the approach using block flights with a lawnmower pattern.
 In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate.
-The potential methods for GNSS-Fusion include using the global fusion features in VINS-Fusion (Qin et al.
-2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g.
-GVINS (Cao, Lu, and Shen 2022).
-Even though the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forests.
+Potential methods for GNSS-Fusion include utilizing the global fusion features in VINS-Fusion (Qin et al.
+2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g., GVINS (Cao, Lu, and Shen 2022).
+Although the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forest.
 It is also notable that the rigorous photogrammetric post-processing effectively compensated for the systematic errors.
 While the real-time trajectory error was 50 cm, the post-processed 3D error was approximately 11 cm without any GCPs.Therefore, for applications relying on post-processing, drift is unlikely to present significant challenges.
-In this study, the performed flight paths were straight flights through the sample plots, and the DBH estimation was performed only for the tree trunks recorded during that path.
+In this study, the flight paths were straight flights through the sample plots, and the DBH values were estimated only for the tree trunks recorded during that path.
 For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides.
-Reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the higher DBH estimation error obtained with the trees with large diameters.
-The global path covering the whole test plot could be implemented either with a simple waypoint generation algorithm as Liang et al.
-(2024) or with more advanced exploration methods such as proposed by Zhou et al.
-(2021a) or Cieslewski, Kaufmann, and Scaramuzza (2017).
-Another way to achieve area coverage would be utilizing a swarm of drones as in the original article by Zhou et al.
+In the study by He et al.
+(2025), the DBH estimation error increased significantly when only part of the shape of the tree trunk was observed.
+This suggests that reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the DBH estimate errors obtained with the system utilized in this study.
+In particular, this is expected to improve the quality of the estimation for trees of large diameter that yielded a higher RMSE in this study than trees with a DBH of less than 30 cm.
+The global path covering the entire test plot could be implemented with a simple waypoint generation algorithm, such as Liang et al.
+(2024), or with more advanced exploration methods such as those proposed by Zhou et al.
+(2021) or Cieslewski, Kaufmann, and Scaramuzza (2017).
+Another way to achieve area coverage would be utilizing a swarm of drones, as in the original article by Zhou et al.
 (2022).
 Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage.
-While this study focused on demonstrating the potential for forest characteristics estimation through DBH estimation and tree detection, the scope of data collection extends well beyond these applications.
+In potential follow-up studies, the DBH estimation potential of the system could also be evaluated in alternative forest environments, for example, in sparser forest environments.
+Although this study focused on demonstrating the potential for forest characteristics estimation through DBH estimation and tree detection, the scope of data collection extends well beyond these applications.
 Including an additional lightweight camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds.
 With the developed camera-based prototype, data collection extends beyond point clouds to encompass applications such as image-based mapping and object detection.
-This opens up a variety of potential uses, including classifying and counting tree species and other vegetation, as well as monitoring bark damages caused by insect pests.
-The prototype also opens possibilities for research of real-time forest characteristics analysis.
+This opens up a variety of potential uses, including classifying and counting tree species and other vegetation, as well as monitoring bark damage caused by insect pests.
+The prototype also opens up possibilities for the study of real-time forest characteristics analysis.
 In this study, the point clouds were obtained through post-processing.
-However, the onboard Intel Realsense D435 camera is also capable of generating 3D point clouds in real time.
-Additionally, image-based object detection can be performed in real-time, e.g., with YOLO (Redmon et al.
+However, Intel RealSense D435 is also capable of generating 3D point clouds in real time.
+Furthermore, image-based object detection can be performed in real time, for example, with YOLO (Redmon et al.
 2016).
-While only computationally light analysis can be conducted on the onboard computer, real-time analysis with higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.
+While only computationally light analysis can be conducted on the onboard computer, real-time analysis with a higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.
 Acknowledgements

### v4

@@
 Robotic; drone; forest; photogrammetry; DBH
-articletype: ARTICLE PREPRINT
+articletype: This is an Accepted Manuscript of an article published by Taylor & Francis in International Journal of Remote Sensing on 03 Dec 2025.
+The published final Version of Record is openly available at: https://doi.org/10.1080/01431161.2025.2579803.
 1 Introduction
 In recent decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information.
-Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al.
-2007) and mobile solutions, e.g., handheld devices (Ryding et al.
-2015; Bauwens et al.
-2016; Marselis et al.
-2016; Balenović et al.
-2021), backpacks (Liang et al.
-2014b, 2018; Hyyppä et al.
-2020b), ground-based vehicles (Forsman, Holmgren, and Olofsson 2016; Liang et al.
-2018; Bienert et al.
-2018) or drones (Chisholm et al.
-2013; Brede et al.
-2017; Wieser et al.
-2017; Hyyppä et al.
-2020a).
-Similarly, the use of structure-from-motion (SfM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., Liang et al.
-2014a; Wallace et al.
-2016; Kuželka and Surovỳ 2018; Piermattei et al.
-2019).
+Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (bienert2007tree) and mobile solutions, e.g., handheld devices (ryding2015assessing; bauwens2016forest; marselis2016deriving; balenovic2021hand), backpacks (liang2014possibilities; liang2018situ; hyyppa2020accurate), ground-based vehicles (forsman2016tree; liang2018situ; bienert2018comparison) or drones (chisholm2013uav; brede2017comparing; wieser2017case; hyyppa2020under).
+Similarly, the use of structure-from-motion (SfM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., liang2014use; wallace2016assessment; kuvzelka2018mapping; piermattei2019terrestrial).
 Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
-Hence, the use of drones in forest research has increased rapidly and offers a practical solution to collect forest data above the forest canopy (Puliti et al.
-2015; Tuominen et al.
-2017).
+Hence, the use of drones in forest research has increased rapidly and offers a practical solution to collect forest data above the forest canopy (puliti2015forestUAS; tuominen2017hyperspectral).
 Modern commercial drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems.
 However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable.
-In dense forests, GNSS-based localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al.
-2010).
+In dense forests, GNSS-based localization becomes unreliable due to signal blockages and multipath effects caused by trees (gnss_multipath).
 Furthermore, navigating within forests requires the ability to autonomously avoid obstacles such as trees and bushes along the flight path.
-Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., Hyyppä et al.
-2020a; Liang et al.
-2019; Krisanski, Taskhiri, and Turner 2020; Prabhu et al.
-2024).
+Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., hyyppa2020under; liang2019forest; krisanski2020photodrone; prabhu2024treeFalcon).
 Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in cluttered GNSS-denied environments.
-Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., Liu et al.
-2022; Loquercio et al.
-2021; Zhou et al.
-2022).
+Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., LiuEtAl; loquercio2021learning; zhou2022swarm).
 However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment.
@@
 Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation.
-Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (Liu et al.
-2022; Liang et al.
-2024; Cheng et al.
-2024).
+Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (LiuEtAl; liang2024forest; cheng2024Treescope).
 Notably, all of these studies have utilized Ouster OS1 3D LiDAR (Ouster, San Francisco, CA, USA) as the main sensor.
-Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (Reddy Cenkeramaddi et al.
-2020).
+Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (sensors).
 In the case of drone development, where optimizing the weight and battery consumption plays a key role, these factors lead to larger and heavier drone platforms and flight batteries that require more space between obstacles to fly safely inside the forest.
@@
 2.1 DBH estimation from under-canopy drone data
-In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total volume of the stem alone or in combination with the three heights (Laasasenaho 1982).
+In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total volume of the stem alone or in combination with the three heights (laasasenaho1982taper).
 In the existing literature, LiDAR-based methods have been extensively studied to efficiently and accurately estimate various tree attributes, such as DBH, tree height, and stem volume.
-To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al.
-2015; Bauwens et al.
-2016; Cabo et al.
-2018; Chen et al.
-2019; Hyyppä et al.
-2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation.
-Chisholm et al.
-(2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy.
+To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (ryding2015assessing; bauwens2016forest; cabo2018comparing; chen2019applicability; hyyppa2020accurate; hyyppa2020under; hyyppa2022direct), though most studies have focused on relatively sparse forests with little understory vegetation.
+chisholm2013uav were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy.
 The utilized custom-built drone system included a Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan).
-Hyyppä et al.
-(2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone in two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha.
+hyyppa2020under collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone in two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha.
 Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory.
-In a follow-up study, Hyyppä et al.
-(2020c) compared various mobile laser scanning methods, including backpack, handheld, Under-Canopy drone, and Above-Canopy drone, for forest data collection.
-In this study, the same manually piloted under-canopy drone system was utilized in the same forest plots as in Hyyppä et al.
-(2020a).
-Wang et al.
-(2021) combined data from above and under canopy flights of a manually piloted drone equipped with a Riegl miniVUX laser scanner (RIEGL Laser Measurement Systems GmbH, Horn, Austria).
+In a follow-up study, hyyppa2020comparison compared various mobile laser scanning methods, including backpack, handheld, Under-Canopy drone, and Above-Canopy drone, for forest data collection.
+In this study, the same manually piloted under-canopy drone system was utilized in the same forest plots as in hyyppa2020under.
+wang2021seamless combined data from above and under canopy flights of a manually piloted drone equipped with a Riegl miniVUX laser scanner (RIEGL Laser Measurement Systems GmbH, Horn, Austria).
 The study area was a boreal forest plot with an approximate tree density of 200 trees/ha.
-Muhojoki et al.
-(2024) compared the accuracy of the estimated tree attributes of several LiDAR-based mobile systems, including two commercial under-canopy drone systems, a Hovermap system (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a Deep Forestry system (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev.
+muhojoki24BENCHMARKING compared the accuracy of the estimated tree attributes of several LiDAR-based mobile systems, including two commercial under-canopy drone systems, a Hovermap system (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a Deep Forestry system (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev.
 5 laser scanner (Ouster, San Francisco, CA, USA).
@@
 In addition to manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation.
-Liang et al.
-(2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha.
+liang2024forest applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha.
 To enable autonomous flight, a predefined global path was planned, in which the drone system made real-time adjustments during the flight according to real-time environmental conditions.
-Liu et al.
-(2022) employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (Chen et al.
-2020) algorithm to calculate the number of trees in the explored area.
+LiuEtAl employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (chen2020sloam) algorithm to calculate the number of trees in the explored area.
 Although in this study autonomous flight was demonstrated without the SLOAM algorithm and tree number calculation was only demonstrated with one manually piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation.
-Prabhu et al.
-(2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones, as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data.
+prabhu2024treeFalcon extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones, as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data.
 However, DBH data collection flights were performed in manual flight mode.
-Cheng et al.
-(2024) used the same drone platform to collect a dataset for tree mapping based on LiDAR.
+cheng2024Treescope used the same drone platform to collect a dataset for tree mapping based on LiDAR.
 In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas.
@@
 During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH.
-Although several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al.
-(2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al.
-(2018); Piermattei et al.
-(2019); Xu, Shen, and Cao (2023); Gao and Kan (2022)) and above-canopy photogrammetric systems (e.g., Wallace et al.
-(2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023; He et al.
-2025).
-Kuželka and Surovỳ (2018) evaluated two manually operated drone systems, a DJI Phantom 4 Pro (DJI, Shenzhen, China) and a DJI Mavic Pro (DJI, Shenzhen, China), and various analysis methods to estimate DBH from photogrammetric point clouds.
+Although several studies have investigated terrestrial photogrammetric systems (e.g., liang2014use; surovy2016accuracy; mokrovs2018high; piermattei2019terrestrial; xu2023sfm_vs_ls; gao2022automaticDBH) and above-canopy photogrammetric systems (e.g., wallace2016assessment; ye2019analysing) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (kuvzelka2018mapping; krisanski2020photodrone; shimabuku2023diameter; he2025estimating360).
+kuvzelka2018mapping evaluated two manually operated drone systems, a DJI Phantom 4 Pro (DJI, Shenzhen, China) and a DJI Mavic Pro (DJI, Shenzhen, China), and various analysis methods to estimate DBH from photogrammetric point clouds.
 The systems were evaluated in two sparse forest plots with tree densities of 270-290 trees/ha.
-Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro in two forest plots with stem densities of 885 trees/ha and 980 to acquire high-density SfM point cloud data and estimated the DBH of individual trees from the point clouds.
-Shimabuku, Konoshima, and Ota (2023) performed DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China).
+krisanski2020photodrone performed manual flights with a DJI Phantom 4 Pro in two forest plots with stem densities of 885 trees/ha and 980 to acquire high-density SfM point cloud data and estimated the DBH of individual trees from the point clouds.
+shimabuku2023diameter performed DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China).
 The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
-Recently, He et al.
-(2025) mounted a 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) as an additional payload to the same LiDAR-based autonomous drone platform used in Liu et al.
-(2022) and post-processed the photogrammetric data to 3D point clouds.
+Recently, he2025estimating360 mounted a 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) as an additional payload to the same LiDAR-based autonomous drone platform used in LiuEtAl and post-processed the photogrammetric data to 3D point clouds.
 However, in the study, it was not mentioned whether the flights were performed in autonomous flight mode or with manual pilot mode.
@@
 From the studies where multiple analysis methods were evaluated, the lowest RMSEs are presented.
-The study by He et al.
-(2025) was omitted from the table, as the study reported only the median relative error (11.5 %) of the accuracy of the DBH estimation instead of the RMSE.
+The study by he2025estimating360 was omitted from the table, as the study reported only the median relative error (11.5 %) of the accuracy of the DBH estimation instead of the RMSE.
 Although some of the relevant studies also estimated tree volumes by flying the flight path several times with different altitudes, with rotating sensors, or with LiDARs with high vertical FOV, only DBH estimation accuracies are included in the table.
@@
 *averaged values from three plots.
-The study by He et al.
-(2025) was omitted from this table since they reported only median errors instead of RMSE.
+The study by he2025estimating360 was omitted from this table since they reported only median errors instead of RMSE.
 Dataset Type Number of trees RMSE (cm) RMSE % Bias (cm) Bias %
@@
 Most importantly, the authors of these methods have also shared their source codes to make them openly available to the research community to encourage more development in the field.
-Although the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al.
-2022; Prabhu et al.
-2024; Ren et al.
-2022; Liu, Ren, and Zhang 2024; Ahmad et al.
-2022) and camera-based methods (Campos-Macías et al.
-2021; Zhou et al.
-2022; Loquercio et al.
-2021; Nguyen et al.
-2023) based on the sensor utilized for the detection of obstacles in the flight path.
+Although the methods utilize various sensor setups, they can be divided into LiDAR-based methods (LiuEtAl; prabhu2024treeFalcon; ren2022BubblePlanner; LiuIPC2024; Ahmad2022Bioinspired) and camera-based methods (CamposEtAl; zhou2022swarm; loquercio2021learning; nguyen2023Unceratainity) based on the sensor utilized for the detection of obstacles in the flight path.
 Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system.
-Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by Loquercio et al.
-(2021) and Nguyen et al.
-(2023) replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories.
-In addition, the methods proposed by Zhou et al.
-(2022) and Ahmad et al.
-(2022) concentrated a swarm of drones instead of the autonomous flying of an individual drone.
+Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by loquercio2021learning and nguyen2023Unceratainity replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories.
+In addition, the methods proposed by zhou2022swarm and Ahmad2022Bioinspired concentrated a swarm of drones instead of the autonomous flying of an individual drone.
 However, most of these methods were validated by a single successful autonomous test flight in a specific forest environment.
-Only four of the articles reported multiple test flights inside real forests (Loquercio et al.
-2021; Ren et al.
-2022; Liu, Ren, and Zhang 2024; Prabhu et al.
-2024).
-From these, (Loquercio et al.
-2021) and (Prabhu et al.
-2024) were the only methods tested at least in two different forest types.
-Furthermore, since the methods by (Loquercio et al.
-2021), (Ren et al.
-2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in these studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation.
-The autonomous trajectory planning method by Prabhu et al.
-(2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data-collecting flights in denser forests were manually piloted.
+Only four of the articles reported multiple test flights inside real forests (loquercio2021learning; ren2022BubblePlanner; LiuIPC2024; prabhu2024treeFalcon).
+From these, (loquercio2021learning) and (prabhu2024treeFalcon) were the only methods tested at least in two different forest types.
+Furthermore, since the methods by (loquercio2021learning), (ren2022BubblePlanner), and (LiuIPC2024) were all targeting high-speed flight, the test forests reported in these studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation.
+The autonomous trajectory planning method by prabhu2024treeFalcon was also tested only in environments where the space between the trees was larger than in typical forests, and the data-collecting flights in denser forests were manually piloted.
 Therefore, open questions remain regarding the applicability and reliability of these algorithms in diverse forest environments, especially challenging boreal forests.
@@
 This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
-Among the open source solutions presented in Section 2.2, the solution by Zhou et al.
-(2022) was chosen as the basis of the prototype.
+Among the open source solutions presented in Section 2.2, the solution by zhou2022swarm was chosen as the basis of the prototype.
 The solution was the only open source camera-based solution that was proven to work with only one stereo camera, making it possible to build a smaller and lighter platform.
@@
 Figure 1: The system architecture used in the data collection.
-In comparison to the original EGO-Planner-v2 article by Zhou et al.
-(2022), UWB-based Drift correction and Wireless trajectory broadcast network modules used for swarm navigation were omitted in this study.
+In comparison to the original EGO-Planner-v2 article by zhou2022swarm, UWB-based Drift correction and Wireless trajectory broadcast network modules used for swarm navigation were omitted in this study.
 Stereo images were downsampled for VIO to guarantee real-time estimation.
@@
 3.1.1 EGO-Planner-v2
-The source code of the planner algorithm by Zhou et al.
-(2022) was named on GitHub as EGO-Planner-v2 (ZJU FAST Lab Team 2022a).
+The source code of the planner algorithm by zhou2022swarm was named on GitHub as EGO-Planner-v2 (EGO-Planner-code).
 Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments.
-The planner runs on top of the Robot Operating System (ROS) (Quigley et al.
-2009).
-The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985).
-The mapping module uses depth images as input and the system adopts the fixed-size circular buffers proposed by Usenko et al.
-(2017) to maintain the local map.
+The planner runs on top of the Robot Operating System (ROS) (quigley2009ros).
+The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (moravec1985high).
+The mapping module uses depth images as input and the system adopts the fixed-size circular buffers proposed by usenko2017 to maintain the local map.
 The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively.
 The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
-In trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al.
-2022) trajectory representation designed for differentially flat systems, such as quadrotors (Mellinger and Kumar 2011).
-The planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (Zhou et al.
-2022).
+In trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (MINCO) trajectory representation designed for differentially flat systems, such as quadrotors (differentialFlatness).
+The planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (zhou2022swarm).
 The trajectories are optimized iteratively by minimizing the weighted sum of metrics defined for trajectory smoothness and flying time.
@@
 The trajectory planning is continuously performed within a defined local planning distance towards a user-defined global goal until the goal is reached.
-Since the safety constraints are modeled as integrals in the penalty functions, the feasibility of the trajectory needs to be verified with a post-check after planning (Zhou et al.
-2022).
+Since the safety constraints are modeled as integrals in the penalty functions, the feasibility of the trajectory needs to be verified with a post-check after planning (zhou2022swarm).
 Due to the non-static probabilistic map, the post-check after trajectory planning guarantees feasibility only at a certain moment.
@@
 3.1.2 VINS-Fusion
-The Visual-inertial odometry (VIO) algorithm used by Zhou et al.
-(2022) was VINS-Fusion (Qin, Li, and Shen 2018; Qin et al.
-2019b)(VINS stands for ’Visual-Inertial system’), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
-VINS-Fusion detects Shi-Tomasi corner features (Shi and Tomasi 1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (Lucas and Kanade 1981).
+The Visual-inertial odometry (VIO) algorithm used by zhou2022swarm was VINS-Fusion (qin2018vins; qin2019general)(VINS stands for ’Visual-Inertial system’), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
+VINS-Fusion detects Shi-Tomasi corner features (shiTomasi1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (lucas1981iterative).
 IMU data is used for preintegration, i.e., position, velocity, and orientation are integrated from IMU measurements between two frames.
@@
 VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate.
-The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al.
-2010) to a DBoW2 (Galvez-López and Tardos 2012) bag-of-word database.
+The loops are detected by storing additional corner features as BRIEF descriptors (BRIEF) to a DBoW2 (DBoW2) bag-of-word database.
 After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses.
@@
 Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
-VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (Qin and Shen 2018).
+VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (onlineTd).
 The method can be used to adjust the time offset $t_{d}$ between the camera and the IMU online.
-In the swarm operations of the original article (Zhou et al.
-2022), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion.
+In the swarm operations of the original article (zhou2022swarm), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion.
 However, with a single drone system in this study, that function is omitted.
@@
 3.1.3 Tracking controller
-The original EGO-Planner-v2 article (Zhou et al.
-2022) did not specify the method for trajectory tracking.
-In this study, an open source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used.
+The original EGO-Planner-v2 article (zhou2022swarm) did not specify the method for trajectory tracking.
+In this study, an open source tracking controller (px4ctrl) by Zhejiang University was used.
 The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints.
-The setpoints are forwarded to the PX4 flight controller software (PX4 Community 2023) responsible for the low-level attitude and angular rate controls.
+The setpoints are forwarded to the PX4 flight controller software (PX4code) responsible for the low-level attitude and angular rate controls.
 3.2 Drone hardware
@@
 The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the MAVROS protocol was used for communication between the onboard computer and the autopilot.
-As a camera, Intel RealSense D435 (Intel Realsense Development Team 2023)(Keselman et al.
-2017) was used.
+As a camera, Intel RealSense D435 (realsense)(realsense_article) was used.
 The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
@@
 Therefore, in this study, the infrared emitter was disabled.
-Since sunlight contains infrared light, which interferes with the artificial features emitted (Kuan, Ee, and Wei 2019), and forests typically do not contain homogeneous surfaces, disabling the emitter should not affect the camera performance in outdoor conditions.
-The camera-IMU system was calibrated with an open source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013).
-In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were estimated from a recorded data sequence, where Aprilgrid (Olson 2011) was used as a calibration target.
-The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (ETH Zürich Autonomous Systems Lab 2023).
+Since sunlight contains infrared light, which interferes with the artificial features emitted (IRoutdoor), and forests typically do not contain homogeneous surfaces, disabling the emitter should not affect the camera performance in outdoor conditions.
+The camera-IMU system was calibrated with an open source camera calibration toolbox, kalibr (kalibr3; kalibr2; kalibr1).
+In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were estimated from a recorded data sequence, where Aprilgrid (apriltag) was used as a calibration target.
+The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (AllanVarianceROS), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (kalibrWikiIMU).
 The computer, the camera, and the autopilot were attached to a 330 mm drone frame.
@@
 The flight tests of the system were performed in two test areas located in Evo, Finland (61.19°N, 25.11°E).
-The flight paths went through well-documented forest sample plots that have been widely used in previous studies, such as in Liang et al.
-(2018, 2019); Wang et al.
-(2019); Hyyppä et al.
-(2020a, c), and are classified into three complexity categories, i.e., ’easy’, ’medium’, and ’difficult’ based on, mainly tree density and amount of vegetation under the canopy in the plots.
+The flight paths went through well-documented forest sample plots that have been widely used in previous studies, such as in liang2018situ; liang2019forest; wang2019field; hyyppa2020under; hyyppa2020comparison, and are classified into three complexity categories, i.e., ’easy’, ’medium’, and ’difficult’ based on, mainly tree density and amount of vegetation under the canopy in the plots.
 The third test area, a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E) was used only to test the loop detection performance of VINS-Fusion.
@@
 Figure 4: View from the takeoff location in the ’Evo-difficult’ test forest.
-For reference purposes, circular GCPs provided by Agisoft Metashape (Agisoft LLC, St.
+For reference purposes, circular Ground Control Points (GCPs) provided by Agisoft Metashape (Agisoft LLC, St.
 Petersburg, Russia) were placed within the forest plots used in the flight tests.
 The targets had a diameter of 24.7 cm and the middle dots had a diameter of 7 cm.
-Relative distances between the GCPs were measured with a Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020).
+Relative distances between the GCPs were measured with a Leica Nova TS60 total station with a precision of approximately 2 mm (TS60Datasheet).
 Figure 5 shows the GCPs in ’Evo-difficult’ during the test flights.
@@
 Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St.
-Petersburg, Russia) (Metashape Development Team 2021) following similar processing steps as introduced by Viljanen et al.
-(2018).
+Petersburg, Russia) (Metashape) following similar processing steps as introduced by Viljanen2018.
 Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m.
@@
 3.6 Tree parameter estimation
-In this study, trees were detected, and their DBHs were estimated using an automatic stem diameter estimation algorithm proposed by Hyyppä et al.
-(2020b) and further developed by Hyyppä et al.
-(2020a) to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS.
-The method described by Hyyppä et al.
-(2020a) was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
+In this study, trees were detected, and their DBHs were estimated using an automatic stem diameter estimation algorithm proposed by hyyppa2020accurate and further developed by hyyppa2020under to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS.
+The method described by hyyppa2020under was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
 The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments and 5-s-long temporal segments.
 The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data.
-The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al.
-1996)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle.
+The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (ester1996DBSCAN)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle.
 The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA).
@@
 Finally, the DBH estimates were calculated by interpolating the stem curve at a height of 1.3 m above the ground.
-For a detailed description of the method, the reader is referred to Hyyppä et al.
-(2020b, a).
+For a detailed description of the method, the reader is referred to hyyppa2020accurate; hyyppa2020under.
 3.7 Performance assessment
@@
 The trajectory was considered smooth if the system did not perform any emergency stops during the flight.
-The accuracy of the VIO localization was evaluated by calculating the Absolute Trajectory Error (ATE), as proposed by Sturm et al.
-(2012), with respect to the post-processed SfM reference trajectories.
+The accuracy of the VIO localization was evaluated by calculating the Absolute Trajectory Error (ATE), as proposed by sturm2012ate, with respect to the post-processed SfM reference trajectories.
 In this study, ATE was calculated only for the position estimate, and the error in the orientation estimate was not evaluated.
-ATE was calculated using the trajectory evaluation toolbox by Zhang and Scaramuzza (2018), which aligns the trajectory estimates with the reference trajectories with the method by Umeyama (1991).
+ATE was calculated using the trajectory evaluation toolbox by rpgToolbox, which aligns the trajectory estimates with the reference trajectories with the method by Umeyma1991.
 After aligning the trajectories, the toolbox calculates the RMSE of the individual state errors to generate a single metric error value representing the accuracy of the entire trajectory.
@@
 where $\Delta\mathbf{p}_{i}$ is the error between the estimated position and reference trajectory position in a single state.
-In the first tests conducted in a simulator as described by (Karjalainen et al.
-2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to a constant, compared to when the online estimation method (Qin and Shen 2018) was used.
+In the first tests conducted in a simulator as described by (karjalainen2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to a constant, compared to when the online estimation method (onlineTd) was used.
 For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr.
@@
 ’Evo-medium’ Old Spruce forest 650 34 - 36 7/7 5/7
-’Evo-difficult’ Young Spruce forest 2000 42 8/9 0/8
+’Evo-difficult’ Young Spruce forest 2000 42 8/9 0/9
 (a) ’Evo-medium’
@@
 The proposed methodology and the accuracy of the DBH estimates obtained were evaluated by comparing them with other relevant studies (Section 2.1, Table 1) considering the precision, autonomy of the system, weight, and complexity of the test environment.
-The highest estimation accuracies in the literature have been obtained in studies using either considerably heavier and larger ULS systems Hyyppä et al.
-(2020a, c); Muhojoki et al.
-(2024) operated in sparse forest environments or in environments with few reference trees Shimabuku, Konoshima, and Ota (2023).
-The estimation results obtained in this study were close to those obtained in most other studies utilizing camera-based under-canopy drone systems (Kuželka and Surovỳ 2018; He et al.
-2025; Krisanski, Taskhiri, and Turner 2020).
-Furthermore, other studies based on SfM (Shimabuku, Konoshima, and Ota 2023; Kuželka and Surovỳ 2018; He et al.
-2025; Krisanski, Taskhiri, and Turner 2020) had more complex flight paths that allowed the full reconstruction most of the tree trunks, whereas in this study, the trunks were only partially reconstructed from image data collected from a single flight line.
-Considering the size of the systems, miniaturized lightweight systems were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al.
-(2024), Cheng et al.
-(2024), He et al.
-(2025), and Liang et al.
-(2024), though the LiDAR-based custom drone systems (Prabhu et al.
-2024; Cheng et al.
-2024; He et al.
-2025; Liang et al.
-2024) were still considerably heavier and larger than the camera-based drone systems.
-From these, LiDAR-based flights in Cheng et al.
-(2024) and Liang et al.
-(2024) were the only ones where the systems were reported to fly without commands from the pilot during the flight.
-Considering the complexity of the environments, the density of the forest in Liang et al.
-(2024) (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha).
-In Cheng et al.
-(2024), the density of the forest was not reported.
-In summary, the approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a level of estimation close to the accuracy of most of the studies with SfM drone systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al.
-2025).
-In particular, existing studies based on SfM of drones were using commercial drone systems piloted manually, except He et al.
-(2025) where the camera was added as an additional payload to a custom LiDAR-based drone platform, and the flight mode was not reported.
-Compared to studies with LiDAR-based miniaturized autonomous drone data collection, the accuracy of the estimation of DBH obtained in this study was higher than in Liang et al.
-(2024) and close to that in Cheng et al.
-(2024) even without any GCPs.
-However, the system utilized in Cheng et al.
-(2024) has yielded slightly better accuracy with manually piloted flights (Cheng et al.
-2024; Prabhu et al.
-2024), suggesting that their LiDAR-based system has the potential for higher accuracy of estimation by optimizing autonomous flight trajectories.
+The highest estimation accuracies in the literature have been obtained in studies using either considerably heavier and larger ULS systems hyyppa2020under; hyyppa2020comparison; muhojoki24BENCHMARKING operated in sparse forest environments or in environments with few reference trees shimabuku2023diameter.
+The estimation results obtained in this study were close to those obtained in most other studies utilizing camera-based under-canopy drone systems (kuvzelka2018mapping; he2025estimating360; krisanski2020photodrone).
+Furthermore, other studies based on SfM (shimabuku2023diameter; kuvzelka2018mapping; he2025estimating360; krisanski2020photodrone) had more complex flight paths that allowed the full reconstruction most of the tree trunks, whereas in this study, the trunks were only partially reconstructed from image data collected from a single flight line.
+Considering the size of the systems, miniaturized lightweight systems were used in the studies by shimabuku2023diameter, krisanski2020photodrone, kuvzelka2018mapping, prabhu2024treeFalcon, cheng2024Treescope, he2025estimating360, and liang2024forest, though the LiDAR-based custom drone systems (prabhu2024treeFalcon; cheng2024Treescope; he2025estimating360; liang2024forest) were still considerably heavier and larger than the camera-based drone systems.
+From these, LiDAR-based flights in cheng2024Treescope and liang2024forest were the only ones where the systems were reported to fly without commands from the pilot during the flight.
+Considering the complexity of the environments, the density of the forest in liang2024forest (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha).
+In cheng2024Treescope, the density of the forest was not reported.
+In summary, the approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a level of estimation close to the accuracy of most of the studies with SfM drone systems (kuvzelka2018mapping; krisanski2020photodrone; he2025estimating360).
+In particular, existing studies based on SfM of drones were using commercial drone systems piloted manually, except he2025estimating360 where the camera was added as an additional payload to a custom LiDAR-based drone platform, and the flight mode was not reported.
+Compared to studies with LiDAR-based miniaturized autonomous drone data collection, the accuracy of the estimation of DBH obtained in this study was higher than in liang2024forest and close to that in cheng2024Treescope even without any GCPs.
+However, the system utilized in cheng2024Treescope has yielded slightly better accuracy with manually piloted flights (cheng2024Treescope; prabhu2024treeFalcon), suggesting that their LiDAR-based system has the potential for higher accuracy of estimation by optimizing autonomous flight trajectories.
 Based on the results, the potential estimation accuracy with a low-cost stereo camera is considerably lower than that with heavy and expensive survey LiDARs but close to the quality of smaller LiDARs suitable for miniaturized drones capable of flying inside dense forests.
@@
 Blind sides in the obstacle avoidance map could be tackled either by sensors with 360° FOV (e.g., 360° camera or multi-camera system) or by storing historical data in the occupancy grid map in addition to the current visible area.
-The virtual floor and ceiling of Ego-planner-V2 (Zhou et al.
-2022) are also defined as static with respect to the takeoff location of the drone, limiting low-altitude flying to relatively flat areas.
+The virtual floor and ceiling of Ego-planner-V2 (zhou2022swarm) are also defined as static with respect to the takeoff location of the drone, limiting low-altitude flying to relatively flat areas.
 To capture images at the DBH estimation altitude in uneven forest plots, the flight altitude limits should be modified to be dynamic with respect to the local terrain level.
@@
 In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate.
-Potential methods for GNSS-Fusion include utilizing the global fusion features in VINS-Fusion (Qin et al.
-2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g., GVINS (Cao, Lu, and Shen 2022).
+Potential methods for GNSS-Fusion include utilizing the global fusion features in VINS-Fusion (qin2019global) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g., GVINS (Cao2022GVINS).
 Although the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forest.
@@
 For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides.
-In the study by He et al.
-(2025), the DBH estimation error increased significantly when only part of the shape of the tree trunk was observed.
+In the study by he2025estimating360, the DBH estimation error increased significantly when only part of the shape of the tree trunk was observed.
 This suggests that reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the DBH estimate errors obtained with the system utilized in this study.
 In particular, this is expected to improve the quality of the estimation for trees of large diameter that yielded a higher RMSE in this study than trees with a DBH of less than 30 cm.
-The global path covering the entire test plot could be implemented with a simple waypoint generation algorithm, such as Liang et al.
-(2024), or with more advanced exploration methods such as those proposed by Zhou et al.
-(2021) or Cieslewski, Kaufmann, and Scaramuzza (2017).
-Another way to achieve area coverage would be utilizing a swarm of drones, as in the original article by Zhou et al.
-(2022).
-Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage.
+The global path covering the entire test plot could be implemented with a simple waypoint generation algorithm, such as liang2024forest, or with more advanced exploration methods such as those proposed by Zhou2021FUEL or Cieslewski2017exploration.
+Another way to achieve area coverage would be utilizing a swarm of drones, as in the original article by zhou2022swarm.
+Furthermore, recent exploration algorithms (Zhou2023RACER)(Bartolomei2023swarmExploration) are able to utilize drone swarms to further improve and speed up the area coverage.
 In potential follow-up studies, the DBH estimation potential of the system could also be evaluated in alternative forest environments, for example, in sparser forest environments.
@@
 However, Intel RealSense D435 is also capable of generating 3D point clouds in real time.
-Furthermore, image-based object detection can be performed in real time, for example, with YOLO (Redmon et al.
-2016).
+Furthermore, image-based object detection can be performed in real time, for example, with YOLO (Redmon2016yolo).
 While only computationally light analysis can be conducted on the onboard computer, real-time analysis with a higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.
-Acknowledgements
+Funding
 This research was funded by the Research Council of Finland within projects ’Learning techniques for autonomous drone based hyperspectral analysis of forest vegetation’ (decision no.
@@
 Part of the data recorded during the field tests is available from the corresponding author upon reasonable request.
+Declaration of AI-assisted technologies in the writing process
+During the revision of this work, the author(s) used free versions of Grammarly and Writefull (included in the Overleaf LaTeX editor) to improve the writing style in a non-native language manuscript.
+These tools were only used to detect grammar mistakes and improve the clarity of the text.
+All suggestions of these writing assistants were manually accepted by the author(s) and no text was autogenerated with generative AI tools.
+After using these tools, the author(s) reviewed the content and take(s) full responsibility for the content of the publication.